### v1

# Kairos: Towards Adaptive and Generalizable Time Series Foundation Models

Abstract
Time series foundation models (TSFMs) have emerged as a powerful paradigm for time series analysis, driven by large-scale pretraining on diverse data corpora. However, time series inherently exhibit heterogeneous information density over time, influenced by system states and signal complexity, presenting significant modeling challenges especially in a zero-shot scenario. Current TSFMs rely on non-adaptive processing pipelines that fail to capture this dynamic nature. For example, common tokenization strategies such as fixed-size patching enforce rigid observational granularity, limiting their ability to adapt to varying information densities. Similarly, conventional positional encodings impose a uniform temporal scale, making it difficult to model diverse periodicities and trends across series. To overcome these limitations, we propose Kairos, a flexible TSFM framework that integrates a dynamic patching tokenizer and an instance-adaptive positional embedding. Kairos adaptively selects tokenization granularity and tailors positional encodings to the unique characteristics of each time series instance. Trained on a large-scale Predictability-Stratified Time Series (PreSTS) corpus comprising over 300 billion time points and adopting a multi-patch prediction strategy in the inference stage, Kairos achieves superior performance with much fewer parameters on two common zero-shot benchmarks, GIFT-Eval and the Time-Series-Library benchmark, consistently outperforming established methods across diverse tasks. The project page is at https://foundation-model-research.github.io/Kairos.
1 Introduction
Time series forecasting is a core component of many real-world applications, including cloud services, traffic management, retail, finance, and energy (Box et al., 2015; Hyndman & Athanasopoulos, 2018). Although time series data are ubiquitous, many individual datasets suffer from scarcity, making it difficult to train models independently, particularly for deep neural networks that require large amounts of data. This challenge has long motivated the development of few-shot and zero-shot forecasting methods (Oreshkin et al., 2020). More recently, inspired by the success of large-scale foundation models (FMs) (Achiam et al., 2023; Kirillov et al., 2023), researchers have turned to building general-purpose time series foundation models (TSFMs). These models leverage vast and diverse time series corpora to acquire strong generalization capabilities, showing promising results across a wide range of downstream tasks (Ansari et al., 2024; Woo et al., 2024; Shi et al., 2024).
Despite this progress, a key limitation persists: current TSFMs rely on non-adaptive processing pipelines that fail to reflect the dynamic and heterogeneous nature of time series. As shown in Figure 1(b) and exemplified in Figure 1(c), the statistical information density measured via spectral entropy on sliding windows (see Appendix D) varies not only across datasets but also across samples within the same series. However, existing models typically impose a uniform and invariant scale across all series, thereby overlooking internal variability. This holds true whether they tokenize each time point individually (Ansari et al., 2024; Shi et al., 2024), tokenize using fixed-size patches (Das et al., 2024), or model temporal dependencies using standard positional encodings (Liu et al., 2025).
This non-adaptive paradigm gives rise to two major challenges. (i) Local information extraction: Current tokenization strategies enforce fixed granularity, requiring the model to observe whole series with the same temporal resolution. Such rigidity prevents adaptation to heterogeneous or fluctuating information density, both across different series and within a single sequence. For instance, models cannot zoom in on fine-grained details during sudden market shocks, nor can they abstract more efficiently when the data remain stable. (ii) Temporal relationship modeling: Conventional positional encodings impose a unified temporal scale across sequences, neglecting differences in periodicity, trend and seasonality. As a result, TSFMs struggle to adapt to heterogeneous domains, for example, power consumption data measured hourly versus retail sales data measured daily, each exhibiting distinct temporal dynamics. Figure 1(b) and 1(d) illustrate these challenges, highlighting both the variability of information density within time series and the rigidity of current methods.
Figure 1: (a) The trade-off between performance (normalized MASE) and the number of parameters on GIFT-Eval benchmark (Aksu et al., 2024) for existing TSFMs. Our Kairos achieves a superior performance at a comparable parameter scale. (b) (c) Significant variation exists in information density across and within different time series datasets. (d) Existing TSFMs primarily use tokenization methods like point-wise or fixed-size patching, while our Kairos utilizes a Mixture of Patch Tokenization to address dynamic changes in information density.
To address these challenges, we propose Kairos, a foundation model designed to handle the complexities of heterogeneous time series data. The name Kairos is derived from its mythological significance as the Greek god of the “opportune moment”, reflecting its adaptability. For local information extraction, we introduce a Mixture-of-Size Dynamic Patching (MoS-DP) mechanism that adaptively models time series at multiple granularities by selecting patch sizes based on local information density. Inspired by the computationally free null experts in Mixture-of-Experts architectures (Zeng et al., 2024; Jin et al., 2024), we further incorporate null patch sizes, which allow the model to flexibly vary the number of active patches. This design moves beyond static patching strategies, enabling more expressive modeling of nuanced, time-varying dynamics. For temporal relationship modeling, we propose Instance-adaptive Rotary Position Embedding (IARoPE), an extension of RoPE (Su et al., 2024), tailored for diverse temporal dynamics. Unlike standard RoPE, which represents positional information with a fixed temporal scale, our IARoPE modulates these scales using spectral features extracted from each series, thereby generating positional encodings uniquely adapted to individual instances.
In addition to our core architectural innovations, we further introduce two key optimizations in the inference and training phases, respectively. For inference, we propose a multi-patch prediction strategy that simultaneously forecasts multiple future patches. This approach enhances forecasting robustness by mitigating cumulative errors and offers greater flexibility for variable-length prediction horizons. To further enhance training, we construct the Predictability-Stratified Time Series (PreSTS) corpus, a large-scale and diverse pretraining dataset curated through a targeted sampling strategy. By prioritizing more predictable sequences while maintaining broad coverage, PreSTS provides high-quality supervision that supports efficient model scaling. As demonstrated by the results on the GIFT-Eval benchmark (Aksu et al., 2024) in Figure 1(a), these combined innovations allow our Kairos to achieve superior performance while using fewer parameters. This strong performance is consistently observed on Time-Series-Library (TSLib) (Wang et al., 2024b) benchmarks as well. In summary, our main contributions are as follows:
  • We propose Kairos, a novel foundation model specifically designed to handle dynamically changing time series data. Kairos demonstrates strong resilience to commonly observed data variability in real-world time series.
  • The two key architectural innovations introduced, i.e., MoS-DP and IARoPE, address the challenges of varying information density and instance-specific temporal dynamics, thereby enhancing the generalization capability of foundation models for time series.
  • Through extensive empirical evaluation, we demonstrate the superior performance of Kairos with significantly fewer parameters on zero-shot forecasting benchmarks, highlighting its flexibility and adaptability across a wide range of diverse and non-stationary time series datasets.
2 Related Work
Time Series Foundation Models (TSFMs). The scarcity of domain-specific time series data has long motivated research into models with strong few-shot or zero-shot learning capabilities (Oreshkin et al., 2020; 2021). Recently, inspired by the strong generalization of large language models, researchers are devoting more effort to TSFMs leveraging vast data corpora. Some works attempt to directly transfer the sequence modeling capabilities of large language models to the time series domain (Gruver et al., 2023). Others leverage the scalability of Transformer (Vaswani et al., 2017) and directly train unified foundation models on extensive time series corpora (Ansari et al., 2024; Das et al., 2024; Woo et al., 2024; Cohen et al., 2025; Liu et al., 2025). A few studies have also explored lightweight architectures, including Multi-Layer Perceptrons (MLPs) such as TTM (Ekambaram et al., 2024). However, these models often apply fixed and unified tokenization strategies and position embedding to time series from diverse domains with distinct characteristics, limiting their cross-domain generalization. Kairos, in contrast, employs a dynamic patching tokenizer and instance-adaptive position embedding to enhance its generalizability for diverse time series data.
Time Series Tokenization. Leveraging Transformer model backbone requires an effective mechanism for local information extraction, i.e., to transform raw data into sequential elements (tokens), discretely or continuously (Sennrich et al., 2016; Agarwal et al., 2025). The intuitive approach for time series tokenization is to map each time point to a token. However, models employ this approach (Ansari et al., 2024; Rasul et al., 2023; Shi et al., 2024) suffer from sensitivity to noise and computational inefficiency. To address this, Nie et al. (2023) introduces a fixed-size patching method that became a widely adopted paradigm in many subsequent TSFMs (Woo et al., 2024; Liu et al., 2025). However, the fixed-size patches failed to flexibly adapt to time series with varying information densities. While models like Pathformer (Chen et al., 2024) and ElasTST (Zhang et al., 2024) attempt to address these limitations by improving feature fusion strategies, they still simply partition the entire sequence into a few predefined patch sizes, failing to dynamically adjust to the varying information densities in single time series data. Furthermore, as they are designed for dataset-specific training rather than as foundation models, they lack cross-domain generalization capabilities. In contrast, Kairos improves upon the existing methods by dynamically selecting appropriate patch sizes for each segment of a time series, rather than relying on one or more fixed patch sizes for the entire sequence.
Position Embedding. Due to the position-unaware characteristic of self-attention mechanism, Transformers relies on position embeddings (PE) to model temporal information. Most TSFMs adopt existing designs for PE in Natural Language Processing (NLP) domains (Das et al., 2024; Shi et al., 2024), which typically emphasize long-term decay through methods like sinusoidal calculations or by suppressing long-range positional information (Vaswani et al., 2017; Su et al., 2024). By imposing a uniform temporal scale, these PEs struggle to model the heterogeneous temporal relationships of time series. Existing adaptive PE methods in NLP (Zheng et al., 2024; Lin et al., 2024) are also insufficient, as they target context extrapolation rather than the complex temporal structures of time series. In the time series domain, ElasTST (Zhang et al., 2024) proposed a tunable RoPE to better adapt to time series data. However, its adaptation uses per-dataset training rather than dynamic adjustment. To address this limitation, we propose to dynamically modulate PE tailored to the intrinsic characteristics of each input time series, enhancing the effectiveness of time series foundation models.
Figure 2: The architecture of KAIROS, which including (1) Mixture-of-Size Dynamic Patching (MoS-DP): This module adaptively tokenizes the time series by fusing features from multiple granularities. As detailed in the expanded view, a Dynamic Patch Router first selects active experts (each corresponding to a patch size) for a coarsest patch and determines the finest patch size $p_{k}$ for tokenization. The final embedding for each resulting finest patch is then created via a hierarchical fusion process: it is the weighted sum of outputs from all activated experts that correspond to the granularity of the finest patch itself or its ancestors. For instance, the diagram illustrates how the embedding for a finest patch of size $p_{1}$ (marked ①) aggregates information from its ancestors of size $p_{2}$. This results in a sequence of fused embeddings rich with multi-scale information that is fed into the encoder. (2) Instance-Adaptive RoPE (IARoPE): This module (left) adjusts positional encodings for the Transformer by modulating them based on the unique frequency profile of each input series.
3 Methodology
3.1 Overall Framework
We present the forecasting problem formulation: given the historical observations of a time series $\bm{x}_{1:T}=(x_{1},\ldots,x_{T})\in\mathbb{R}^{T}$, the objective is to forecast the future values $\bm{x}_{T+1:T+H}=(x_{T+1},\ldots,x_{T+H})\in\mathbb{R}^{H}$, where $T$ is the lookback window and $H$ denotes the forecast horizon.
The architecture of Kairos, shown in Figure 2, consists of three main components. First, the input time series is tokenized into a sequence of embeddings by our Mixture-of-Size Dynamic Patching (MoS-DP) module to extract multi-granularity local information. These embeddings are then processed by a Transformer encoder, which uses our proposed Instance-Adaptive Rotary Position Embedding (IARoPE) to model complex temporal relationships. Finally, a Transformer decoder utilizes a multi-patch prediction strategy for forecasting. The details of each key component, including MoS-DP, IARoPE, and the multi-patch prediction strategy, are presented in the following sections.
3.2 Mixture-of-Size Dynamic Patching (MoS-DP)
In contrast to time-series foundation models such as Time-MoE (Shi et al., 2024) and Moirai-MoE (Liu et al., 2024b), which directly replace the original feed-forward network (FFN) layers in Transformer blocks with Mixture-of-Experts (MoE) modules (Jacobs et al., 1991; Shazeer et al., 2017), Kairos takes inspiration from MoE to tackle the intrinsic heterogeneity of time series. To improve local information extraction, we introduce the MoS-DP module, which performs routing at the level of each time series patch and thereby effectively addresses the challenge of adapting foundation models to diverse datasets with varying information granularities. MoS-DP unfolds in three stages: Patchify at Coarsest Granularity, Dynamic Patch Router, and Dynamic Patch Fusion.
Patchify at Coarsest Granularity. In this stage, the entire time series is partitioned based on the maximum patch size. We begin with an initial partition of the sequence. A set of $S$ patch sizes is defined in advance, denoted as $\{p_{1},\ldots,p_{S}\}\in\mathbb{R}^{S}$, where $p_{1}<p_{2}<\ldots<p_{S}$. We first tokenize the input time series $\bm{x}_{1:T}\in\mathbb{R}^{T}$ using non-overlapping patching using the coarsest patch size $p_{S}$. The series is divided into $N=\lceil\frac{T}{p_{S}}\rceil$ coarsest patches. If the original series length $T$ is not an exact multiple of $p_{S}$, $\bm{x}_{1:T}$ is augmented with left-padding of zeros as needed. This process yields a sequence of $N$ coarsest patches, $\{\bm{x}_{n}^{p_{S}}\}_{n=1}^{N}$, where each patch has a uniform length of $p_{S}$.
Dynamic Patch Router. To perform the final multi-granularity information fusion, we select a set of appropriate granularities for each coarsest patch $\bm{x}_{n}^{p_{S}}$. Inspired by the MoE paradigm, we employ multiple FFNs as experts, with each expert dedicated to extracting features at a unique granularity level from the set $\{p_{1},\ldots,p_{S}\}$. Furthermore, we introduce $Z$ null experts, which perform no computation when selected, allowing the number of activated experts to be dynamic for each input, facilitating more flexible modeling of heterogeneous time series. The routing process is formulated as
$g_{n,i}=\begin{cases}s^{\prime}_{n,i},&s^{\prime}_{n,i}\in\mathrm{TopK}(\{s^{\prime}_{n,j}|1\leqslant j\leqslant S+Z\},K)\ \text{and}\ 1\leqslant i\leqslant S\\ 0,&\text{otherwise.}\end{cases}$ (1)
$s^{\prime}_{n,i}=\frac{\exp(s_{n,i}+\bm{b}_{i})}{\sum_{j=1}^{S}\exp(s_{n,j}+\bm{b}_{j})},\quad\text{where }s_{n,i}=\mathbf{W}_{i}\bm{x}_{n}^{p_{S}}.$ (2)
Here $g_{n,i}$ represents the gating value for the $i$-th expert with respect to the $n$-th coarsest patch; $s_{n,i}$ is the affinity score between patch and the expert, and $\mathbf{W}_{i}$ is the trainable weight matrix for the $i$-th expert. The bias term $\bm{b}_{i}$ is utilized for an Auxiliary-Loss-Free Load Balancing method (Liu et al., 2024a) and is discussed in Appendix C.3.1.
The coarsest patch $\bm{x}_{n}^{p_{S}}$ will then be partitioned into smaller finest patches. The size of these finest patches, $p_{k}$ is determined by the finest granularity selected by the router for this specific coarsest patch, i.e., $p_{k}=\min\{p_{i}|g_{n,i}>0\}$. Note that this finest granularity can differ for each coarsest patch $\bm{x}_{n}^{p_{S}}$. All $N$ coarsest patches are re-segmented in parallel based on their respective routing decisions. Compared to uniformly segmenting the entire sequence with a fixed size (Chen et al., 2024; Zhang et al., 2024) as discussed in Section 2, our method offers greater flexibility to adapt to temporal sequences with varying information densities.
Dynamic Patch Fusion. In the final stage, we yield the embedding for each finest patch by fusing information from multiple granularities. This process ensures that the final token representations are aware of the context at various temporal scales. Specifically, for the $n$-th coarsest patch $\bm{x}_{n}^{p_{S}}$, which was determined in the previous stage, it is first partitioned into $M_{n}=\frac{p_{S}}{p_{k}}$ non-overlapping finest patches. Here, $p_{k}$ is the finest patch size selected by the Dynamic Patch Router for this coarsest patch. We denote the $m$-th finest patch within the $n$-th coarsest patch as $\bm{x}_{n,m}^{p_{k}}$. Each such finest patch corresponds to a single token for the subsequent Transformer encoder. The embedding $e_{n,m}$ for the finest patch $\bm{x}_{n,m}^{p_{k}}$ is calculated as the weighted summation of semantic information originating from all activated granularities, formulated as:
$\bm{e}_{n,m}=\sum_{i=k}^{S}\alpha_{n,i}\cdot\text{MLP}_{i}(\mathcal{A}(x_{n,m}^{p_{k}},p_{i})),\ \text{where}\ \ \alpha_{n,i}=\frac{g_{n,i}}{\sum_{j=k}^{S}g_{n,j}},\ \ \bm{e}_{n,m}\in\mathbb{R}^{D_{h}},$ (3)
where $\mathcal{A}(x_{n,m}^{p_{k}},p_{i})$ identifies the ancestor patch of size $p_{i}$ that contains the finest patch $x_{n,m}^{p_{k}}$, as illustrated in Figure 2. We provide a formal definition of $\mathcal{A}(\cdot,\cdot)$ and a detailed example in Appendix C.3.2. $\text{MLP}_{i}$ denotes the expert that extracts the embedding for patch size $p_{i}$. The term $\alpha_{n,i}$ represents the normalized gating value from the router for the $n$-th coarsest patch, which dynamically weighs the importance of each granularity. $D_{h}$ is the hidden dimension. Following Eq. 3, $M$ embeddings are extracted from the coarsest patch $\bm{x}_{n}^{p_{S}}$ via the Dynamic Patch Fusion (DPF) module as:
$\text{DPF}(\bm{x}_{n}^{p_{S}},\bm{g}_{n})=[\bm{e}_{n,1},\bm{e}_{n,2},\ldots,\bm{e}_{n,M_{n}}]\in\mathbb{R}^{M_{n}\times D_{h}},\quad\bm{g}_{n}=(g_{n,1},g_{n,2}\ldots,g_{n,S}).$ (4)
Eventually, the original time series $\bm{x}_{1:T}$ is transformed into a sequence of embeddings by concatenating all DPF module outputs as $\mathbf{E}=\text{DPF}(\bm{x}^{p_{S}}_{1},\bm{g}_{1})||\text{DPF}(\bm{x}^{p_{S}}_{2},\bm{g}_{2})||\ldots||\text{DPF}(\bm{x}^{p_{S}}_{N},\bm{g}_{N})$ with sequence length $T^{\prime}=\sum_{n=1}^{N}M_{n}$, where $M_{n}$ represents the number of finest patches that the $n$-th patch in $\bm{x}^{p_{S}}_{1:N}$ is segmented into by the Dynamic Patch Router. This final sequence $\mathbf{E}\in\mathbb{R}^{T^{\prime}\times D_{h}}$ would serve as the input to the subsequent Transformer module of Kairos. This design allows the model to use small patches for in-depth analysis where local details are rich and large patches for efficient modeling where information is stable, as demonstrated in Section 4.4.1.
3.3 Instance-Adaptive Rotary Position Embedding (IARoPE)
Rotary Position Embedding (RoPE) (Su et al., 2024) encodes relative positions by rotating hidden vectors within Transformer layer, such as a query ($\bm{q}$) or a key ($\bm{k}$), which we denote generically as $\bm{z}$. The operation is formulated as $f_{\text{RoPE}}(\bm{z},t)=(\bm{z}_{2j}+i\bm{z}_{2j+1})e^{it\theta_{\text{init},j}}$, where the base RoPE frequencies $\theta_{\text{init},j}=b^{-2j/D_{h}}$ are determined by a fixed hyperparameter $b$. A detailed derivation of the RoPE calculation is provided in Appendix C.4. However, these fixed frequencies impose a single, uniform temporal scale across all data. This rigidity prevents the model from effectively modeling temporal relationships for series with diverse intrinsic structures, such as different seasonalities or trends. To address this limitation, we propose Instance-Adaptive Rotary Position Embedding (IARoPE), which dynamically tailors customized positional encodings for each time series instance.
The core idea of IARoPE is to modulate the base RoPE frequencies $\theta_{\text{init}}$ using spectral features extracted from each instance. Specifically, Given an input time series instance $\mathbf{x}_{1:T}$, we first extract its frequency features by applying the Fast Fourier Transform (FFT) and select the first $\omega$ low-frequency components (their magnitudes) to form a feature vector $\mathbf{x}_{\text{fft}}$, where $\omega$ is a hyperparameter. We focus on low frequencies as they often capture dominant trends and periodic patterns crucial for understanding the temporal structure of instance (Zhou et al., 2022). Then we feed $\mathbf{x}_{\text{fft}}$ into a small MLP to predict instance-specific modulation parameters: a scaling factor $\mathbf{\gamma}\in\mathbb{R}^{D_{h}/2}$ and a bias factor $\mathbf{\beta}\in\mathbb{R}^{D_{h}/2}$. These parameters then adapt the base RoPE frequencies $\theta_{\text{init}}\in\mathbb{R}^{D_{h}/2}$ via an element-wise affine transformation to produce the adaptive frequencies $\theta^{\prime}\in\mathbb{R}^{D_{h}/2}$:
$\theta^{\prime}_{j}=\gamma_{j}\odot\theta_{\text{init},j}+\beta_{j}\quad\text{for }j=0,1,\ldots,D_{h}/2-1,$ (5)
where $\odot$ denotes the Hadamard product. The resulting $\theta^{\prime}$ incorporates learned information from the instance’s frequency characteristics. Finally, these adaptive angles $\theta^{\prime}$ replace the fixed $\theta$ in the RoPE calculation for the current instance:
$f_{\text{IARoPE}}(\bm{z},t)=(\bm{z}_{2j}+i\bm{z}_{2j+1})e^{it\theta^{\prime}_{j}},$ (6)
This mechanism adapts the positional encoding to the frequency profile of each time series instance, potentially offering a more accurate representation of relative temporal positions. We demonstrate the effectiveness of this adaptation in Section 4.4.2.
3.4 Multi-Patch Prediction
Kairos employs multi-patch prediction to effectively mitigate the cumulative errors inherent in autoregressive processes. Specifically, after the encoder outputs the hidden representation $\bm{h}^{\text{en}}\in\mathbb{R}^{T^{\prime}\times D_{h}}$, a Transformer decoder is employed to forecast future values. We introduce a set of $J$ learnable forecast tokens, denoted as $Q_{\text{forecast}}=\{o_{1},\dots,o_{J}\}$. These tokens are passed as input to decoder, which processes them simultaneously to produce a sequence of output embeddings $\bm{h}^{\text{de}}\in\mathbb{R}^{J\times D_{h}}$:
$\bm{h}^{\text{de}}=\text{Decoder}(Q_{\text{forecast}},\bm{h}^{\text{en}}).$ (7)
Each output embedding $\bm{h}^{\text{de}}_{j}$ is then passed through a shared prediction head (a residual FFN block (Das et al., 2024)) to forecast a patch of the future of length $H_{s}=H/J$ as
$\hat{\bm{x}}_{T+jH_{s}+1:T+(j+1)H_{s}}=\text{ResidualFFN}(\bm{h}^{\text{de}}_{j}),\quad j=\{1,\dots,J\}$ (8)
This multi-token design is motivated by the limitations of two common forecasting paradigms. First, compared to conventional single-step autoregressive models (Ansari et al., 2024; Liu et al., 2024c), predicting $J$ patches at once enhances computational efficiency and mitigates cumulative error by reducing iterative steps. Second, unlike models that forecast a larger patch (Das et al., 2024), our approach provides greater flexibility for shorter prediction horizons. It can construct a precise forecast by aggregating outputs from the necessary number of tokens, avoiding the potentially imprecise truncation of a larger-than-needed output. We provide a detailed analysis of this design in Appendix A.
3.5 Training Corpus
For pre-training, we curated Predictability-Stratified Time Series (PreSTS) corpus, a corpus of over 300 billion time points. PreSTS comprises real-world time series spanning multiple domains and is augmented by a synthetic dataset we constructed to ensure comprehensive coverage. Additionally, we have excluded all test sets found within the GIFT-Eval benchmark (Aksu et al., 2024) to maintain evaluation integrity. While a prevailing consensus regarding TSFMs highlights the necessity of a diverse training dataset, research has also shown that the presence of anomalies in the training set can severely degrade a model’s predictive capabilities (Cheng et al., 2024). Therefore, the real-world datasets in PreSTS are stratified into five tiers based on their degree of predictability. Datasets exhibiting higher predictability are assigned a greater sampling probability during training, a strategy that enables the model to undergo a more robust, high-quality training regimen without sacrificing the comprehensiveness of the data. Comprehensive descriptions of our training datasets along with synthetic data generation algorithm, are provided in Appendix E.1.
4 Experiment
In this section, we conduct a comprehensive evaluation to assess the performance of Kairos and the effect of proposed components. Specifically, we investigate the following three key research questions. RQ1: Does Kairos achieve better generalization and zero-shot prediction than previous methods? (Section 4.2) RQ2: How effectively does MoS-DP handle time series heterogeneity by adapting its patch size according to local information densities? (Section 4.3 and 4.4.1) RQ3: Does IARoPE effectively generate customized position embedding tailored to the unique temporal structure of each time series instance? (Section 4.3 and 4.4.2)
4.1 Experiment Settings
4.1.1 Evaluation Datasets and Metric
To comprehensively evaluate the performance of our proposed method, we conducted zero-shot assessments on two well-known public benchmarks: GIFT-Eval (Aksu et al., 2024) and Time-Series-Library (TSLib) (Wang et al., 2024b). We train Kairos in three sizes: mini (10M), small (23M) and base (50M). The training details are provided in Appendix C.1.
Benchmarks. GIFT-Eval benchmark consists of 97 prediction tasks, with 55 tasks dedicated to short-term predictions, 21 for medium-term predictions, and 21 for long-term predictions, providing a comprehensive means of assessing the model’s predictive capabilities. For the TSLib benchmark, we followed the common practice in Time-MoE (Shi et al., 2024) and selected the ETT datasets (Zhou et al., 2021a) and Weather (Zhou et al., 2021b). Additionally, to test the model’s adaptability to diverse sampling frequencies, we incorporated the Saugeen datasets (Godahewa et al., 2021), which include data with both daily and weekly sampling frequencies. These were referred to as Saugeen (D) and Saugeen (W), respectively. The detailed description of the evaluation datasets, their splitting configurations, and the selection of the evaluation length are provided in Appendix E.2 and E.3.
Metric. In our evaluation of GIFT-Eval, consistent with prior research (Auer et al., 2025; Liu et al., 2025), we employ the official Mean Absolute Scaled Error (MASE) and Continuous Ranked Probability Score (CRPS) metrics to assess point and probabilistic forecasting performance, respectively. These metrics are first normalized by the Seasonal Naïve baseline for each individual task. The final score is then computed as the geometric mean of these normalized values. For the TSLib benchmark, we followed previous works (Liu et al., 2023; 2024c; Nie et al., 2023) by adopting Mean Squared Error (MSE) and Mean Absolute Error (MAE) to evaluate time series forecasting performance. The metric calculations are detailed in Appendix C.2.1.
4.1.2 Compared Methods
We compare Kairos with several state-of-the-art models, including TSFMs such as Sundial (Liu et al., 2025), Toto (Cohen et al., 2025), YingLong (Wang et al., 2025), VisionTS (Chen et al., 2025), Time-MoE (Shi et al., 2024), ChronosBolt (Ansari et al., 2024), TTM (Ekambaram et al., 2024), Moirai (Woo et al., 2024), Timer-XL (Liu et al., 2024c), TimesFM-2.0 (Das et al., 2024), and Chronos (Ansari et al., 2024), as well as advanced full-shot deep learning models, including PatchTST (Nie et al., 2023), DLinear (Zeng et al., 2022), iTransformer (Liu et al., 2023), Pathformer (Chen et al., 2024), and TimesNet (Wu et al., 2022). .
4.2 Zero-shot Evaluation
Finding 1: Kairos demonstrates remarkable zero-shot forecasting capabilities across multiple benchmarks. On the GIFT-Eval benchmark, as presented in Table 1, $\textsc{Kairos}_{\text{base}}$ (50M) achieves the best MASE and the second-best CRPS when compared to state-of-the-art task-specific deep learning methods and other TSFMs. Notably, $\textsc{Kairos}_{\text{small}}$ (23M) surpasses both Toto and Sundial in MASE, despite having a parameter count that is 6.6 and 5.6 times smaller, respectively. On the TSLib benchmark, as shown in Figure 3, our lightweight $\textsc{Kairos}_{\text{mini}}$ (10M) surpasses the performance of both recent advanced TSFMs and the majority of full-shot deep learning models. Compared with other TSFMs, Kairos achieves a significant improvement in predictive performance.
Table 1: Performance evaluation on the GIFT-Eval. We evaluated model performance using the normalized MASE and CRPS metrics delineated in Section 4.1.1, where lower values indicate higher forecasting accuracy. The baseline models fall into three categories: statistical methods, deep learning (DL) models, and TSFMs. Following the official classification of GIFT-Eval, TSFMs are further subdivided into TestData Leakage and Zero-Shot. TestData Leakage signifies that the model’s training set included test data, whereas Zero-Shot indicates that the model was trained without any data from the test set or its corresponding training split. Baseline results are officially reported by GIFT-Eval.
Type Statistical DL (Full-Shot) TSFMs (TestData Leakage) TSFMs (Zero-Shot)
Method
Seasonal
Naïve
DLinear
PTST. TTM Chronos
Chronos
Bolt
TimesFM Moirai VisionTS Ying. Toto Sundial
$\textsc{Kairos}_{s}$
$\textsc{Kairos}_{b}$
#Params - - - 5M 709M 205M 500M 311M 112M 300M 151M 128M 23M 50M
MASE 1.000 1.061 0.849 1.020 0.870 0.808 0.758 0.875 0.863 0.798 0.750 0.750 0.748 0.742
CRPS 1.000 0.846 0.587 0.873 0.574 0.574 0.550 0.599 0.755 0.548 0.517 0.559 0.554 0.548
Figure 3: Zero-shot forecasting performance on TSLib. Results are averaged across prediction lengths {96, 192, 336, 720}. The subscripts $l$, $b$, $s$, and $a$ represent model sizes of large, base, small, and advanced, respectively. The complete experimental results are presented in Appendix F.
4.3 Ablation Study
Table 2: Ablation study comparing MoS-DP and RoPE variants. We evaluated ablation settings using the normalized MASE from Section 4.1.1, assessing performance on individual prediction horizons and in aggregate across all tasks.
MoS-DP RoPE Variant Short Medium Long AVG
IARoPE 0.719 0.759 0.789 0.742
IARoPE 0.725 0.779 0.813 0.755
RoPE 0.726 0.783 0.828 0.759
RoPE 0.731 0.784 0.829 0.763
Finding 2: MoS-DP and IARoPE collaboratively enable Kairos to adapt to heterogeneous time series, achieving powerful predictive performance. To investigate the effectiveness of our proposed MoS-DP and IARoPE,we conducted a comprehensive ablation study, with the results summarized in Table 2. For a fair comparison, Kairos without MoS-DP uses a fixed patch size of 64, which was determined to be optimal via hyperparameter search. For IARoPE, we compare against the original RoPE (Su et al., 2024), which employs a fixed, predefined $\theta$. The integration of MoS-DP and IARoPE into our method leads to the best performance on the 97 tasks of the GIFT-Eval benchmark, which validates the superiority of our designs. Furthermore, the improvement of Kairos gets more significant compared to Kairos without MoS-DP or IARoPE as horizon gets longer, exhibiting better long-term forecasting ability.
4.4 Model Analysis
4.4.1 Patch Distribution Analysis
Finding 3: Kairos demonstrates adaptive capability to varying information densities, enabling it to model complex temporal dynamics with appropriate granularity. For each coarsest patch, Kairos dynamically fuses information by selecting from a set of granularities (patch sizes). To quantify this dynamic routing, we calculate a weighted average patch size by multiplying the weight assigned to each granularity by its corresponding patch size. Consequently, a smaller value for this metric indicates the model’s preference for finer-grained processing of a given coarsest patch, while a larger value suggests a preference for coarser granularity. In Figure 4, this weighted average patch size is visualized by the background color, with darker shades corresponding to smaller values. As the figure illustrates, Kairos strategically employs smaller patch sizes for sequences with abrupt changes and larger patch sizes for stable sequences.
Figure 4: Patch size preferences in GIFT-Eval test datasets. Darker shades indicate a smaller weighted average patch size, signifying the model’s preference for finer-grained processing in that region.
4.4.2 IARoPE Analysis
Figure 5: Causal analysis of adaptive modulation by IARoPE. This experiment validates the criticality of matching positional encodings to the unique characteristics of each time series instance by disrupting or removing this adaptation. We test this by manipulating the RoPE frequencies $\theta$ under several conditions: IARoPE (standard), Intra-Dataset Shuffle ($\theta$ modulations permuted between different instances within the same dataset), Inter-Dataset Shuffle ($\theta$ modulations from instances of other datasets), and Fixed RoPE (no modulation).
Finding 4: Modulating RoPE frequencies $\theta$ instance-wise can indeed better model time series temporal relationships. To verify that the performance gains of IARoPE are indeed caused by its instance-specific RoPE frequencies $\theta$ modulation, we designed a series of control experiments that impair or ablate this mechanism by shuffling the modulation parameters (details in Figure 5; Appendix E.4). The results confirm IARoPE’s superior forecasting accuracy, demonstrating the advantage of tailoring PE to individual instance frequency characteristics. The Intra-Dataset Shuffle ($\theta$ modulations permuted between different instances within the same dataset) performed second best, suggesting that while instance-specific modulation is optimal, leveraging dataset-level frequency similarities still offers utility. Conversely, both Inter-Dataset Shuffle ($\theta$ modulations from instances of other datasets) and Fixed RoPE (no modulation) yielded poorer performance, highlighting that mismatched or static $\theta$ parameters are suboptimal as they fail to capture the specific temporal nature of each series.
5 Conclusions
We introduce Kairos, a time series foundation model that advances the design of modules specifically adapted to the unique characteristics of time series data. Kairos achieves this through key innovations, including Mixture-of-Size Dynamic Patching, which considers varying information densities for more precise and efficient modeling, and Instance-Adaptive RoPE, which provides more accurate positional information tailored to time series nuances. Our comprehensive experiments reveal superior performance achieved by Kairos, highlighting the importance of designing foundational models that are inherently aligned with the structural nuances of time series data.
Appendix A Analysis of Multi-Patch Prediction
In this section, we present an analysis of the multi-patch prediction strategy introduced in Section 3.4. During the training phase, we experimented with a range of forecast patch numbers, specifically $J=\{1,2,4,6,12\}$. The corresponding evaluation results on the GIFT-Eval benchmark are illustrated in Figure 6. Our observations reveal that when $J=1$, which corresponds to the conventional approach of predicting a single patch (Ansari et al., 2024; Liu et al., 2024c; Das et al., 2024), the model achieves optimal performance in short-term forecasting. However, this method necessitates multiple iterations of autoregressive prediction, leading to a significant degradation in performance for medium- and long-term forecasting.
Conversely, as we increase the forecast patch number $J$, the number of required autoregressive steps is markedly reduced. For instance, when $J=2$, the autoregression frequency for medium- and long-term predictions is halved compared to the $J=1$ case. This reduction yields a substantial improvement in forecasting accuracy over these longer horizons, demonstrating that our proposed multi-patch prediction strategy effectively mitigates the cumulative error inherent in autoregressive processes for medium- and long-term forecasting.
Nevertheless, we noted that the forecasting performance does not improve indefinitely with an increasing the forecast patch number $J$. We attribute this phenomenon to the escalated difficulty of the prediction task, which hinders the model’s ability to optimize effectively. Ultimately, by evaluating the mean normalized MASE, we identified the optimal trade-off, selecting a forecast patch number of $J=2$ for Kairos.
Figure 6: Performance analysis of multi-patch prediction on the GIFT-Eval benchmark across short, medium, and long horizons. We vary the number of forecast patches and find that a forecast patch number of 2 achieves the optimal trade-off, resulting in the best overall performance (lowest normalized MASE).
Appendix B Analysis of Inference Speed
To provide a comprehensive view of our model’s efficiency, we benchmarked the single-batch inference speed of Kairos against several state-of-the-art models. The experimental setup involved an input sequence length of 2048 and a prediction horizon of 96 on a single NVIDIA TITAN RTX GPU. For TTM-Advanced, its maximum supported input length of 1536 was used.
As shown in Table 3, the inference time of Kairos is on the same order of magnitude as other highly efficient models like ChronosBolt and Moirai. This demonstrates that Kairos maintains competitive computational efficiency while delivering the superior forecasting accuracy detailed in the main paper. Notably, it is significantly faster than models like Time-MoE, TimesFM, and especially Chronos.
Table 3: Comparison of single-batch inference speeds. All models were tested with an input length of 2048 and an output length of 96, with the exception of TTM-Advanced (*), which used its maximum input length of 1536. Times are averaged per batch and reported in seconds.
Model Inference Time (s)
TTM-Advanced* 0.012
Timer-XL 0.021
ChronosBolt-Base 0.062
Kairos (Ours) 0.065
Moirai-Large 0.077
Time-MoE-Large 0.156
TimesFM 0.210
Chronos-Large 4.901
Appendix C Implementation Details
C.1 Training Details
C.1.1 Model Configurations
We train Kairos in three sizes: mini (10M), small (23M) and base (50M) with the detailed model configurations are in Table 4. The base model are trained for 300,000 steps with batch sizes of 512. We employ the AdamW optimizer, a linear decay learning rate adjustment strategy for model optimization. The learning rate for parameters related to IARoPE is set to 1e-5, while the learning rate for others is set to 1e-3. Training is conducted on 4 × NVIDIA A100 GPUs using TF32 precision, which takes only 15 hours for base size.
Table 4: Details of Kairos model configurations.
Layers Heads $d_{\text{model}}$ $d_{\text{ff}}$ $d_{\text{expert}}$ $P$ $K$ $S$ $Z$ $\tau$ $\eta_{b}$ $w$ $\theta_{\text{init}}$ Params
$\textsc{Kairos}_{\text{mini}}$ 4 4 256 1024 1408 $\{32,64,128\}$ 3 3 2 $\{0.55,0.1,0.05,0.15,0.15\}$ 0.01 128 $10000^{-2j/64}$ 10M
$\textsc{Kairos}_{\text{small}}$ 4 8 384 1536 1408 $\{32,64,128\}$ 3 3 2 $\{0.55,0.1,0.05,0.15,0.15\}$ 0.01 128 $10000^{-2j/64}$ 23M
$\textsc{Kairos}_{\text{base}}$ 6 8 512 2048 1408 $\{32,64,128\}$ 3 3 2 $\{0.55,0.1,0.05,0.15,0.15\}$ 0.01 128 $10000^{-2j/64}$ 50M
C.1.2 Loss Function
To better accommodate varied forecast horizons, and following the methodology of ElasTST (Zhang et al., 2024), we build upon the standard quantile loss by assigning distinct weights to each timestep, such that earlier predictions are given greater importance. The model’s parameters are optimized by minimizing the quantile loss with weight decay, formulated as:
$\mathcal{L}=\frac{1}{B}\sum_{i=1}^{B}\sum_{t=1}^{H}\frac{1}{K}\sum_{k=1}^{K}\omega(t)L_{\alpha_{k}}(y_{i,t},q_{i,t}(\alpha_{k})),$ (9)
$\omega(t)=\frac{1}{H}(\ln(H)-\ln(t)),\quad L_{\alpha}(y,q)=(\alpha-\mathbf{1}_{\{y<q\}})(y-q),$ (10)
where $B,H,K$ are the batch size, forecast horizon, and number of quantiles, respectively; $y_{i,t}$ is the ground truth value, $q_{i,t}(\alpha_{k})$ is the $k$-th quantile forecast, $\omega(t)$ is the weight at each time step $t$, and $L_{\alpha_{k}}$ denotes the pinball loss function. Empirically, we set $K=9$ and use quantile levels of $\alpha=0.1,0.2,\ldots,0.9.$
C.2 Evaluation Details
C.2.1 Metric
GIFT-Eval. GIFT-Eval employs the Mean Absolute Scaled Error (MASE) and the Continuous Ranked Probability Score (CRPS) to assess the performance of point and probabilistic forecasts, respectively. Following the official evaluation protocol, we normalize the metrics for each task using a seasonal naïve baseline and subsequently aggregate the scores across all tasks via the geometric mean.
TSLib. We adopt mean square error (MSE) and mean absolute error (MAE) as evaluation metrics. These metrics are calculated as follows:
$\text{MSE}=\frac{1}{H}\sum_{i=1}^{H}(x_{i}-\hat{x})^{2},\quad\text{MAE}=\frac{1}{H}\sum_{i=1}^{H}\lvert x_{i}-\hat{x}\rvert$ (11)
where $x_{i}$ is the ground truth and $\hat{x}_{i}$ s the prediction for the $i$-th future time point.
C.2.2 Stride
Due to the significant inference latency of our baseline, Chronos (Ansari et al., 2024), we set the evaluation stride to 96 to enhance evaluation efficiency without compromising fairness. This setting is consistent with the protocols of Moirai (Woo et al., 2024), Moirai-MoE (Liu et al., 2024b), and the GIFT-Eval benchmark (Aksu et al., 2024), all of which set the stride equal to the prediction length. Additionally, the original Chronos paper evaluates the model on only a single window per dataset. Therefore, we wish to emphasize a critical point: while the specific numerical results would likely vary with a different stride, our chosen protocol is applied uniformly to all models under evaluation, ensuring a fair and equitable comparison.
C.3 Mixture-of-Size Dynamic Patching (MoS-DP)
C.3.1 Auxiliary-Loss-Free Load Balancing
In this section, we elaborate on the details of bias previously introduced in Section 3.2. In order to ensure that different experts are adequately trained and to control the distribution ratios of various patch sizes, we employ an Auxiliary-Loss-Free Load Balancing method similar to that used in DeepSeek-V3 (Liu et al., 2024a). Specifically, we sum the normalized gating value of the $i$-th expert activated throughout the entire sequence to obtain $L_{i}$:
$L_{i}=\sum_{n=1}^{N}\alpha_{n,i},$ (12)
where
$\alpha_{n,i}=\frac{g_{n,i}}{\sum_{j=1}^{S}g_{n,j}}.$ (13)
We define a target load distribution $\tau=(\tau_{1},\tau_{2},\ldots,\tau_{S})$, where $\tau_{i}$ is the desired proportion of the total load for expert $i$, satisfying $\sum_{j=1}^{S}\tau_{j}=1$. Next, we update the bias term $\bm{b}_{i}$ using $L_{i}$ and $\tau$:
$\bm{b}_{i}\leftarrow\bm{b}_{i}+\eta_{b}\cdot\frac{\tau_{i}\cdot\sum_{j=1}^{S}L_{j}-L_{i}}{\sum_{j=1}^{S}L_{j}},$ (14)
where $\eta_{b}$ is a hyper-parameter to govern the magnitude of this adjustment and is referred to as the bias update speed.
In practical implementation, $\bm{b}_{i}$ is updated once per batch, i.e., $L_{i}$ denotes the sum of the normalized gating values for the $i$-th expert across all sequences within the entire batch. At the end of each step, the bias term $\bm{b}_{i}$ for each expert is adjusted. This dynamic adjustment of $\bm{b}_{i}$ aims to balance the workload across the experts according to the desired distribution by influencing future top-K selections, while also steering the patches-to-expert affinity scores $s^{\prime}_{n,i}$ in subsequent batches, ensuring that the actual load distribution $L_{i}$ progressively aligns with the target distribution $\tau_{i}$, thereby promoting balanced expert utilization over time.
C.3.2 Definition of the ancestor function
In this section, we provide the formal definition for the function $\mathcal{A}(\cdot,\cdot)$ within the Dynamic Patch Fusion (DPF) module, as introduced in Section 3.2.
Consider the $n$-th coarsest patch, which is initially partitioned at the coarsest granularity, $p_{S}$. As this coarsest patch is processed by the Dynamic Patch Router, $K$ experts are activated. Let $p_{k}$ denote the minimum patch size among these activated experts. The coarsest patch is subsequently partitioned into $M_{n}=\frac{p_{S}}{p_{k}}$ non-overlapping finest patches. We denote the $m$-th such finest patch as $\bm{x}_{n,m}^{p_{k}}$.
The function $\mathcal{A}(\bm{x}_{n,m}^{p_{k}},p_{i})$ is then defined as the operation that retrieves the ancestor patch of size $p_{i}$ that contains the finest patch $\bm{x}_{n,m}^{p_{k}}$. Formally, this is expressed as:
$\mathcal{A}(\bm{x}_{n,m}^{p_{k}},p_{i})=[\bm{x}_{\lfloor\frac{(m-1)\cdot p_{k}}{p_{i}}\rfloor\frac{p_{i}}{p_{k}}+1}^{p_{k}},\dots,\bm{x}_{\lfloor\frac{(m-1)\cdot p_{k}}{p_{i}}+1\rfloor\frac{p_{i}}{p_{k}}}^{p_{k}}].$ (15)
We now illustrate the computational process with the following example. Let the setup be as follows:
  • Set of available patch sizes: $\{p_{1},\dots,p_{S}\}=\{32,64,128\}$.
  • Number of null experts: $Z=2$.
  • Number of activated experts: $K=3$.
  • Sequence context length: $T=2048$.
The sequence is partitioned using the maximum patch size $p_{S}=128$, resulting in $N=\lceil\frac{T}{p_{S}}\rceil=16$ coarsest patches. For the first coarsest patch ($n=1$), suppose the activated expert indices are {1, 3, 4}. Since experts 4 and 5 are designated as null experts, the fusion granularities are determined by the non-null activated experts, corresponding to patch sizes $p_{1}=32$ and $p_{3}=128$. The minimum patch size among these is $p_{k}=32$. Consequently, this patch is partitioned into $M_{n}=\frac{p_{S}}{p_{k}}=\frac{128}{32}=4$ non-overlapping finest patches. For the third finest patch ($m=3$), the ancestor patches are computed as follows:
  • For $i=1$ and $p_{i}=p_{1}=32$:
    $\mathcal{A}(\bm{x}_{1,3}^{32},32)=\left[\bm{x}_{\lfloor\frac{(3-1)\cdot 32}{32}\rfloor\frac{32}{32}+1}^{32},\dots,\bm{x}_{\lfloor\frac{(3-1)\cdot 32}{32}+1\rfloor\frac{32}{32}}^{32}\right]=[\bm{x}_{3}^{32}].$ (16)
    The ancestor patch is the finest patch itself.
  • For $i=2$ and $p_{i}=p_{2}=64$:
    $\mathcal{A}(\bm{x}_{1,3}^{32},64)=\left[\bm{x}_{\lfloor\frac{(3-1)\cdot 32}{64}\rfloor\frac{64}{32}+1}^{32},\dots,\bm{x}_{\lfloor\frac{(3-1)\cdot 32}{64}+1\rfloor\frac{64}{32}}^{32}\right]=[\bm{x}_{3}^{32},\bm{x}_{4}^{32}].$ (17)
    Expert 2 is not activated, so its normalized gating value, $\alpha_{n,i}$, in Equation 3 would be zero, meaning it does not contribute to the information fusion.
  • For $i=3$ and $p_{i}=p_{3}=128$:
    $\mathcal{A}(\bm{x}_{1,3}^{32},128)=\left[\bm{x}_{\lfloor\frac{(3-1)\cdot 32}{128}\rfloor\frac{128}{32}+1}^{32},\dots,\bm{x}_{\lfloor\frac{(3-1)\cdot 32}{128}+1\rfloor\frac{128}{32}}^{32}\right]=[\bm{x}_{1}^{32},\bm{x}_{2}^{32},\bm{x}_{3}^{32},\bm{x}_{4}^{32}].$ (18)
This mechanism, enabled by the function $\mathcal{A}(\cdot,\cdot)$, allows for the dynamic selection of different granularities for information fusion, tailored to the information density of each coarsest patch. Consequently, it effectively addresses the challenge of time series heterogeneity inherent in TSFMs.
C.4 Instance-Adaptive Rotary Position Embedding (IARoPE)
In this section, we provide a more detailed description of the IARoPE implementation discussed in Section 3.3. We first introduce the principles of Rotary Position Embedding (RoPE) (Su et al., 2024) in detail, and then provide corresponding supplementary information on the IARoPE implementation.
C.4.1 Details of RoPE
The core idea of RoPE is to rotate segments of the query and key vectors by an angle proportional according to their absolute position in the sequence. This allows the model to discern relative positions through the geometry of these rotations, without needing explicit calculation of relative distances.
Specifically, RoPE operates on vectors of an even dimension, denoted as $D_{h}$. For a vector $\bm{e}$ (representing a query $\mathbf{q}$ or a key $\mathbf{k}$ in self-attention) at position $m$, it is transformed by a rotation matrix $\mathbf{R}_{m}$. This matrix is block-diagonal, composed of $D_{h}/2$ individual $2\times 2$ rotation blocks. Each block $\mathbf{R}_{m,j}$ acts on a pair of dimensions $(e_{2j},e_{2j+1})$ of the vector:
As introduced in Section 3.3, RoPE applies a rotation to input vector. This rotation is applied to pairs of dimensions $(e_{2j},e_{2j+1})$, and can be concisely expressed using complex form:
$f_{\text{RoPE}}(\bm{e},m)=(e_{2j}+ie_{2j+1})e^{im\theta_{j}},$ (19)
where $\theta_{j}$ is the angular frequency. This multiplication by $e^{im\theta_{j}}$ corresponds to a rotation in the complex plane. For the real-valued components $e_{2j}$ and $e_{2j+1}$, this operation is equivalent to applying the following $2\times 2$ rotation matrix $\mathbf{R}_{m,j}$:
$\mathbf{R}_{m,j}=\begin{bmatrix}\cos(m\theta_{j})&-\sin(m\theta_{j})\\ \sin(m\theta_{j})&\cos(m\theta_{j})\end{bmatrix}$, (20)
here, $m$ is the absolute position of the token. The term $\theta_{j}$ represents a predefined angular frequency for the $j$-th pair of dimensions ($j\in[0,D_{h}/2-1]$), typically defined as $\theta_{j}=b^{-2j/D_{h}}$ where $b=10000$ in RoPE (Su et al., 2024) original setting.
The full rotation matrix $\mathbf{R}_{m}$ for position $m$ is thus:
$\mathbf{R}_{m}$ $=\text{diag}(\mathbf{R}_{m,0},\mathbf{R}_{m,1},\dots,\mathbf{R}_{m,D_{h}/2-1})$ (21)
$=\begin{pmatrix}\cos m\theta_{0}&-\sin m\theta_{0}&0&0&\cdots&0&0\\ \sin m\theta_{0}&\cos m\theta_{0}&0&0&\cdots&0&0\\ 0&0&\cos m\theta_{1}&-\sin m\theta_{1}&\cdots&0&0\\ 0&0&\sin m\theta_{1}&\cos m\theta_{1}&\cdots&0&0\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&0&\cdots&\cos m\theta_{D_{h}/2-1}&-\sin m\theta_{D_{h}/2-1}\\ 0&0&0&0&\cdots&\sin m\theta_{D_{h}/2-1}&\cos m\theta_{D_{h}/2-1}\end{pmatrix}$. (22)
Let $\mathbf{q}_{m}$ and $\mathbf{k}_{n}$ be the original query and key vectors for tokens at positions $m$ and $n$ respectively. After applying RoPE, their new representations become $\mathbf{q}^{\prime}_{m}=\mathbf{R}_{m}\mathbf{q}_{m}$ and $\mathbf{k}^{\prime}_{n}=\mathbf{R}_{n}\mathbf{k}_{n}$. The dot product for attention is then:
$(\mathbf{q}^{\prime}_{m})^{\top}(\mathbf{k}^{\prime}_{n})=(\mathbf{R}_{m}\mathbf{q}_{m})^{\top}(\mathbf{R}_{n}\mathbf{k}_{n})=\mathbf{q}_{m}^{\top}\mathbf{R}_{m}^{\top}\mathbf{R}_{n}\mathbf{k}_{n},$ (23)
due to the properties of rotation matrices, $\mathbf{R}_{m}^{\top}\mathbf{R}_{n}=\mathbf{R}_{n-m}$ we can get:
$(\mathbf{q}^{\prime}_{m})^{\top}(\mathbf{k}^{\prime}_{n})=\mathbf{q}_{m}^{\top}\mathbf{R}_{n-m}\mathbf{k}_{n}.$ (24)
This equation shows that the dot product between a query and key vector after rotation inherently depends on their original values ($\mathbf{q}_{m},\mathbf{k}_{n}$) and their relative positions (nmn-m). This allows RoPE to integrate relative positional information into the self-attention without any additional learnable parameters or explicit relative position computations.
C.4.2 Details of IARoPE
The initial RoPE frequencies $\theta_{\text{init},j}$ typically exhibit a wide numerical range. For instance, in our setting $\theta_{\text{init},0}=1.0\ (j=0)$, while $\theta_{\text{init},31}\approx 1.33\times 10^{-4}$ ($j=31$, corresponding to $D_{h}/2-1$ and $D_{h}=64$). Directly applying affine modulation to these values could lead to numerical instability or disproportionate adjustments due to the vast scale differences.
To address this and ensure stable and effective modulation across the entire range of base frequencies, our IARoPE performs the adaptation in log-space. The layer-specific modulation parameters, $\gamma^{(l)}_{j}$ and $\beta^{(l)}_{j}$, predicted by the Multilayer Perceptron (MLP) for layer $l$ following Algorithm 1, are applied to the log-transformed base frequencies $\log\theta_{\text{init},j}$ via an element-wise affine transformation as
$\log\theta^{\prime(l)}_{j}=\gamma^{(l)}_{j}\odot\log\theta_{\text{init},j}+\beta^{(l)}_{j}$, (25)
where $\log\theta^{\prime(l)}_{j}$ represents the modulated log-frequencies for layer $l$ and dimension pair $j$. Then, these modulated log-frequencies are transformed back to their original scale by exponentiation to obtain the adaptive rotation frequencies $\theta^{\prime(l)}_{j}$ used in the IARoPE calculation for layer $l$ as
$\theta^{\prime(l)}_{j}=\exp(\log\theta^{\prime(l)}_{j})$. (26)
Algorithm 1 Generating Instance Adaptive Parameters ($\gamma^{(l)},\beta^{(l)}$)
1:Time series instance $X\in\mathbb{R}^{B\times L}$, mask $M\in\{0,1\}^{B\times L}$, batch size $B$, sequence length $L$, FFT feature dimension $w$, MLP network $f_{\text{MLP}}$
2:Adaptive parameters $\gamma^{(l)}\in\mathbb{R}^{B\times D_{h}/2}$, $\beta^{(l)}\in\mathbb{R}^{B\times D_{h}/2}$
[…]
9:return $\gamma^{(l)},\beta^{(l)}$
Appendix D Calculation of information density
In this section, we present the formula for the information density illustrated in Figure 1(b). To effectively characterize and compare different time series datasets, we introduce a method to quantify their information density and the variation of this density over time. For this, similar to TimeMixer (Wang et al., 2024a), we utilize spectral entropy, a metric that measures the complexity and compressibility of a signal. In this context, a signal with low spectral entropy, such as one with a few dominant periodic components, is considered to have low information density due to its redundant and predictable nature. Conversely, a signal with high spectral entropy, resembling white noise, has its power spread broadly across the frequency spectrum, indicating a high degree of randomness and thus a higher information density.
Given the observations $\mathbf{x}_{1:T}=(x_{1},\ldots,x_{T})\in\mathbb{R}^{T}$ of a specific dataset, we analyze the time series using a sliding window approach. The series is segmented into $N$ windows, $\mathbf{w}_{i}$, each of size $M$ with a step size of $S$. In our analysis, we use a non-overlapping configuration where both the window size and the step size are set to 128 (i.e., $M=128$, $S=128$). The total number of windows is $N=\lfloor(T-M)/S\rfloor+1$. To mitigate spectral leakage from windowing, we apply a Hamming window to each segment $\mathbf{w}_{i}$:
$h[n]=0.5-0.5\cos\left(\frac{2\pi n}{M-1}\right),\quad 0\leq n\leq M-1.$ (27)
For each resulting windowed segment $\mathbf{w}^{\prime}_{i}=\mathbf{w}_{i}\cdot h$, we compute its normalized power spectral density, $p_{i}[k]$, which describes how the signal’s power is distributed over different frequencies. The spectral entropy for the $i$-th window, $H_{SE}(\mathbf{w}_{i})$, is then calculated using the Shannon entropy formula:
$H_{SE}(\mathbf{w}_{i})=-\sum_{k=0}^{M-1}p_{i}[k]\log_{2}(p_{i}[k]).$ (28)
This process yields a sequence of entropy values, $(H_{SE}(\mathbf{w}_{1}),\ldots,H_{SE}(\mathbf{w}_{N}))$, where each value represents the localized information density of its corresponding time segment.
To obtain a holistic view of the dataset’s characteristics, we compute two key statistics from this entropy sequence.
First, the mean spectral entropy ($\mu_{SE}$) serves as a measure of the average information density of the entire dataset. A higher $\mu_{SE}$ suggests that the dataset, on average, contains more complex and less predictable patterns. This allows for a direct comparison of the overall information content between different datasets.
$\mu_{SE}=\frac{1}{N}\sum_{i=1}^{N}H_{SE}(\mathbf{w}_{i}).$ (29)
Second, the standard deviation of spectral entropy ($\sigma_{SE}$) quantifies the variability of information density within the dataset. A small $\sigma_{SE}$ indicates that the dataset is stationary in its complexity, with a consistent level of information density throughout. A large $\sigma_{SE}$, however, reveals a non-stationary character, signifying substantial fluctuations in the signal’s complexity and information content over time.
$\sigma_{SE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(H_{SE}(\mathbf{w}_{i})-\mu_{SE})^{2}}.$ (30)
Together, $\mu_{SE}$ and $\sigma_{SE}$ provide a concise yet powerful summary of a dataset’s informational characteristics, enabling a quantitative assessment of its average complexity and internal dynamics.
Appendix E Additional Details of Experiment Setting
E.1 Pre-training Datasets
We trained Kairos on the Predictability-Stratified Time Series (PreSTS) corpus, which consists of over 300B real-world time series observations from Chronos (Ansari et al., 2024) and Moirai (Woo et al., 2024) in conjunction with 15B synthetic time points. Following (Das et al., 2024), the training loader samples 80% real data and 20% synthetic data.
Real-world data. The real-world datasets were stratified into five tiers based on their predictability. This hierarchical structure dictates the sampling probability during model training, assigning a higher likelihood of selection to datasets with greater predictability. Such a strategy ensures that the model is preferentially trained on high-quality data while preserving its capacity to predict corner cases. Specifically, Tier 1 comprises datasets characterized by pronounced periodicity and trends with low noise. Tier 2 contains datasets with similarly distinct patterns but high noise, whereas Tier 3 includes those with subtle trends and considerable noise. The remaining datasets were classified into Tiers 4 and 5, based on a composite assessment of their size and pattern regularity. The specific details of these datasets, categorized by their respective sampling frequencies, are presented in Tables 5-8.
Synthetic data. We build a synthetic data generator that produces two distinct types of time series, each with a length of 4096. The first type consists of composite series, created by the additive combination of seasonal, trend, and noise components. For seasonality, we sample one or two components, where the primary period is drawn from {24, 48, 288, 360}, and a potential second period is a fixed seven-fold multiple of the first; the seasonal patterns manifest as either spike trains or smooth non-sinusoidal templates generated via interpolation. An optional trend is chosen from linear, exponential, or an ARIMA-like process derived from cumulating a stationary ARMA model’s output. High-probability white Gaussian noise is also added, with each series guaranteed to contain at least one seasonal or trend component and all magnitudes bounded for stability. The generation process for these composite series is formally detailed in Algorithm 2. Complementing these are idealized industrial signals, which simulate perfectly regular machine cycles. These feature a constant baseline with repeating events like trapezoidal spikes or inverted-U shaped dips, where the period, amplitude, and width of the events remain fixed across the entire series with no random jitter, as outlined in Algorithm 3.
Table 5: Detailed descriptions of second-level, minute-level, and hourly datasets.
Dataset Domain Frequency # Time Series # Time points
Wind Power Energy 4S 1 7,397,147
Residential Load Power Energy T 813 437,983,677
Residential PV Power Energy T 699 376,016,850
Loop Seattle Transport 5T 323 33,953,760
Los-Loop Transport 5T 207 7,094,304
PEMS03 Transport 5T 358 9,382,464
PEMS04 Transport 5T 921 15,649,632
PEMS07 Transport 5T 883 24,921,792
PEMS08 Transport 5T 510 9,106,560
PEMS Bay Transport 5T 325 16,941,600
Alibaba Cluster Trace 2018 CloudOps 5T 116,818 190,385,060
Azure VM Traces 2017 CloudOps 5T 159,472 885,522,908
Borg Cluster Data 2011 CloudOps 5T 286,772 1,075,105,708
LargeST Transport 5T 42,333 4,452,510,528
KDD Cup 2022 Energy 10T 134 4,727,519
HZMetro Transport 15T 160 380,320
Q-Traffic Transport 15T 45,148 264,386,688
SHMetro Transport 15T 576 5,073,984
SZ-Taxi Transport 15T 156 464,256
M DENSE Transport 15T 30 525,600
Electricity (15 Min.) Energy 15T 370 41,936,458
Beijing Subway Transport 30T 552 867,744
Elecdemand Energy 30T 1 17,520
Australian Electricity Demand Energy 30T 5 1,155,264
London Smart Meters Energy 30T 5,560 166,528,896
Taxi Transport 30T 2428 3,589,798
Solar Energy H 5,166 543,049,920
BDG-2 Bear Energy H 91 1,482,312
BDG-2 Fox Energy H 135 2,324,568
BDG-2 Panther Energy H 105 919,800
BDG-2 Rat Energy H 280 4,728,288
Borealis Energy H 15 83,269
BDG-2 Bull Energy H 41 719,304
China Air Quality Nature H 2,622 34,435,404
BDG-2 Cockatoo Energy H 1 17,544
Covid19 Energy Energy H 1 31,912
ELF Energy H 1 21,792
GEF12 Energy H 20 788,280
GEF14 Energy H 1 17,520
GEF17 Energy H 8 140,352
BDG-2 Hog Energy H 24 421,056
IDEAL Energy H 217 1,255,253
KDD Cup 2018 Energy H 270 2,897,004
Low Carbon London Energy H 713 9,543,553
Oikolab Weather Climate H 8 800,456
PDB Energy H 1 17,520
Sceaux Energy H 1 34,223
SMART Energy H 5 95,709
Spanish Energy and Weather Energy H 1 35,064
ERCOT Load Energy H 8 1,238,976
M4 Hourly Econ/Fin H 414 373,372
Mexico City Bikes Transport H 494 38,687,004
Electricity (Hourly) Energy H 321 8,443,584
Beijing Air Quality Nature H 132 4,628,448
KDD Cup 2018 Nature H 270 2,942,364
Pedestrian Counts Transport H 66 3,132,346
Rideshare Transport H 2,304 859,392
Traffic Transport H 862 15,122,928
Solar (Hourly) Energy H 5,166 45,254,160
Taxi (Hourly) Transport H 2,428 1,794,292
Uber TLC (Hourly) Transport H 262 1,138,128
Wind Farms (Hourly) Energy H 337 2,869,414
Weatherbench (Hourly) Nature H 225,280 78,992,150,528
Buildings900K Energy H 1,795,256 15,728,237,816
ERA5 Climate H 11,059,200 96,613,171,200
CMIP6 Climate 6H 14,327,808 104,592,998,400
Table 6: Detailed descriptions of daily datasets.
Dataset Domain Frequency # Time Series # Time points
Bitcoin Econ/Fin D 18 81,918
Covid Mobility Transport D 362 148,602
Extended Web Traffic Web D 145,063 370,926,091
Favorita Sales Sales D 111,840 139,179,538
Favorita Transactions Sales D 54 84,408
Hierarchical Sales Sales D 118 212,164
Restaurant Sales D 809 289,402
Subseasonal Climate D 3,448 56,788,560
Subseasonal Precipitation Climate D 862 9,760,426
Sunspot Nature D 1 73,894
Vehicle Trips Transport D 329 32,512
Wiki-Rolling Web D 47,675 40,619,100
Dominick Retail D 100,014 29,652,492
M4 Daily Econ/Fin D 4,227 10,023,836
M5 Sales D 30,490 47,649,940
COVID Deaths Healthcare D 266 56,392
Temperature Rain Nature D 32,072 22,290,040
Monash Weather Climate D 3,010 43,032,000
NN5 Daily Econ/Fin D 111 87,801
Uber TLC Daily Transport D 262 47,422
Weatherbench (Daily) Nature D 225,280 3,291,336,704
Wiki Daily (100k) Web D 100,000 274,100,000
Wind Farms (Daily) Energy D 337 119,549
Exchange Rate Finance D 8 60,704
Table 7: Detailed descriptions of weekly datasets.
Dataset Domain Frequency # Time Series # Time points
CDC Fluview ILINet Healthcare W 375 319,515
CDC Fluview WHO NREVSS Healthcare W 296 167,040
Kaggle Web Traffic Weekly Web W 145,063 16,537,182
Project Tycho Healthcare W 1,258 1,377,707
Traffic Weekly Transport W 862 82,752
M4 Weekly Econ/Fin W 359 371,579
Electricity (Weekly) Energy W 321 50,076
NN5 Weekly Econ/Fin W 111 12,543
Weatherbench (Weekly) Nature W 225,280 470,159,360
Table 8: Detailed descriptions of monthly, quarterly, and yearly datasets.
Dataset Domain Frequency # Time Series # Time points
GoDaddy Econ/Fin M 6,270 257,070
M4 Monthly Econ/Fin M 48,000 11,246,411
Car Parts Sales M 2,674 136,374
CIF 2016 Econ/Fin M 72 7,108
FRED MD Econ/Fin M 107 77,896
Hospital Healthcare M 767 64,428
M1 Monthly Econ/Fin M 617 55,998
M3 Monthly Econ/Fin M 1,428 167,562
Tourism Monthly Econ/Fin M 366 109,280
M3 Other Econ/Fin Q 174 11,933
M4 Quarterly Econ/Fin Q 24,000 2,406,108
M1 Quarterly Econ/Fin Q 203 9,944
M3 Quarterly Econ/Fin Q 756 37,004
Tourism Quarterly Econ/Fin Q 427 42,544
M4 Yearly Econ/Fin Y 23,000 858,458
M1 Yearly Econ/Fin Y 181 4,515
M3 Yearly Econ/Fin Y 645 18,319
Tourism Yearly Econ/Fin Y 518 12,757
Algorithm 2 Composite Time Series Generation
1:Time series length $L=4096$, Primary period set $\mathcal{P}_{1}=\{24,48,288,360\}$, Harmonic multiplier $n=7$, Trend types $\mathcal{T}_{\text{types}}=\{\text{linear, exp, ARMA}\}$, Seasonal patterns $\mathcal{S}_{\text{patterns}}=\{\text{spike, interpolated segment}\}$.
2:A synthetic time series $\bm{x}_{1:L}$.
[…]
4:Sample flags $f_{s},f_{t},f_{n}$ $\triangleright$ Determine component inclusion, ensuring $f_{s}\lor f_{t}$ is true
5:if $f_{s}$ is true then
6:  $k\sim\text{Bernoulli}(0.2)$ $\triangleright$ $k=1$ for double period (20% prob), $k=0$ for single
7:  $p_{1}\sim\mathcal{U}(\mathcal{P}_{1})$ $\triangleright$ Sample primary period
8:  $\mathcal{P}_{\text{active}}\leftarrow\{p_{1}\}$
9:  if $k=1$ then
10:   $\mathcal{P}_{\text{active}}\leftarrow\mathcal{P}_{\text{active}}\cup\{n\cdot p_{1}\}$
11:  end if
12:  for $p$ in $\mathcal{P}_{\text{active}}$ do
13:   $a\sim\mathcal{U}(1.0,3.0)$ $\triangleright$ Sample amplitude for this component
14:   $pattern\sim\mathcal{U}(\mathcal{S}_{\text{patterns}})$
15:   $\bm{c}\leftarrow\text{GeneratePattern}(pattern,p,a)$ $\triangleright$ Create a single cycle of the pattern
16:   $\bm{s}\leftarrow\bm{s}+\text{Tile}(\bm{c},L)$ $\triangleright$ Tile the cycle to length $L$ and add to seasonal component
17:  end for
18:end if
19:if $f_{t}$ is true then
20:  $type\sim\mathcal{U}(\mathcal{T}_{\text{types}})$
21:  $\bm{t}\leftarrow\text{GenerateTrend}(type,L)$
22:  if $f_{s}$ is true then
23:   $\lambda\sim\mathcal{U}(0.1,0.3)$ $\triangleright$ Reduce trend strength when seasonality is present
24:   $\bm{t}\leftarrow\lambda\cdot\bm{t}$
25:  end if
26:end if
27:if $f_{n}$ is true then
28:  $\sigma_{n}\sim\mathcal{U}(0.01,0.1)$
29:  $\bm{n}\sim\mathcal{N}(\bm{0},\sigma_{n}^{2}\mathbf{I})$ $\triangleright$ Generate white Gaussian noise
30:end if
[…]
32:return $\bm{x}_{1:L}$
Algorithm 3 Idealized Industrial Signal Generation
1:Time series length $L=4096$, Pattern types $\mathcal{P}_{\text{types}}=\{\text{inverted\_u, spikes}\}$.
2:A synthetic time series $\bm{x}_{1:L}$.
[…]
7:if $type=\text{inverted\_u}$ then
8:  $sign\leftarrow-1$
9:else
10:  $sign\leftarrow+1$
11:end if
12:for $i\leftarrow 0,p,2p,\dots$ up to L1L-1 do
13:  $start\leftarrow i$, $end\leftarrow\min(i+w,L)$
14:  $\bm{x}_{start:end}\leftarrow\bm{x}_{start:end}+sign\cdot\bm{e}_{1:end-start}$ $\triangleright$ Add or subtract event shape at periodic intervals
15:end for
16:if $\sigma_{n}>0$ then
17:  $\bm{x}\leftarrow\bm{x}+\mathcal{N}(\bm{0},\sigma_{n}^{2}\mathbf{I})$ $\triangleright$ Add global Gaussian noise
18:end if
19:return $\bm{x}_{1:L}$
E.2 Evaluation Datasets
We select datasets from diverse domains and with varying sampling frequencies as evaluation datasets. The details are summarized in Table 9.
Table 9: Detailed descriptions of evaluation datasets.
Dataset Domain Frequency # Time Series # Target # Time points
ETTh1 Energy H 1 7 17,420
ETTh2 Energy H 1 7 17,420
ETTm1 Energy 15T 1 7 69,680
ETTm2 Energy 15T 1 7 69,680
Weather Nature 10T 1 21 52,696
Saugeen (D) Nature D 1 1 23,741
Saugeen (W) Nature W 1 1 3,391
For the evaluation on the TSLib benchmark, the datasets were split into training, validation, and test sets. The split for the ETT and Weather datasets follows the configuration adopted by iTransformer (Liu et al., 2023). All other datasets use a 70%/10%/20% ratio for the training, validation, and test sets, respectively.
E.3 Evaluation Length Selection
In this section, we provide further details regarding the selection of historical sequence lengths for the various models evaluated in the TSLib benchmark, supplementing the discussion in Section 4.1.1.
To accommodate diverse application scenarios, an increasing number of TSFMs (Liu et al., 2025; Das et al., 2024; Liu et al., 2024c; Woo et al., 2024) have devoted attention to predicting over long contexts. Consequently, we evaluate Kairos and other TSFMs under a long‑context setting. Specifically, we adopt a context length of 2048 time steps and examine four prediction horizons, which are {96,192,336,720}. For TSFMs incapable of processing this context length, we instead employ the context length at which each model achieves its best performance.
For the full-shot deep learning baselines, a hyperparameter search was conducted independently for each dataset to ensure a fair evaluation. We compared the performance of each model using the context length from its original publication against a length of 2048, selecting the superior of the two. To be specific, lengths of 96 and 2048 were compared for DLinear (Zeng et al., 2022), iTransformer (Liu et al., 2023), TimesNet (Wu et al., 2022), and Pathformer, while lengths of 336, 512, and 2048 were compared for PatchTST (Nie et al., 2023).
The definitive context lengths adopted for each model are systematically tabulated in Table 10.
Table 10: Context Lengths for Models on the TSLib Benchmark.
Method DLinear iTrans. TimesNet PatchTST Path. Chronos Moirai TimesFM-2.0 Timer-XL $\text{TTM}_{a}$ ChronosBolt Kairos
Context length {96, 2048} {96, 2048} {96, 2048} {336, 512, 2048} {96, 2048} 512 2048 2048 2048 1536 2048 2048
E.4 Details of IARoPE Analysis
In this section, we explain in more detail the setting of the IARoPE analysis experiment discussed in Section 4.4.2. To more definitively ascertain whether these instance-specific modulations truly capture and leverage beneficial positional information derived from an instance’s FFT features, we designed a series of “shuffle” experiments. These experiments function as a form of causal intervention analysis. The underlying hypothesis is: if the instance-derived $\theta$ modulations are crucial for IARoPE’s performance, then disrupting this linkage by applying $\theta$ modulations from one instance to another (shuffling) should lead to a noticeable degradation in forecasting accuracy.
The details of each group in the experiment are as follows:
  • IARoPE: Our proposed method, where each time series instance utilizes its own FFT-derived features to independently modulate its RoPE $\theta$ parameters at each layer.
  • Intra-Dataset Shuffle: During inference, the learned layer-specific $\theta$ modulation parameters ($\gamma^{(l)},\beta^{(l)}$) are randomly permuted among instances within the same batch and originating from the same dataset. The modulated $\theta$ are always shuffled in the same layer.
  • Inter-Dataset Shuffle: For instances of a target dataset, layer-specific $\theta$ modulation parameters ($\gamma^{(l)},\beta^{(l)}$) are randomly sampled from a pre-computed collection derived from a different source dataset and applied to instances of the target dataset. This process also ensures layer-wise correspondence of the applied modulations.
  • Fixed RoPE: We use the fully trained IARoPE model but disable its instance-adaptive modulation during inference. Specifically, the prediction of $\gamma^{(l)}$ and $\beta^{(l)}$ is bypassed, and all instances revert to using the initial RoPE $\theta_{\text{init}}$ across all layers. This approach keeps the base model architecture and other learned weights identical to IARoPE, isolating the effect of the modulation.
By comparing IARoPE’s performance against these shuffled and fixed $\theta$ configurations, we can better attribute performance gains or losses to the effectiveness of the instance-adaptive $\theta$ modulation process.
Appendix F Full Evaluation Results
In this section, we present the detailed results in Section 4.2. Table 11 presents the full results of the zero-shot forecasting experiments conducted at each forecast horizon. For Time-MoE, we report the results as presented in the original paper (Shi et al., 2024).
Table 11: Full results of zero-shot forecasting experiments.
Method Metric ETTh1 ETTh2 ETTm1 ETTm2 Weather Saugeen(D) Saugeen(W) AVG Total $1^{\textbf{st}}$
96 192 336 720 AVG 96 192 336 720 AVG 96 192 336 720 AVG 96 192 336 720 AVG 96 192 336 720 AVG 96 192 336 720 AVG 96 192 336 AVG
DLinear MSE 0.377 0.416 0.446 0.482 0.430 0.28 0.398 0.505 0.796 0.495 0.312 0.339 0.364 0.401 0.354 0.174 0.228 0.287 0.396 0.271 0.174 0.214 0.256 0.309 0.238 0.829 0.832 0.826 0.858 0.836 0.820 0.849 0.781 0.817 0.480 -
MAE 0.391 0.416 0.44 0.495 0.436 0.358 0.434 0.495 0.634 0.480 0.359 0.375 0.392 0.415 0.385 0.272 0.312 0.355 0.428 0.342 0.238 0.272 0.304 0.346 0.290 0.502 0.503 0.506 0.513 0.506 0.569 0.571 0.551 0.564 0.424
iTransformer MSE 0.383 0.441 0.478 0.47 0.443 0.282 0.367 0.402 0.406 0.364 0.327 0.363 0.390 0.424 0.376 0.336 0.369 0.398 0.428 0.383 0.171 0.218 0.278 0.357 0.256 0.778 0.795 0.800 0.839 0.803 0.988 1.041 1.007 1.012 0.501 -
MAE 0.396 0.429 0.449 0.469 0.436 0.344 0.392 0.422 0.433 0.398 0.355 0.381 0.411 0.442 0.397 0.355 0.381 0.411 0.442 0.397 0.21 0.256 0.299 0.351 0.279 0.457 0.463 0.482 0.490 0.473 0.626 0.64 0.642 0.636 0.423
TimesNet MSE 0.408 0.456 0.508 0.567 0.485 0.315 0.435 0.481 0.455 0.422 0.358 0.428 0.494 0.550 0.458 0.175 0.251 0.307 0.412 0.286 0.174 0.229 0.282 0.360 0.261 1.177 1.268 1.129 1.136 1.178 0.817 0.777 0.751 0.782 0.544 -
MAE 0.426 0.451 0.496 0.521 0.469 0.359 0.422 0.456 0.461 0.425 0.373 0.412 0.447 0.480 0.428 0.255 0.305 0.343 0.407 0.328 0.222 0.265 0.304 0.352 0.286 0.546 0.561 0.534 0.531 0.543 0.517 0.494 0.501 0.504 0.423
PatchTST MSE 0.382 0.416 0.443 0.466 0.427 0.280 0.351 0.388 0.424 0.361 0.278 0.327 0.363 0.414 0.346 0.162 0.221 0.270 0.369 0.256 0.149 0.195 0.254 0.346 0.236 0.827 0.83 0.819 0.848 0.831 0.849 0.847 0.781 0.826 0.456 -
MAE 0.403 0.426 0.444 0.474 0.437 0.349 0.393 0.420 0.445 0.402 0.330 0.361 0.388 0.423 0.376 0.249 0.291 0.323 0.386 0.312 0.205 0.248 0.293 0.354 0.275 0.474 0.482 0.488 0.491 0.484 0.545 0.526 0.500 0.524 0.397
Pathformer MSE 0.405 0.466 0.506 0.494 0.468 0.276 0.365 0.420 0.429 0.373 0.302 0.356 0.395 0.463 0.379 0.162 0.225 0.290 0.414 0.273 0.152 0.200 0.258 0.341 0.238 1.168 1.266 1.142 1.149 1.181 0.932 0.949 0.877 0.919 0.533 -
MAE 0.398 0.427 0.447 0.463 0.434 0.334 0.386 0.428 0.443 0.398 0.325 0.361 0.388 0.429 0.376 0.244 0.288 0.333 0.403 0.317 0.192 0.237 0.281 0.335 0.261 0.501 0.535 0.508 0.499 0.511 0.532 0.532 0.533 0.532 0.399
$\text{Chronos}_{s}$ MSE 0.516 0.599 0.613 0.603 0.583 0.304 0.407 0.427 0.450 0.397 0.504 0.601 0.667 0.728 0.625 0.207 0.277 0.343 0.468 0.324 0.209 0.267 0.328 0.397 0.300 1.108 1.142 1.105 1.143 1.125 1.183 1.227 1.165 1.192 0.629 0
MAE 0.416 0.464 0.487 0.512 0.470 0.347 0.401 0.428 0.457 0.408 0.398 0.457 0.498 0.538 0.473 0.277 0.323 0.364 0.437 0.350 0.213 0.265 0.307 0.356 0.285 0.477 0.503 0.506 0.512 0.500 0.520 0.524 0.526 0.523 0.426
$\text{Chronos}_{b}$ MSE 0.526 0.585 0.593 0.608 0.578 0.322 0.411 0.440 0.443 0.404 0.422 0.523 0.600 0.686 0.558 0.206 0.272 0.329 0.436 0.311 0.184 0.241 0.296 0.375 0.274 1.172 1.224 1.202 1.249 1.212 1.166 1.222 1.157 1.182 0.626 0
MAE 0.413 0.447 0.461 0.491 0.453 0.343 0.392 0.422 0.440 0.399 0.374 0.434 0.477 0.520 0.451 0.270 0.316 0.354 0.420 0.340 0.206 0.256 0.298 0.353 0.278 0.473 0.505 0.502 0.505 0.496 0.515 0.517 0.508 0.513 0.415
$\text{Chronos}_{l}$ MSE 0.493 0.601 0.636 0.653 0.596 0.328 0.434 0.480 0.522 0.442 0.452 0.536 0.613 0.671 0.568 0.203 0.268 0.316 0.424 0.303 0.195 0.254 0.311 0.395 0.289 1.178 1.229 1.194 1.232 1.208 1.213 1.269 1.241 1.241 0.642 0
MAE 0.400 0.449 0.475 0.503 0.457 0.348 0.407 0.448 0.482 0.421 0.379 0.430 0.470 0.512 0.448 0.263 0.309 0.342 0.408 0.331 0.211 0.264 0.305 0.360 0.285 0.481 0.508 0.499 0.499 0.497 0.523 0.534 0.544 0.534 0.420
$\text{Moirai}_{s}$ MSE 0.395 0.430 0.445 0.430 0.425 0.274 0.341 0.364 0.378 0.339 0.362 0.385 0.410 0.450 0.402 0.192 0.258 0.307 0.391 0.287 0.160 0.207 0.258 0.327 0.238 1.082 1.112 1.085 1.105 1.096 1.272 1.284 1.177 1.244 0.551 4
MAE 0.404 0.427 0.436 0.445 0.428 0.324 0.369 0.386 0.405 0.371 0.360 0.379 0.399 0.422 0.390 0.263 0.308 0.340 0.397 0.327 0.194 0.238 0.275 0.324 0.258 0.479 0.493 0.491 0.493 0.489 0.560 0.562 0.535 0.552 0.397
$\text{Moirai}_{b}$ MSE 0.429 0.494 0.531 0.544 0.500 0.275 0.347 0.379 0.388 0.347 0.456 0.459 0.482 0.518 0.479 0.194 0.251 0.292 0.354 0.273 0.170 0.218 0.270 0.344 0.251 1.129 1.175 1.142 1.177 1.156 1.208 1.243 1.171 1.207 0.579 1
MAE 0.412 0.447 0.467 0.501 0.457 0.319 0.365 0.388 0.408 0.370 0.385 0.400 0.415 0.435 0.409 0.266 0.306 0.333 0.375 0.320 0.203 0.249 0.288 0.339 0.270 0.472 0.495 0.485 0.485 0.484 0.557 0.550 0.538 0.548 0.403
$\text{Moirai}_{l}$ MSE 0.464 0.523 0.546 0.549 0.521 0.298 0.403 0.464 0.530 0.424 0.522 0.541 0.563 0.611 0.559 0.195 0.257 0.308 0.395 0.289 0.170 0.222 0.277 0.361 0.258 1.102 1.140 1.102 1.122 1.117 1.213 1.252 1.179 1.215 0.604 0
MAE 0.433 0.470 0.486 0.511 0.475 0.346 0.412 0.448 0.498 0.426 0.401 0.423 0.442 0.465 0.433 0.267 0.311 0.342 0.395 0.329 0.203 0.253 0.293 0.347 0.274 0.464 0.486 0.484 0.486 0.480 0.541 0.543 0.534 0.539 0.418
TimesFM-2.0 MSE 0.424 0.444 0.443 0.438 0.437 0.256 0.328 0.366 0.391 0.335 0.362 0.410 0.429 0.470 0.418 0.175 0.239 0.294 0.369 0.269 - - - - - 1.099 1.106 1.058 1.088 1.088 1.277 1.448 3.971 2.232 - 0
MAE 0.397 0.417 0.424 0.442 0.420 0.308 0.361 0.389 0.413 0.368 0.340 0.373 0.395 0.424 0.383 0.245 0.294 0.332 0.385 0.314 - - - - - 0.449 0.483 0.492 0.522 0.487 0.649 0.724 1.362 0.912 -
Timer-XL MSE 0.394 0.436 0.469 0.436 0.434 0.273 0.342 0.377 0.382 0.344 0.325 0.374 0.414 0.510 0.406 0.187 0.240 0.285 0.374 0.272 0.167 0.213 0.266 0.340 0.247 1.043 1.126 1.127 1.127 1.106 1.112 1.128 1.043 1.094 0.537 0
MAE 0.395 0.419 0.437 0.448 0.425 0.336 0.380 0.402 0.416 0.384 0.354 0.385 0.409 0.455 0.401 0.272 0.310 0.338 0.392 0.328 0.221 0.264 0.304 0.354 0.286 0.574 0.612 0.615 0.606 0.602 0.626 0.620 0.594 0.613 0.427
$\text{Time-MoE}_{b}$ MSE 0.357 0.384 0.411 0.449 0.400 0.305 0.351 0.391 0.419 0.367 0.338 0.353 0.381 0.504 0.394 0.201 0.258 0.324 0.488 0.318 0.160 0.210 0.274 0.418 0.266 - - - - - - - - - - 3
MAE 0.381 0.404 0.434 0.477 0.424 0.359 0.386 0.418 0.454 0.404 0.368 0.388 0.413 0.493 0.416 0.291 0.334 0.373 0.464 0.366 0.214 0.260 0.309 0.405 0.297 - - - - - - - - - -
$\text{Time-MoE}_{l}$ MSE 0.350 0.388 0.411 0.427 0.394 0.302 0.364 0.417 0.537 0.405 0.309 0.346 0.373 0.475 0.376 0.197 0.250 0.337 0.480 0.316 0.159 0.215 0.291 0.415 0.270 - - - - - - - - - - 2
MAE 0.382 0.412 0.430 0.455 0.420 0.354 0.385 0.425 0.496 0.415 0.357 0.381 0.408 0.477 0.406 0.286 0.322 0.375 0.461 0.361 0.213 0.266 0.322 0.400 0.300 - - - - - - - - - -
$\text{TTM}_{a}$ MSE 0.373 0.396 0.406 0.411 0.397 0.262 0.326 0.354 0.376 0.330 0.329 0.364 0.381 0.413 0.372 0.167 0.227 0.277 0.360 0.258 0.146 0.193 0.249 0.328 0.229 0.929 0.926 0.904 0.939 0.925 1.114 1.134 1.051 1.156 0.494 11
MAE 0.397 0.415 0.424 0.441 0.419 0.334 0.378 0.400 0.421 0.383 0.350 0.374 0.390 0.410 0.381 0.254 0.296 0.329 0.382 0.315 0.197 0.242 0.284 0.336 0.265 0.511 0.518 0.523 0.541 0.523 0.628 0.628 0.608 0.621 0.408
$\text{ChronosBolt}_{s}$ MSE 0.405 0.475 0.514 0.516 0.478 0.255 0.339 0.388 0.400 0.346 0.309 0.375 0.420 0.512 0.404 0.166 0.231 0.289 0.400 0.272 0.152 0.205 0.264 0.357 0.245 0.863 0.956 1.001 1.103 0.981 0.762 0.802 0.749 0.771 0.489 3
MAE 0.388 0.426 0.453 0.485 0.438 0.307 0.362 0.398 0.418 0.371 0.317 0.357 0.389 0.441 0.376 0.237 0.284 0.324 0.393 0.310 0.186 0.237 0.279 0.338 0.260 0.410 0.458 0.492 0.551 0.478 0.433 0.446 0.473 0.451 0.381
$\text{ChronosBolt}_{b}$ MSE 0.420 0.486 0.517 0.493 0.479 0.255 0.333 0.381 0.393 0.341 0.303 0.370 0.410 0.497 0.395 0.164 0.233 0.299 0.414 0.278 0.150 0.197 0.255 0.344 0.237 0.905 0.959 0.957 1.049 0.968 0.752 0.791 0.725 0.756 0.483 14
MAE 0.388 0.424 0.445 0.457 0.429 0.305 0.355 0.389 0.408 0.364 0.311 0.352 0.382 0.428 0.368 0.232 0.280 0.323 0.391 0.307 0.183 0.230 0.272 0.329 0.254 0.417 0.439 0.441 0.467 0.441 0.423 0.437 0.462 0.441 0.369
$\textsc{Kairos}_{m}$ MSE 0.386 0.436 0.460 0.426 0.427 0.257 0.332 0.369 0.382 0.335 0.304 0.358 0.383 0.408 0.363 0.168 0.228 0.279 0.366 0.260 0.144 0.193 0.254 0.353 0.236 0.976 0.997 0.982 1.015 0.993 0.789 0.840 0.788 0.806 0.477 4
MAE 0.384 0.411 0.422 0.425 0.411 0.314 0.364 0.391 0.407 0.369 0.329 0.362 0.384 0.407 0.371 0.246 0.290 0.323 0.378 0.309 0.181 0.229 0.273 0.331 0.254 0.424 0.439 0.443 0.450 0.439 0.437 0.444 0.426 0.436 0.367
$\textsc{Kairos}_{s}$ MSE 0.385 0.429 0.446 0.422 0.421 0.269 0.340 0.380 0.394 0.346 0.305 0.356 0.381 0.404 0.362 0.163 0.226 0.278 0.382 0.262 0.151 0.197 0.253 0.340 0.235 0.920 0.943 0.942 1.002 0.952 0.808 0.851 0.815 0.825 0.473 4
MAE 0.383 0.408 0.418 0.424 0.408 0.320 0.368 0.397 0.416 0.375 0.326 0.360 0.382 0.404 0.368 0.238 0.284 0.320 0.384 0.307 0.186 0.232 0.274 0.329 0.255 0.410 0.426 0.429 0.442 0.427 0.447 0.446 0.429 0.441 0.366
$\textsc{Kairos}_{b}$ MSE 0.396 0.434 0.452 0.426 0.427 0.276 0.353 0.385 0.386 0.350 0.295 0.336 0.367 0.393 0.348 0.160 0.220 0.271 0.357 0.252 0.146 0.192 0.248 0.338 0.231 0.927 0.955 0.945 0.989 0.954 0.817 0.843 0.804 0.821 0.471 28
MAE 0.385 0.408 0.418 0.427 0.410 0.319 0.369 0.394 0.412 0.374 0.322 0.355 0.380 0.402 0.365 0.238 0.283 0.317 0.373 0.303 0.182 0.228 0.270 0.330 0.253 0.414 0.425 0.427 0.432 0.425 0.446 0.449 0.434 0.443 0.364
Appendix G Showcases
We present several prediction examples generated by Kairos during testing, as illustrated in Figure 7.
(a) ETTh1-1
(b) ETTh1-2
(c) ETTh2-1
(d) ETTh2-2
(e) ETTm1-1
(f) ETTm1-2
(g) ETTm2-1
(h) ETTm2-2
(i) Weather-1
(j) Weather-2
(k) Saugeen-1
(l) Saugeen-2
Figure 7: Example of forecasts from $\textsc{Kairos}_{b}$ on the test datasets used in experiments.
Appendix H Broader Impacts
Our work on Kairos is foundational research focused on advancing time series modeling. While Kairos itself is not designed for direct societal applications with immediate negative impacts, any powerful predictive technology could be misused. So beyond general risks of advanced AI, our model has no specific negative societal impacts need to be discussed here.

### v2

@@
-Kairos: Towards Adaptive and Generalizable Time Series Foundation Models
+
 Abstract
-Time series foundation models (TSFMs) have emerged as a powerful paradigm for time series analysis, driven by large-scale pretraining on diverse data corpora.
-However, time series inherently exhibit heterogeneous information density over time, influenced by system states and signal complexity, presenting significant modeling challenges especially in a zero-shot scenario.
-Current TSFMs rely on non-adaptive processing pipelines that fail to capture this dynamic nature.
-For example, common tokenization strategies such as fixed-size patching enforce rigid observational granularity, limiting their ability to adapt to varying information densities.
-Similarly, conventional positional encodings impose a uniform temporal scale, making it difficult to model diverse periodicities and trends across series.
-To overcome these limitations, we propose Kairos, a flexible TSFM framework that integrates a dynamic patching tokenizer and an instance-adaptive positional embedding.
-Kairos adaptively selects tokenization granularity and tailors positional encodings to the unique characteristics of each time series instance.
-Trained on a large-scale Predictability-Stratified Time Series (PreSTS) corpus comprising over 300 billion time points and adopting a multi-patch prediction strategy in the inference stage, Kairos achieves superior performance with much fewer parameters on two common zero-shot benchmarks, GIFT-Eval and the Time-Series-Library benchmark, consistently outperforming established methods across diverse tasks.
+Inherent temporal heterogeneity, such as varying sampling densities and periodic structures, has posed substantial challenges in zero-shot generalization for Time Series Foundation Models (TSFMs).
+Existing TSFMs predominantly rely on massive parameterization to absorb such heterogeneity, as their static tokenization and positional encoding schemes entangle diverse temporal patterns into a fixed representation space, encouraging memorization rather than adaptation.
+To address this limitation, we propose Kairos, a flexible and parameter-efficient TSFM that decouples temporal heterogeneity from model capacity through a novel tokenization perspective.
+Kairos introduces a dynamic patching tokenizer and a mixture-of-size encoding that adapt observational granularity to local information density, enabling fine-grained temporal abstraction without increasing model width or depth.
+In addition, we design a multi-granularity positional embedding based on dynamic rotary encodings, which conditions on instance-level spectral features and temporal structure induced by dynamic patching tokenization, allowing robust modeling of diverse temporal dependencies.
+Trained on a novel Predictability-Stratified Time-Series (PreSTS) corpus, Kairos achieves superior zero-shot performance with substantially fewer parameters on two mainstream benchmarks, GIFT-Eval and Time-Series-Library.
 The project page is at https://foundation-model-research.github.io/Kairos.
+Figure 1: (a) Comparison on the GIFT-Eval benchmark (Aksu et al., 2024) demonstrates that our Kairos achieves superior zero-shot forecasting performance (lower normalized MASE) while requiring significantly fewer parameters than existing TSFMs.
+(b) (c) Significant variation exists in information density (i.e., signal complexity) across and within different time series datasets.
+(d) Existing TSFMs primarily use tokenization methods like point-wise or fixed-size patching, while our Kairos uses Mixture-of-Size Tokenization to dynamically adapt to information density, enabling the learning of generalizable rules rather than memorization.
 1 Introduction
-Time series forecasting is a core component of many real-world applications, including cloud services, traffic management, retail, finance, and energy (Box et al., 2015; Hyndman & Athanasopoulos, 2018).
+Time series forecasting is a core component of many real-world applications (Box et al., 2015; Wang et al., 2024b).
 Although time series data are ubiquitous, many individual datasets suffer from scarcity, making it difficult to train models independently, particularly for deep neural networks that require large amounts of data.
 This challenge has long motivated the development of few-shot and zero-shot forecasting methods (Oreshkin et al., 2020).
-More recently, inspired by the success of large-scale foundation models (FMs) (Achiam et al., 2023; Kirillov et al., 2023), researchers have turned to building general-purpose time series foundation models (TSFMs).
-These models leverage vast and diverse time series corpora to acquire strong generalization capabilities, showing promising results across a wide range of downstream tasks (Ansari et al., 2024; Woo et al., 2024; Shi et al., 2024).
-Despite this progress, a key limitation persists: current TSFMs rely on non-adaptive processing pipelines that fail to reflect the dynamic and heterogeneous nature of time series.
-As shown in Figure 1(b) and exemplified in Figure 1(c), the statistical information density measured via spectral entropy on sliding windows (see Appendix D) varies not only across datasets but also across samples within the same series.
-However, existing models typically impose a uniform and invariant scale across all series, thereby overlooking internal variability.
-This holds true whether they tokenize each time point individually (Ansari et al., 2024; Shi et al., 2024), tokenize using fixed-size patches (Das et al., 2024), or model temporal dependencies using standard positional encodings (Liu et al., 2025).
-This non-adaptive paradigm gives rise to two major challenges.
-(i) Local information extraction: Current tokenization strategies enforce fixed granularity, requiring the model to observe whole series with the same temporal resolution.
-Such rigidity prevents adaptation to heterogeneous or fluctuating information density, both across different series and within a single sequence.
-For instance, models cannot zoom in on fine-grained details during sudden market shocks, nor can they abstract more efficiently when the data remain stable.
-(ii) Temporal relationship modeling: Conventional positional encodings impose a unified temporal scale across sequences, neglecting differences in periodicity, trend and seasonality.
-As a result, TSFMs struggle to adapt to heterogeneous domains, for example, power consumption data measured hourly versus retail sales data measured daily, each exhibiting distinct temporal dynamics.
-Figure 1(b) and 1(d) illustrate these challenges, highlighting both the variability of information density within time series and the rigidity of current methods.
-Figure 1: (a) The trade-off between performance (normalized MASE) and the number of parameters on GIFT-Eval benchmark (Aksu et al., 2024) for existing TSFMs.
-Our Kairos achieves a superior performance at a comparable parameter scale.
-(b) (c) Significant variation exists in information density across and within different time series datasets.
-(d) Existing TSFMs primarily use tokenization methods like point-wise or fixed-size patching, while our Kairos utilizes a Mixture of Patch Tokenization to address dynamic changes in information density.
-To address these challenges, we propose Kairos, a foundation model designed to handle the complexities of heterogeneous time series data.
-The name Kairos is derived from its mythological significance as the Greek god of the “opportune moment”, reflecting its adaptability.
-For local information extraction, we introduce a Mixture-of-Size Dynamic Patching (MoS-DP) mechanism that adaptively models time series at multiple granularities by selecting patch sizes based on local information density.
-Inspired by the computationally free null experts in Mixture-of-Experts architectures (Zeng et al., 2024; Jin et al., 2024), we further incorporate null patch sizes, which allow the model to flexibly vary the number of active patches.
-This design moves beyond static patching strategies, enabling more expressive modeling of nuanced, time-varying dynamics.
-For temporal relationship modeling, we propose Instance-adaptive Rotary Position Embedding (IARoPE), an extension of RoPE (Su et al., 2024), tailored for diverse temporal dynamics.
-Unlike standard RoPE, which represents positional information with a fixed temporal scale, our IARoPE modulates these scales using spectral features extracted from each series, thereby generating positional encodings uniquely adapted to individual instances.
-In addition to our core architectural innovations, we further introduce two key optimizations in the inference and training phases, respectively.
-For inference, we propose a multi-patch prediction strategy that simultaneously forecasts multiple future patches.
-This approach enhances forecasting robustness by mitigating cumulative errors and offers greater flexibility for variable-length prediction horizons.
+More recently, inspired by the success of large-scale foundation models (Achiam et al., 2023; Kirillov et al., 2023), researchers have turned to building Time Series Foundation Models (TSFMs) for forecasting.
+These models leverage vast and diverse time series corpora to acquire strong generalization capabilities, showing promising results across various downstream tasks (Ansari et al., 2024; Woo et al., 2024; Shi et al., 2024).
+Despite this progress, a key limitation persists: current TSFMs rely on non-adaptive processing pipelines that fail to reflect the dynamic and heterogeneous nature of time series, severely compromising their parameter efficiency.
+As shown in Figure 1(b)(c), the statistical information density (details in Appendix G) varies not only across datasets but also across segments within the same series.
+By applying rigid architectures (exemplified in Figure 1(d)) to this inherently heterogeneous data, models are forced to use massive parameters to memorize specific patterns rather than acquiring adaptive, generalizable rules.
+Unfortunately, this strategy diminishes the overall utility of the model’s capacity, rendering it less effective for real-world deployment.
+This deficiency manifests primarily in two fundamental aspects of model design: (i) Representation learning: TSFMs have yet to adopt dynamic tokenization strategies analogous to those in Large Language Models (LLMs).
+While LLMs utilize subword tokenization (Gage, 1994) to efficiently compress semantic units of varying lengths, TSFMs predominantly persist with static patching, whether through point-wise tokenization (Ansari et al., 2024; Shi et al., 2024) or fixed-size patches (Das et al., 2024).
+By coupling token generation to fixed time intervals rather than information content, these approaches forcibly compress diverse temporal variations into a rigid representation space.
+Consequently, the model wastes computational capacity on low-information segments while struggling to capture high-frequency details within the same fixed budget, leading to significant parameter inefficiency.
+(ii) Temporal dependency modeling: Conventional positional encodings in TSFMs impose a unified temporal scale across sequences, neglecting differences in periodicity, trend and seasonality.
+As a result, existing TSFMs like (Liu et al., 2025) struggle to adapt to heterogeneous domains, for example, power consumption data measured hourly versus retail sales data measured daily, each exhibiting distinct temporal dynamics, as we empirically validate in Section 4.3.
+Consequently, the model requires a deeper, more complex architecture to compensate for this misalignment, further exacerbating parameter inefficiency and hindering practical utility.
+To address these limitations, we propose Kairos, a parameter-efficient model comprising a Mixture-of-Size Encoder, a Heterogeneity-Aware Transformer, and a Multi-Patch Decoder.
+Kairos addresses time series heterogeneity through two primary mechanisms: (i) For intra-patch representation learning, the Mixture-of-Size Encoder adaptively models time series at multiple granularities based on local characteristics.
+Inspired by null experts in Mixture-of-Experts (Zeng et al., 2024; Jin et al., 2024), we introduce null patch sizes to dynamically adjust the number of active granularities.
+This approach moves beyond static tokenization strategies, enabling more expressive modeling of nuanced, time-varying dynamics.
+(ii) For inter-patch dependency modeling, we enhance the Transformer with Dynamic Rotary Position Embedding (DRoPE).
+Unlike standard RoPE (Su et al., 2024), which represents positional information with a fixed temporal scale, DRoPE modulates temporal scales using spectral features and granularity to generate instance-specific positional encodings.
+Flexible intra-patch and inter-patch modeling enable Kairos to effectively characterize heterogeneous time series with a substantially reduced parameter count (see Figure 1(a)).
+Additionally, the Multi-Patch Decoder uses learnable tokens to predict future patches in parallel, which mitigates cumulative errors in autoregressive generation and offers greater flexibility for variable-length prediction horizons.
 To further enhance training, we construct the Predictability-Stratified Time Series (PreSTS) corpus, a large-scale and diverse pretraining dataset curated through a targeted sampling strategy.
@@
 This strong performance is consistently observed on Time-Series-Library (TSLib) (Wang et al., 2024b) benchmarks as well.
-In summary, our main contributions are as follows:
-  • We propose Kairos, a novel foundation model specifically designed to handle dynamically changing time series data.
-Kairos demonstrates strong resilience to commonly observed data variability in real-world time series.
-  • The two key architectural innovations introduced, i.e., MoS-DP and IARoPE, address the challenges of varying information density and instance-specific temporal dynamics, thereby enhancing the generalization capability of foundation models for time series.
-  • Through extensive empirical evaluation, we demonstrate the superior performance of Kairos with significantly fewer parameters on zero-shot forecasting benchmarks, highlighting its flexibility and adaptability across a wide range of diverse and non-stationary time series datasets.
+In summary, our main contributions are as follows: (i) We present Kairos, a parameter-efficient time series foundation model designed to handle the dynamic and heterogeneous nature of time series data.
+(ii) We introduce three key architectural innovations: the Mixture-of-Size Encoder for adaptive multi-granularity tokenization, DRoPE for instance-specific temporal calibration, and the Multi-Patch Decoder for flexible and efficient parallel forecasting.
+(iii) Extensive empirical evaluations on the GIFT-Eval and Time-Series-Library benchmarks confirm the efficacy of Kairos, highlighting its superior zero-shot forecasting performance.
 2 Related Work
 Time Series Foundation Models (TSFMs).
-The scarcity of domain-specific time series data has long motivated research into models with strong few-shot or zero-shot learning capabilities (Oreshkin et al., 2020; 2021).
-Recently, inspired by the strong generalization of large language models, researchers are devoting more effort to TSFMs leveraging vast data corpora.
-Some works attempt to directly transfer the sequence modeling capabilities of large language models to the time series domain (Gruver et al., 2023).
-Others leverage the scalability of Transformer (Vaswani et al., 2017) and directly train unified foundation models on extensive time series corpora (Ansari et al., 2024; Das et al., 2024; Woo et al., 2024; Cohen et al., 2025; Liu et al., 2025).
-A few studies have also explored lightweight architectures, including Multi-Layer Perceptrons (MLPs) such as TTM (Ekambaram et al., 2024).
-However, these models often apply fixed and unified tokenization strategies and position embedding to time series from diverse domains with distinct characteristics, limiting their cross-domain generalization.
-Kairos, in contrast, employs a dynamic patching tokenizer and instance-adaptive position embedding to enhance its generalizability for diverse time series data.
+Inspired by the strong generalization of Large Language Models (LLMs), researchers have devoted increasing attention to TSFMs.
+Existing approaches typically involve adapting LLMs to time series (Gruver et al., 2023), training scalable Transformers on large corpora (Ansari et al., 2024; Das et al., 2024; Woo et al., 2024; Cohen et al., 2025; Liu et al., 2025), or exploring alternative architectures, including non-Transformer models (Auer et al., 2025; Graf et al., 2025) and lightweight Multi-Layer Perceptrons (Ekambaram et al., 2024).
+However, these models often apply fixed tokenization strategies and position embedding to time series from diverse domains with distinct characteristics, limiting their cross-domain generalization.
+Kairos, in contrast, employs a dynamic patching tokenizer and instance-adaptive, granularity-aware position embedding to enhance its generalizability for diverse time series data.
 Time Series Tokenization.
-Leveraging Transformer model backbone requires an effective mechanism for local information extraction, i.e., to transform raw data into sequential elements (tokens), discretely or continuously (Sennrich et al., 2016; Agarwal et al., 2025).
-The intuitive approach for time series tokenization is to map each time point to a token.
-However, models employ this approach (Ansari et al., 2024; Rasul et al., 2023; Shi et al., 2024) suffer from sensitivity to noise and computational inefficiency.
-To address this, Nie et al.
-(2023) introduces a fixed-size patching method that became a widely adopted paradigm in many subsequent TSFMs (Woo et al., 2024; Liu et al., 2025).
-However, the fixed-size patches failed to flexibly adapt to time series with varying information densities.
-While models like Pathformer (Chen et al., 2024) and ElasTST (Zhang et al., 2024) attempt to address these limitations by improving feature fusion strategies, they still simply partition the entire sequence into a few predefined patch sizes, failing to dynamically adjust to the varying information densities in single time series data.
-Furthermore, as they are designed for dataset-specific training rather than as foundation models, they lack cross-domain generalization capabilities.
-In contrast, Kairos improves upon the existing methods by dynamically selecting appropriate patch sizes for each segment of a time series, rather than relying on one or more fixed patch sizes for the entire sequence.
+Transformer-based models require converting raw data into discrete or continuous tokens (Sennrich et al., 2016; Agarwal et al., 2025).
+Current strategies generally fall into two static paradigms: (i) Fixed-size: The dominant paradigm relies on a uniform patch size across the sequence.
+This ranges from fine-grained point-wise tokenization (Ansari et al., 2024; Rasul et al., 2023; Shi et al., 2024) to coarser uni-size patching (Nie et al., 2023), which has been adopted by recent foundation models (Woo et al., 2024; Liu et al., 2025).
+While the latter improves upon the computational inefficiency of point-wise methods, both remain rigid.
+(ii) Multi-size: To capture broader temporal dynamics, models like Pathformer (Chen et al., 2024) and TTM (Ekambaram et al., 2024) incorporate multiple predefined patch sizes.
+However, these scales are applied statically, rather than adapting to the time series instance.
+All the above approaches partition sequences based on fixed time intervals rather than local information.
+This rigidity forces the model to allocate equal parameters to segments of varying complexity.
+This imbalance inevitably leads to suboptimal parameter efficiency and limited generalization.
+In contrast, Kairos employs a mixture-of-size strategy.
+By allocating fewer tokens to simple patterns and more to complex regions, Kairos significantly enhances parameter efficiency while maintaining robust generalization.
 Position Embedding.
-Due to the position-unaware characteristic of self-attention mechanism, Transformers relies on position embeddings (PE) to model temporal information.
-Most TSFMs adopt existing designs for PE in Natural Language Processing (NLP) domains (Das et al., 2024; Shi et al., 2024), which typically emphasize long-term decay through methods like sinusoidal calculations or by suppressing long-range positional information (Vaswani et al., 2017; Su et al., 2024).
+Due to the position-unaware characteristic of self-attention mechanism, Transformers Vaswani et al.
+(2017) rely on position embeddings (PEs) to model temporal information.
+However, most TSFMs adopt existing designs for PE in Natural Language Processing (NLP) domains (Das et al., 2024; Shi et al., 2024), which typically emphasize long-term decay through methods like sinusoidal calculations or by suppressing long-range positional information (Vaswani et al., 2017; Su et al., 2024).
 By imposing a uniform temporal scale, these PEs struggle to model the heterogeneous temporal relationships of time series.
 Existing adaptive PE methods in NLP (Zheng et al., 2024; Lin et al., 2024) are also insufficient, as they target context extrapolation rather than the complex temporal structures of time series.
-In the time series domain, ElasTST (Zhang et al., 2024) proposed a tunable RoPE to better adapt to time series data.
+Therefore, ElasTST (Zhang et al., 2024) proposed a tunable RoPE to better adapt to time series data.
 However, its adaptation uses per-dataset training rather than dynamic adjustment.
-To address this limitation, we propose to dynamically modulate PE tailored to the intrinsic characteristics of each input time series, enhancing the effectiveness of time series foundation models.
-Figure 2: The architecture of KAIROS, which including (1) Mixture-of-Size Dynamic Patching (MoS-DP): This module adaptively tokenizes the time series by fusing features from multiple granularities.
-As detailed in the expanded view, a Dynamic Patch Router first selects active experts (each corresponding to a patch size) for a coarsest patch and determines the finest patch size $p_{k}$ for tokenization.
-The final embedding for each resulting finest patch is then created via a hierarchical fusion process: it is the weighted sum of outputs from all activated experts that correspond to the granularity of the finest patch itself or its ancestors.
-For instance, the diagram illustrates how the embedding for a finest patch of size $p_{1}$ (marked ①) aggregates information from its ancestors of size $p_{2}$.
-This results in a sequence of fused embeddings rich with multi-scale information that is fed into the encoder.
-(2) Instance-Adaptive RoPE (IARoPE): This module (left) adjusts positional encodings for the Transformer by modulating them based on the unique frequency profile of each input series.
+To address this limitation, we propose to dynamically modulate PE tailored to the intrinsic characteristics of each input time series, enhancing the parameter efficiency and overall performance of TSFMs.
+Figure 2: The architecture of Kairos, which includes (i) The Mixture-of-Size Encoder (right) adaptively tokenizes the input series by routing segments to optimal granularity experts and fusing their embeddings into a unified sequence $\boldsymbol{H}$.
+(ii) The Heterogeneity-Aware Transformer (middle) processes these tokens using Dynamic Rotary Position Embedding (DRoPE) (left), which modulates frequencies based on instance-level spectral features and calibrates positions to account for varying patch sizes.
+(iii) The Multi-Patch Decoder (top-middle) employs learnable forecast tokens to predict multiple future patches in parallel via cross-attention.
 3 Methodology
 3.1 Overall Framework
-We present the forecasting problem formulation: given the historical observations of a time series $\bm{x}_{1:T}=(x_{1},\ldots,x_{T})\in\mathbb{R}^{T}$, the objective is to forecast the future values $\bm{x}_{T+1:T+H}=(x_{T+1},\ldots,x_{T+H})\in\mathbb{R}^{H}$, where $T$ is the lookback window and $H$ denotes the forecast horizon.
-The architecture of Kairos, shown in Figure 2, consists of three main components.
-First, the input time series is tokenized into a sequence of embeddings by our Mixture-of-Size Dynamic Patching (MoS-DP) module to extract multi-granularity local information.
-These embeddings are then processed by a Transformer encoder, which uses our proposed Instance-Adaptive Rotary Position Embedding (IARoPE) to model complex temporal relationships.
-Finally, a Transformer decoder utilizes a multi-patch prediction strategy for forecasting.
-The details of each key component, including MoS-DP, IARoPE, and the multi-patch prediction strategy, are presented in the following sections.
-3.2 Mixture-of-Size Dynamic Patching (MoS-DP)
-In contrast to time-series foundation models such as Time-MoE (Shi et al., 2024) and Moirai-MoE (Liu et al., 2024b), which directly replace the original feed-forward network (FFN) layers in Transformer blocks with Mixture-of-Experts (MoE) modules (Jacobs et al., 1991; Shazeer et al., 2017), Kairos takes inspiration from MoE to tackle the intrinsic heterogeneity of time series.
-To improve local information extraction, we introduce the MoS-DP module, which performs routing at the level of each time series patch and thereby effectively addresses the challenge of adapting foundation models to diverse datasets with varying information granularities.
-MoS-DP unfolds in three stages: Patchify at Coarsest Granularity, Dynamic Patch Router, and Dynamic Patch Fusion.
-Patchify at Coarsest Granularity.
-In this stage, the entire time series is partitioned based on the maximum patch size.
-We begin with an initial partition of the sequence.
-A set of $S$ patch sizes is defined in advance, denoted as $\{p_{1},\ldots,p_{S}\}\in\mathbb{R}^{S}$, where $p_{1}<p_{2}<\ldots<p_{S}$.
-We first tokenize the input time series $\bm{x}_{1:T}\in\mathbb{R}^{T}$ using non-overlapping patching using the coarsest patch size $p_{S}$.
-The series is divided into $N=\lceil\frac{T}{p_{S}}\rceil$ coarsest patches.
-If the original series length $T$ is not an exact multiple of $p_{S}$, $\bm{x}_{1:T}$ is augmented with left-padding of zeros as needed.
-This process yields a sequence of $N$ coarsest patches, $\{\bm{x}_{n}^{p_{S}}\}_{n=1}^{N}$, where each patch has a uniform length of $p_{S}$.
-Dynamic Patch Router.
-To perform the final multi-granularity information fusion, we select a set of appropriate granularities for each coarsest patch $\bm{x}_{n}^{p_{S}}$.
-Inspired by the MoE paradigm, we employ multiple FFNs as experts, with each expert dedicated to extracting features at a unique granularity level from the set $\{p_{1},\ldots,p_{S}\}$.
-Furthermore, we introduce $Z$ null experts, which perform no computation when selected, allowing the number of activated experts to be dynamic for each input, facilitating more flexible modeling of heterogeneous time series.
-The routing process is formulated as
-$g_{n,i}=\begin{cases}s^{\prime}_{n,i},&s^{\prime}_{n,i}\in\mathrm{TopK}(\{s^{\prime}_{n,j}|1\leqslant j\leqslant S+Z\},K)\ \text{and}\ 1\leqslant i\leqslant S\\ 0,&\text{otherwise.}\end{cases}$ (1)
-$s^{\prime}_{n,i}=\frac{\exp(s_{n,i}+\bm{b}_{i})}{\sum_{j=1}^{S}\exp(s_{n,j}+\bm{b}_{j})},\quad\text{where }s_{n,i}=\mathbf{W}_{i}\bm{x}_{n}^{p_{S}}.$ (2)
-Here $g_{n,i}$ represents the gating value for the $i$-th expert with respect to the $n$-th coarsest patch; $s_{n,i}$ is the affinity score between patch and the expert, and $\mathbf{W}_{i}$ is the trainable weight matrix for the $i$-th expert.
-The bias term $\bm{b}_{i}$ is utilized for an Auxiliary-Loss-Free Load Balancing method (Liu et al., 2024a) and is discussed in Appendix C.3.1.
-The coarsest patch $\bm{x}_{n}^{p_{S}}$ will then be partitioned into smaller finest patches.
-The size of these finest patches, $p_{k}$ is determined by the finest granularity selected by the router for this specific coarsest patch, i.e., $p_{k}=\min\{p_{i}|g_{n,i}>0\}$.
-Note that this finest granularity can differ for each coarsest patch $\bm{x}_{n}^{p_{S}}$.
-All $N$ coarsest patches are re-segmented in parallel based on their respective routing decisions.
-Compared to uniformly segmenting the entire sequence with a fixed size (Chen et al., 2024; Zhang et al., 2024) as discussed in Section 2, our method offers greater flexibility to adapt to temporal sequences with varying information densities.
-Dynamic Patch Fusion.
-In the final stage, we yield the embedding for each finest patch by fusing information from multiple granularities.
-This process ensures that the final token representations are aware of the context at various temporal scales.
-Specifically, for the $n$-th coarsest patch $\bm{x}_{n}^{p_{S}}$, which was determined in the previous stage, it is first partitioned into $M_{n}=\frac{p_{S}}{p_{k}}$ non-overlapping finest patches.
-Here, $p_{k}$ is the finest patch size selected by the Dynamic Patch Router for this coarsest patch.
-We denote the $m$-th finest patch within the $n$-th coarsest patch as $\bm{x}_{n,m}^{p_{k}}$.
-Each such finest patch corresponds to a single token for the subsequent Transformer encoder.
-The embedding $e_{n,m}$ for the finest patch $\bm{x}_{n,m}^{p_{k}}$ is calculated as the weighted summation of semantic information originating from all activated granularities, formulated as:
-$\bm{e}_{n,m}=\sum_{i=k}^{S}\alpha_{n,i}\cdot\text{MLP}_{i}(\mathcal{A}(x_{n,m}^{p_{k}},p_{i})),\ \text{where}\ \ \alpha_{n,i}=\frac{g_{n,i}}{\sum_{j=k}^{S}g_{n,j}},\ \ \bm{e}_{n,m}\in\mathbb{R}^{D_{h}},$ (3)
-where $\mathcal{A}(x_{n,m}^{p_{k}},p_{i})$ identifies the ancestor patch of size $p_{i}$ that contains the finest patch $x_{n,m}^{p_{k}}$, as illustrated in Figure 2.
-We provide a formal definition of $\mathcal{A}(\cdot,\cdot)$ and a detailed example in Appendix C.3.2.
-$\text{MLP}_{i}$ denotes the expert that extracts the embedding for patch size $p_{i}$.
-The term $\alpha_{n,i}$ represents the normalized gating value from the router for the $n$-th coarsest patch, which dynamically weighs the importance of each granularity.
-$D_{h}$ is the hidden dimension.
-Following Eq.
-3, $M$ embeddings are extracted from the coarsest patch $\bm{x}_{n}^{p_{S}}$ via the Dynamic Patch Fusion (DPF) module as:
-$\text{DPF}(\bm{x}_{n}^{p_{S}},\bm{g}_{n})=[\bm{e}_{n,1},\bm{e}_{n,2},\ldots,\bm{e}_{n,M_{n}}]\in\mathbb{R}^{M_{n}\times D_{h}},\quad\bm{g}_{n}=(g_{n,1},g_{n,2}\ldots,g_{n,S}).$ (4)
-Eventually, the original time series $\bm{x}_{1:T}$ is transformed into a sequence of embeddings by concatenating all DPF module outputs as $\mathbf{E}=\text{DPF}(\bm{x}^{p_{S}}_{1},\bm{g}_{1})||\text{DPF}(\bm{x}^{p_{S}}_{2},\bm{g}_{2})||\ldots||\text{DPF}(\bm{x}^{p_{S}}_{N},\bm{g}_{N})$ with sequence length $T^{\prime}=\sum_{n=1}^{N}M_{n}$, where $M_{n}$ represents the number of finest patches that the $n$-th patch in $\bm{x}^{p_{S}}_{1:N}$ is segmented into by the Dynamic Patch Router.
-This final sequence $\mathbf{E}\in\mathbb{R}^{T^{\prime}\times D_{h}}$ would serve as the input to the subsequent Transformer module of Kairos.
-This design allows the model to use small patches for in-depth analysis where local details are rich and large patches for efficient modeling where information is stable, as demonstrated in Section 4.4.1.
-3.3 Instance-Adaptive Rotary Position Embedding (IARoPE)
-Rotary Position Embedding (RoPE) (Su et al., 2024) encodes relative positions by rotating hidden vectors within Transformer layer, such as a query ($\bm{q}$) or a key ($\bm{k}$), which we denote generically as $\bm{z}$.
-The operation is formulated as $f_{\text{RoPE}}(\bm{z},t)=(\bm{z}_{2j}+i\bm{z}_{2j+1})e^{it\theta_{\text{init},j}}$, where the base RoPE frequencies $\theta_{\text{init},j}=b^{-2j/D_{h}}$ are determined by a fixed hyperparameter $b$.
-A detailed derivation of the RoPE calculation is provided in Appendix C.4.
-However, these fixed frequencies impose a single, uniform temporal scale across all data.
-This rigidity prevents the model from effectively modeling temporal relationships for series with diverse intrinsic structures, such as different seasonalities or trends.
-To address this limitation, we propose Instance-Adaptive Rotary Position Embedding (IARoPE), which dynamically tailors customized positional encodings for each time series instance.
-The core idea of IARoPE is to modulate the base RoPE frequencies $\theta_{\text{init}}$ using spectral features extracted from each instance.
-Specifically, Given an input time series instance $\mathbf{x}_{1:T}$, we first extract its frequency features by applying the Fast Fourier Transform (FFT) and select the first $\omega$ low-frequency components (their magnitudes) to form a feature vector $\mathbf{x}_{\text{fft}}$, where $\omega$ is a hyperparameter.
-We focus on low frequencies as they often capture dominant trends and periodic patterns crucial for understanding the temporal structure of instance (Zhou et al., 2022).
-Then we feed $\mathbf{x}_{\text{fft}}$ into a small MLP to predict instance-specific modulation parameters: a scaling factor $\mathbf{\gamma}\in\mathbb{R}^{D_{h}/2}$ and a bias factor $\mathbf{\beta}\in\mathbb{R}^{D_{h}/2}$.
-These parameters then adapt the base RoPE frequencies $\theta_{\text{init}}\in\mathbb{R}^{D_{h}/2}$ via an element-wise affine transformation to produce the adaptive frequencies $\theta^{\prime}\in\mathbb{R}^{D_{h}/2}$:
-$\theta^{\prime}_{j}=\gamma_{j}\odot\theta_{\text{init},j}+\beta_{j}\quad\text{for }j=0,1,\ldots,D_{h}/2-1,$ (5)
-where $\odot$ denotes the Hadamard product.
-The resulting $\theta^{\prime}$ incorporates learned information from the instance’s frequency characteristics.
-Finally, these adaptive angles $\theta^{\prime}$ replace the fixed $\theta$ in the RoPE calculation for the current instance:
-$f_{\text{IARoPE}}(\bm{z},t)=(\bm{z}_{2j}+i\bm{z}_{2j+1})e^{it\theta^{\prime}_{j}},$ (6)
-This mechanism adapts the positional encoding to the frequency profile of each time series instance, potentially offering a more accurate representation of relative temporal positions.
-We demonstrate the effectiveness of this adaptation in Section 4.4.2.
-3.4 Multi-Patch Prediction
-Kairos employs multi-patch prediction to effectively mitigate the cumulative errors inherent in autoregressive processes.
-Specifically, after the encoder outputs the hidden representation $\bm{h}^{\text{en}}\in\mathbb{R}^{T^{\prime}\times D_{h}}$, a Transformer decoder is employed to forecast future values.
-We introduce a set of $J$ learnable forecast tokens, denoted as $Q_{\text{forecast}}=\{o_{1},\dots,o_{J}\}$.
-These tokens are passed as input to decoder, which processes them simultaneously to produce a sequence of output embeddings $\bm{h}^{\text{de}}\in\mathbb{R}^{J\times D_{h}}$:
-$\bm{h}^{\text{de}}=\text{Decoder}(Q_{\text{forecast}},\bm{h}^{\text{en}}).$ (7)
-Each output embedding $\bm{h}^{\text{de}}_{j}$ is then passed through a shared prediction head (a residual FFN block (Das et al., 2024)) to forecast a patch of the future of length $H_{s}=H/J$ as
-$\hat{\bm{x}}_{T+jH_{s}+1:T+(j+1)H_{s}}=\text{ResidualFFN}(\bm{h}^{\text{de}}_{j}),\quad j=\{1,\dots,J\}$ (8)
-This multi-token design is motivated by the limitations of two common forecasting paradigms.
-First, compared to conventional single-step autoregressive models (Ansari et al., 2024; Liu et al., 2024c), predicting $J$ patches at once enhances computational efficiency and mitigates cumulative error by reducing iterative steps.
-Second, unlike models that forecast a larger patch (Das et al., 2024), our approach provides greater flexibility for shorter prediction horizons.
-It can construct a precise forecast by aggregating outputs from the necessary number of tokens, avoiding the potentially imprecise truncation of a larger-than-needed output.
-We provide a detailed analysis of this design in Appendix A.
+We formulate our task as learning a time series forecasting function $f_{\phi}:\mathbb{R}^{T}\to\mathbb{R}^{H}$ parameterized by $\phi$.
+Given historical time series $\mathbf{X}\in\mathbb{R}^{T}$, time series forecasting is to predict the future horizon $\mathbf{Y}\in\mathbb{R}^{H}$.
+The learning objective is to minimize the discrepancy between the ground truth $\mathbf{Y}$ and the prediction $\hat{\mathbf{Y}}=f_{\phi}(\mathbf{X})$.
+We detail it in Appendix D.1.2.
+The architecture of Kairos, shown in Figure 2, consists of three components.
+First, the input time series is tokenized by the Mixture-of-Size Encoder to extract multi-granularity local representations via dynamic granularity selection.
+These embeddings are then processed by a Heterogeneity-Aware Transformer equipped with Dynamic Rotary Position Embedding (DRoPE), a granularity-aware positional encoding that performs instance-wise modulation and adapts to dynamic patching tokenization.
+Finally, the Multi-Patch Decoder generates forecasts by predicting multiple future patches in parallel, mitigating cumulative errors.
+The details of these components are presented in the following sections.
+3.2 Mixture-of-Size Encoder
+Prior to the Heterogeneity-Aware Transformer (Section 3.3), the Mixture-of-Size Encoder adaptively tokenizes the raw time series $\mathbf{X}$ through a three-stage process.
+First, the series is partitioned into fixed-size segments, where a router dynamically selects optimal temporal granularities for each segment.
+Next, for each activated granularity, the segment is simultaneously patchified and projected into the latent space by the expert network corresponding to that granularity.
+Finally, these multi-granularity embeddings are fused into a unified token sequence representing the input series.
+Notably, the Mixture-of-Size Encoder is both sparse and dynamic.
+A gating mechanism equipped with null experts (described below) adaptively selects relevant granularities for each segment, enabling sparse expert activation and segment-specific granularity adaptation.
+Coarse Segmentation.
+As shown in Figure 1(b) in Section 1 where time series exhibits inherent temporal heterogeneity, we first apply coarse segmentation to the input time series $\mathbf{X}$ and process each segment independently in subsequent tokenization and encoding stages.
+Specifically, $\mathbf{X}$ is partitioned into a sequence of non-overlapping coarse segments $\mathbf{X}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{N}]$, where each segment $\mathbf{x}$ has a fixed length $P$ and the total number of segments is $N=\lceil{T}/{P}\rceil$.
+In the following, we operate within each segment $\mathbf{x}$ and therefore omit the segment index for notational simplicity.
+Granularity.
+For a given time series segment $\mathbf{x}$, we adopt a multi-granularity modeling perspective.
+Granularity refers to the temporal resolution at which the segment is tokenized and encoded, as defined below.
+Definition 3.1 (Granularity Level).
+A granularity level $i\in\{1,\ldots,G\}$ is defined by a specific patch size $P_{i}$ and its corresponding encoding network $\pi_{i}$, under which the time series segment $\mathbf{x}$ is patchified and mapped into a latent representation.
+Intuitively, finer granularity levels capture high-frequency local variations, while coarser granularity levels emphasize low-frequency global trends.
+At each granularity level $i$, the patchification and embedding operations described below are applied independently and in parallel, producing representations that characterize the segment at the corresponding temporal resolution.
+Step 1: Top-$K$ Granularity Routing.
+Applying encoding at an appropriate granularity level for each time series segment is non-trivial due to the inherent temporal heterogeneity of time series data.
+To address this challenge, we introduce a dynamic routing mechanism equipped with a lightweight router that adaptively selects an optimal subset of granularity levels for subsequent encoding of a given segment.
+Inspired by the Mixture-of-Experts (MoE) paradigm (Jacobs et al., 1991; Shazeer et al., 2017), we instantiate $G$ granularity-specific experts, each corresponding to a distinct granularity level, together with $Z\in\{1,\dots,K-1\}$ null experts that perform no computation.
+The inclusion of null experts allows the model to implicitly skip encoding at certain granularities, enabling adaptive control over the effective number of active granularities for each input segment.
+The routing weights are computed via a learnable scoring function.
+Specifically, for each expert $i\in\{1,\dots,G+Z\}$, an affinity score is first computed as $s_{i}=\mathbf{w}_{i}^{\top}\mathbf{x},$ which is then normalized with a bias-corrected softmax $\tilde{s}_{i}=\left(\exp(s_{i}+\boldsymbol{b}_{i})\right)/\left(\sum_{j=1}^{G+Z}\exp(s_{j}+\boldsymbol{b}_{j})\right),$ where $\mathbf{w}_{i}$ denotes learnable parameters and $\boldsymbol{b}_{i}$ is a bias term introduced for auxiliary-loss-free load balancing (Liu et al., 2024a) (details in Appendix D.3.1).
+The router thus produces a set of active granularity indices $\mathcal{I}_{\mathrm{valid}}=\operatorname*{argtop\text{-}K}_{i\in[G+Z]}\penalty 10000\ \tilde{s}_{i}\cap\{1,\dots,G\}.$ That is, only the Top-$K$ granularity experts are activated, while null experts contribute no computation even when selected.
+By selectively activating granularity experts and allowing null experts to absorb routing probability mass, Kairos dynamically adapts the effective granularity resolution for each time series segment according to its local information, while maintaining a sparse and parameter-efficient architecture.
+Step 2: Granularity-wise Encoding.
+For each selectively activated granularity expert at level $i$, we apply the corresponding patching and embedding operations.
+Specifically, a given time series segment $\mathbf{x}$ is re-partitioned into a sequence of patches at granularity $i$:
+$\boldsymbol{p}_{i,1},\dots,\boldsymbol{p}_{i,M_{i}}]=\mathrm{Patchify}(\mathbf{x},i)\in\mathbb{R}^{M_{i}\times P_{i}},$ (1)
+where $P_{i}$ denotes the patch size at granularity $i$, and $M_{i}=\left\lceil{P}/{P_{i}}\right\rceil$ is the resulting number of patches obtained via non-overlapping segmentation, with zero-padding applied when necessary.
+Each patch is then projected into the latent space to obtain an embedding sequence $\boldsymbol{E}_{i}=[\boldsymbol{e}_{i,1},\dots,\boldsymbol{e}_{i,M_{i}}]$:
+$\quad\boldsymbol{e}_{i,m}=\pi_{i}(\boldsymbol{p}_{i,m})\in\mathbb{R}^{D_{h}},m\in[1,M_{i}]\penalty 10000\ .$ (2)
+where $\pi_{i}$ denotes a granularity-specific projection function, implemented as a two-layer multilayer perceptron (MLP).
+Step 3: Mixture-of-Size Fusion.
+In the final encoder stage, we fuse the embeddings produced by the activated granularity experts into a unified token sequence, serving as the input to the Heterogeneity-Aware Transformer.
+First, for a time-series segment $\mathbf{x}$, we align all granularity-specific embeddings $\boldsymbol{E}_{i}$ to the finest activated resolution $P_{\min}=\min\{P_{i}\mid i\in\mathcal{I}_{\mathrm{valid}}\},$ yielding a target sequence length $M=\lceil P/P_{\min}\rceil$.
+This alignment is crucial to eliminate representation bias arising from unequal sequence lengths induced by different patch sizes across granularities.
+To enable efficient alignment, we design the patch sizes to be nested, i.e., $P_{i+1}$ is divisible by $P_{i}$.
+Under this constraint, each embedding sequence $\boldsymbol{E}_{i}\in\mathbb{R}^{M_{i}\times D_{h}}$ can be deterministically expanded to the target resolution via repetition (details and an example in Appendix D.3.2): $\tilde{\boldsymbol{E}}_{i}=\mathrm{Expand}(\boldsymbol{E}_{i})\in\mathbb{R}^{M\times D_{h}}.$ The fused representation is obtained by a gated weighted summation over the activated granularities: $\bar{\boldsymbol{E}}=\sum_{i\in\mathcal{I}_{\mathrm{valid}}}\alpha_{i}\,\tilde{\boldsymbol{E}}_{i},\alpha_{i}=\tilde{s}_{i}/{\sum_{j\in\mathcal{I}_{\mathrm{valid}}}\tilde{s}_{j}}.$
+Sample-level Result. Finally, the representations from all coarse segments are concatenated to form the encoder output
+$\boldsymbol{H}=\mathrm{Concat}(\bar{\boldsymbol{E}}_{1},\ldots,\bar{\boldsymbol{E}}_{N})\in\mathbb{R}^{\tilde{T}\times D_{h}},$ (3)
+where $\tilde{T}=\sum_{n=1}^{N}M_{n}$ and $M_{n}$ is the embedding sequence length of the $n$-th segment, forming the representation sequence for the given time series sample $\mathbf{X}$.
+Unlike prior approaches such as Chen et al.
+(2024); Zhang et al.
+(2024), which impose a static multi-scale partitioning over the entire time series, our method (i) dynamically adapts the tokenization granularity at the segment level.
+(ii) decouples temporal heterogeneity from a fixed representational scale by explicitly and dynamically assigning patch size via routing and multi-granularity encoding, enabling more effective intra-patch representation learning.
+This adaptive behavior cannot be achieved via globally predefined scales applied statically and is empirically validated in Section 4.4.1.
+As a result, varying temporal patterns are modeled via the lightweight Mixture-of-Size Encoder, reducing the burden of subsequent model parameterization.
+3.3 Heterogeneity-Aware Transformer
+Following Mixture-of-Size Encoder, the fused representation $\boldsymbol{H}\in\mathbb{R}^{\tilde{T}\times D_{h}}$ is processed by a Heterogeneity-Aware Transformer, a stack of standard Transformer (Vaswani et al., 2017) encoder blocks, yielding representations $\boldsymbol{H}^{\text{en}}\in\mathbb{R}^{\tilde{T}\times D_{h}}$ that are subsequently fed to the prediction head.
+To inject temporal information into the self-attention mechanism, we propose Dynamic Rotary Position Embedding (DRoPE).
+Unlike static encodings such as Vaswani et al.
+(2017); Su et al.
+(2024), DRoPE adaptively calibrates temporal structure to account for instance-specific periodicities and the varying physical durations of dynamic patches.
+Base Structure of DRoPE.
+Our dynamic position encoding is based on the standard Rotary Position Embedding (RoPE) Su et al.
+(2024), where a hidden vector $\boldsymbol{z}\in\mathbb{R}^{D_{h}}$ (a query or key) at token index $t\in[1,\tilde{T}$] is treated as a set of $D_{h}/2$ complex numbers.
+Each pair is rotated by an angle proportional to the position $t$ and a fixed frequency $\theta_{d}$ as below, where $d\in[0,D_{h}/2-1]$ indexes the dimension:
+$f_{\text{DRoPE}}(\boldsymbol{z},t)=(\boldsymbol{z}_{2d}+i\boldsymbol{z}_{2d+1})e^{it\theta_{d}}$ (4)
+Here the frequencies are traditionally pre-defined as $\theta_{d}=b^{-2d/D_{h}}$.
+DRoPE generalizes this formulation by dynamically transforming both the frequency $\theta$ and the token position $t$.
+A detailed derivation of the RoPE calculation is provided in Appendix D.4.
+Instance-Level Frequency Modulation.
+Standard RoPE uses a uniform frequency $\theta$ for all sequences, failing to capture the unique periodic structures of heterogeneous time series (analysis in Theorem D.1).
+We propose to modulate the base frequencies using instance-specific spectral features.
+Specifically, given a time series input $\mathbf{X}$, we extract the magnitudes of the first $\omega$ low-frequency components via Fast Fourier Transform, denoted as $\mathbf{X}_{\mathrm{fft}}$.
+A lightweight MLP maps these features to a scaling factor $\gamma$ and a bias $\beta$, to produce adaptive frequencies $\theta_{\mathrm{inst}}$:
+$\theta_{\text{inst},d}=\gamma_{d}\cdot\theta_{d}+\beta_{d}\penalty 10000\ ,\text{where\penalty 10000\ }\gamma,\beta=\mathrm{MLP}(\mathbf{X}_{\mathrm{fft}})$ (5)
+This affine transformation allows Kairos to re-scale the temporal rotation based on the dominant information like trends and seasonalities in the specific time series instance.
+Granularity-Aware Position Calibration.
+In our multi-granularity framework, the discrete token index $t$ is no longer an accurate proxy for time because different tokens represent different physical durations, a direct consequence of the granularity routing and tokenization in our Mixture-of-Size Encoder, as described in Section 3.2.
+To maintain temporal consistency across mixed patch sizes, we replace the integer index $t$ with a calibrated position $\tilde{t}=\sum_{k=1}^{t-1}{\bar{P}_{k}}/{\bar{P}_{\mathrm{g\_min}}}$, where $\bar{P}_{k}$ is the patch size of the $k$-th token and $\bar{P}_{\mathrm{g\_min}}=\min\{P_{i}\}_{i=1}^{G}$ is the globally minimum patch size of the finest granularity.
+The resulting DRoPE operation unifies these two components, rotating the hidden vector $\boldsymbol{z}$ at position $t$ as follows:
+$f_{\text{DRoPE}}(\boldsymbol{z},t)=(\boldsymbol{z}_{2d}+i\boldsymbol{z}_{2d+1})e^{i\cdot\tilde{t}\cdot\theta_{\text{inst},d}}$ (6)
+By measuring relative distances in terms of absolute physical time and instance-specific frequencies, DRoPE enables the Transformer to learn consistent temporal dependencies considering the diverse encoding granularity and the intrinsic properties of the data sample.
+3.4 Multi-Patch Decoder
+To mitigate the error accumulation typical of patch-wise autoregressive decoding while maintaining structural flexibility, Kairos utilizes a Multi-Patch Decoder.
+This component forecasts $J$ future patches in parallel, effectively decoupling the prediction horizon from iterative dependencies.
+We first introduce a sequence of $J$ learnable forecast tokens, $\mathbf{Q}_{\text{forecast}}=\{\mathbf{q}_{1},\dots,\mathbf{q}_{J}\}$.
+Multi-Patch Decoder processes these token embeddings with the latent representations $\boldsymbol{H}^{\text{en}}\in\mathbb{R}^{\tilde{T}\times D_{h}}$ from the above Heterogeneity-Aware Transformer through cross-attention, to produce a sequence of latent forecasting embeddings $\boldsymbol{h}^{\text{de}}\in\mathbb{R}^{J\times D_{h}}$.
+Each embedding $\boldsymbol{h}^{\text{de}}_{j}$ is then projected into the observation space via a shared prediction head, which is implemented as a residual FFN block Das et al.
+(2024), to generate a future patch of length $H_{s}=\lceil H/J\rceil$ as
+$\hat{\mathbf{Y}}_{(j-1)H_{s}+1:jH_{s}}=\text{ResidualFFN}(\boldsymbol{h}^{\text{de}}_{j}),\quad j\in[1,J]$ (7)
+This multi-token design is strategically for two objectives regarding predominant forecasting paradigms.
+(i) Error mitigation and efficiency: Compared to conventional single-step autoregressive models (Ansari et al., 2024; Liu et al., 2024c), predicting $J$ patches in parallel improves computational efficiency and mitigates cumulative error by reducing iterative decoding steps.
+(ii) Horizon flexibility: Unlike models that generate a single larger patch (Das et al., 2024), our approach offers superior granularity for variable prediction horizons.
+By assembling the forecast from the specific subset of tokens required, Kairos ensures high-fidelity alignment with the target length and avoids the imprecise truncation associated with oversized outputs.
+We provide a detailed analysis of this design in Appendix B.
+By combining the Mixture-of-Size Encoder for adaptive input processing with this Multi-Patch Decoder for flexible output generation, Kairos achieves a unified and parameter-efficient framework for zero-shot time series forecasting.
 3.5 Training Corpus
-For pre-training, we curated Predictability-Stratified Time Series (PreSTS) corpus, a corpus of over 300 billion time points.
-PreSTS comprises real-world time series spanning multiple domains and is augmented by a synthetic dataset we constructed to ensure comprehensive coverage.
-Additionally, we have excluded all test sets found within the GIFT-Eval benchmark (Aksu et al., 2024) to maintain evaluation integrity.
-While a prevailing consensus regarding TSFMs highlights the necessity of a diverse training dataset, research has also shown that the presence of anomalies in the training set can severely degrade a model’s predictive capabilities (Cheng et al., 2024).
-Therefore, the real-world datasets in PreSTS are stratified into five tiers based on their degree of predictability.
-Datasets exhibiting higher predictability are assigned a greater sampling probability during training, a strategy that enables the model to undergo a more robust, high-quality training regimen without sacrificing the comprehensiveness of the data.
-Comprehensive descriptions of our training datasets along with synthetic data generation algorithm, are provided in Appendix E.1.
+For pre-training, we curated the Predictability-Stratified Time Series (PreSTS) corpus, comprising over 300 billion time points.
+PreSTS integrates large-scale real-world time series from diverse domains and a complementary synthetic dataset to ensure broad coverage.
+All test sets from the GIFT-Eval (Aksu et al., 2024) and TSLib (Wang et al., 2024b) benchmarks are explicitly excluded to prevent data leakage.
+While dataset diversity is widely regarded as essential for TSFM training, recent studies indicate that low-predictability or anomalous sequences can substantially impair forecasting (Cheng et al., 2024).
+To address this, we stratify real-world datasets into five tiers based on predictability and assign higher sampling probabilities to more predictable data during training.
+This strategy promotes stable and high-quality pre-training without compromising data diversity.
+Further details are provided in Appendix E.1.
 4 Experiment
-In this section, we conduct a comprehensive evaluation to assess the performance of Kairos and the effect of proposed components.
-Specifically, we investigate the following three key research questions.
-RQ1: Does Kairos achieve better generalization and zero-shot prediction than previous methods?
-(Section 4.2) RQ2: How effectively does MoS-DP handle time series heterogeneity by adapting its patch size according to local information densities?
-(Section 4.3 and 4.4.1) RQ3: Does IARoPE effectively generate customized position embedding tailored to the unique temporal structure of each time series instance?
+In this section, we assess the performance of Kairos and its components by addressing three research questions.
+RQ1: Does Kairos achieve superior zero-shot generalization capabilities while maintaining parameter efficiency compared to previous methods?
+(Section 4.2) RQ2: Does the Mixture-of-Size Encoder effectively handle time series heterogeneity by adapting its granularity based on local information density?
+(Section 4.3 and 4.4.1) RQ3: Does DRoPE effectively generate customized position embeddings tailored to the unique temporal structure of each time series instance?
 (Section 4.3 and 4.4.2)
+Table 1: Performance evaluation on the GIFT-Eval.
+We evaluate using normalized MASE and CRPS (Section 4.1.1).
+Baselines encompass statistical, Deep Learning (DL), and TSFM methods, with TSFMs categorized into TestData Leakage and Zero-Shot following official protocols.
+TestData Leakage signifies that the model’s training set included test data, whereas Zero-Shot indicates that the model was trained without any data from the test set or its corresponding training split.
+Baseline results are officially reported by GIFT-Eval.
+Type Statistical DL (Full-Shot) TSFMs (TestData Leakage) TSFMs (Zero-Shot)
+Method Seasonal Naïve DLinear PTST. TTM Chronos Chronos Bolt TimesFM Moirai VisionTS Ying. Toto Sundial $\textsc{Kairos}_{s}$ (ours) $\textsc{Kairos}_{b}$ (ours)
+#Params - - - 5M 709M 205M 500M 311M 112M 300M 151M 128M 23M 53M
+MASE $\downarrow$ 1.000 1.061 0.849 1.020 0.870 0.808 0.758 0.875 0.863 0.798 0.750 0.750 0.748 0.738
+CRPS $\downarrow$ 1.000 0.846 0.587 0.873 0.574 0.574 0.550 0.599 0.755 0.548 0.517 0.559 0.554 0.548
+Figure 3: Zero-shot performance on TSLib averaged over prediction lengths {96, 192, 336, 720}.
+The subscripts $l$, $b$, $s$, $m$, and $a$ represent model sizes of large, base, small, mini and advanced, respectively.
+The complete experimental results are presented in Appendix F.
 4.1 Experiment Settings
 4.1.1 Evaluation Datasets and Metric
-To comprehensively evaluate the performance of our proposed method, we conducted zero-shot assessments on two well-known public benchmarks: GIFT-Eval (Aksu et al., 2024) and Time-Series-Library (TSLib) (Wang et al., 2024b).
-We train Kairos in three sizes: mini (10M), small (23M) and base (50M).
-The training details are provided in Appendix C.1.
+To comprehensively evaluate our proposed method, we conducted zero-shot assessments on two well-known public benchmarks: GIFT-Eval (Aksu et al., 2024) and Time-Series-Library (TSLib) (Wang et al., 2024b).
+We train Kairos in three sizes: mini (10M), small (23M) and base (53M).
+The training details are provided in Appendix D.1.
 Benchmarks.
-GIFT-Eval benchmark consists of 97 prediction tasks, with 55 tasks dedicated to short-term predictions, 21 for medium-term predictions, and 21 for long-term predictions, providing a comprehensive means of assessing the model’s predictive capabilities.
-For the TSLib benchmark, we followed the common practice in Time-MoE (Shi et al., 2024) and selected the ETT datasets (Zhou et al., 2021a) and Weather (Zhou et al., 2021b).
-Additionally, to test the model’s adaptability to diverse sampling frequencies, we incorporated the Saugeen datasets (Godahewa et al., 2021), which include data with both daily and weekly sampling frequencies.
-These were referred to as Saugeen (D) and Saugeen (W), respectively.
-The detailed description of the evaluation datasets, their splitting configurations, and the selection of the evaluation length are provided in Appendix E.2 and E.3.
+The GIFT-Eval benchmark comprises 97 tasks spanning short- (55), medium- (21), and long-term (21) horizons, facilitating a comprehensive assessment.
+For TSLib, we selected the ETT and Weather (Zhou et al., 2021) datasets, consistent with Time-MoE (Shi et al., 2024).
+To test adaptability to diverse frequencies, we also included the Saugeen datasets (Godahewa et al., 2021), denoted as Saugeen (D) and Saugeen (W) for daily and weekly intervals, respectively.
+The detailed description of the evaluation datasets, splitting configurations, and evaluation length selection are provided in Appendix E.2 and E.3.
 Metric.
@@
 For the TSLib benchmark, we followed previous works (Liu et al., 2023; 2024c; Nie et al., 2023) by adopting Mean Squared Error (MSE) and Mean Absolute Error (MAE) to evaluate time series forecasting performance.
-The metric calculations are detailed in Appendix C.2.1.
+The metric calculations are detailed in Appendix D.2.1.
 4.1.2 Compared Methods
-We compare Kairos with several state-of-the-art models, including TSFMs such as Sundial (Liu et al., 2025), Toto (Cohen et al., 2025), YingLong (Wang et al., 2025), VisionTS (Chen et al., 2025), Time-MoE (Shi et al., 2024), ChronosBolt (Ansari et al., 2024), TTM (Ekambaram et al., 2024), Moirai (Woo et al., 2024), Timer-XL (Liu et al., 2024c), TimesFM-2.0 (Das et al., 2024), and Chronos (Ansari et al., 2024), as well as advanced full-shot deep learning models, including PatchTST (Nie et al., 2023), DLinear (Zeng et al., 2022), iTransformer (Liu et al., 2023), Pathformer (Chen et al., 2024), and TimesNet (Wu et al., 2022).
-.
+We compare Kairos with several state-of-the-art models, including TSFMs such as Sundial (Liu et al., 2025), Toto (Cohen et al., 2025), YingLong (Wang et al., 2025), VisionTS (Chen et al., 2025), Time-MoE (Shi et al., 2024), ChronosBolt (Ansari et al., 2024), TTM (Ekambaram et al., 2024), Moirai (Woo et al., 2024), Timer-XL (Liu et al., 2024c), TimesFM-2.0 (Das et al., 2024), Chronos (Ansari et al., 2024), and advanced full-shot deep learning models, including PatchTST (Nie et al., 2023), DLinear (Zeng et al., 2022), iTransformer (Liu et al., 2023), Pathformer (Chen et al., 2024), and TimesNet (Wu et al., 2022).
 4.2 Zero-shot Evaluation
-Finding 1: Kairos demonstrates remarkable zero-shot forecasting capabilities across multiple benchmarks.
-On the GIFT-Eval benchmark, as presented in Table 1, $\textsc{Kairos}_{\text{base}}$ (50M) achieves the best MASE and the second-best CRPS when compared to state-of-the-art task-specific deep learning methods and other TSFMs.
+Finding 1: Kairos achieves superior zero-shot forecasting performance with significantly higher parameter efficiency.
+On the GIFT-Eval benchmark, as presented in Table 1, $\textsc{Kairos}_{\text{base}}$ (53M) achieves the best MASE and the second-best CRPS among state-of-the-art task-specific deep learning methods and other TSFMs.
 Notably, $\textsc{Kairos}_{\text{small}}$ (23M) surpasses both Toto and Sundial in MASE, despite having a parameter count that is 6.6 and 5.6 times smaller, respectively.
 On the TSLib benchmark, as shown in Figure 3, our lightweight $\textsc{Kairos}_{\text{mini}}$ (10M) surpasses the performance of both recent advanced TSFMs and the majority of full-shot deep learning models.
-Compared with other TSFMs, Kairos achieves a significant improvement in predictive performance.
-Table 1: Performance evaluation on the GIFT-Eval.
-We evaluated model performance using the normalized MASE and CRPS metrics delineated in Section 4.1.1, where lower values indicate higher forecasting accuracy.
-The baseline models fall into three categories: statistical methods, deep learning (DL) models, and TSFMs.
-Following the official classification of GIFT-Eval, TSFMs are further subdivided into TestData Leakage and Zero-Shot.
-TestData Leakage signifies that the model’s training set included test data, whereas Zero-Shot indicates that the model was trained without any data from the test set or its corresponding training split.
-Baseline results are officially reported by GIFT-Eval.
-Type Statistical DL (Full-Shot) TSFMs (TestData Leakage) TSFMs (Zero-Shot)
-Method
-Seasonal
-Naïve
-DLinear
-PTST. TTM Chronos
-Chronos
-Bolt
-TimesFM Moirai VisionTS Ying. Toto Sundial
-$\textsc{Kairos}_{s}$
-$\textsc{Kairos}_{b}$
-#Params - - - 5M 709M 205M 500M 311M 112M 300M 151M 128M 23M 50M
-MASE 1.000 1.061 0.849 1.020 0.870 0.808 0.758 0.875 0.863 0.798 0.750 0.750 0.748 0.742
-CRPS 1.000 0.846 0.587 0.873 0.574 0.574 0.550 0.599 0.755 0.548 0.517 0.559 0.554 0.548
-Figure 3: Zero-shot forecasting performance on TSLib.
-Results are averaged across prediction lengths {96, 192, 336, 720}.
-The subscripts $l$, $b$, $s$, and $a$ represent model sizes of large, base, small, and advanced, respectively.
-The complete experimental results are presented in Appendix F.
+These results confirm that Kairos delivers a significant improvement in predictive capability without relying on massive scale.
 4.3 Ablation Study
-Table 2: Ablation study comparing MoS-DP and RoPE variants.
-We evaluated ablation settings using the normalized MASE from Section 4.1.1, assessing performance on individual prediction horizons and in aggregate across all tasks.
-MoS-DP RoPE Variant Short Medium Long AVG
-IARoPE 0.719 0.759 0.789 0.742
-IARoPE 0.725 0.779 0.813 0.755
-RoPE 0.726 0.783 0.828 0.759
-RoPE 0.731 0.784 0.829 0.763
-Finding 2: MoS-DP and IARoPE collaboratively enable Kairos to adapt to heterogeneous time series, achieving powerful predictive performance.
-To investigate the effectiveness of our proposed MoS-DP and IARoPE,we conducted a comprehensive ablation study, with the results summarized in Table 2.
-For a fair comparison, Kairos without MoS-DP uses a fixed patch size of 64, which was determined to be optimal via hyperparameter search.
-For IARoPE, we compare against the original RoPE (Su et al., 2024), which employs a fixed, predefined $\theta$.
-The integration of MoS-DP and IARoPE into our method leads to the best performance on the 97 tasks of the GIFT-Eval benchmark, which validates the superiority of our designs.
-Furthermore, the improvement of Kairos gets more significant compared to Kairos without MoS-DP or IARoPE as horizon gets longer, exhibiting better long-term forecasting ability.
+Finding 2: The synergistic integration of Mixture-of-Size Encoding, DRoPE, and the Multi-Patch Decoding is critical for adapting to heterogeneous time series and achieving superior forecasting performance.
+Table 2 details the component-wise contributions.
+(i) Encoder Design: Kairos outperforms variants using fixed patch sizes or excluding null experts, confirming that dynamic, sparse multi-granularity modeling is critical for capturing diverse temporal patterns.
+(ii) Positional Embedding: DRoPE proves superior to standard RoPE (Su et al., 2024).
+Our ablation further isolates the gains from Instance-Level Frequency Modulation and Granularity-Aware Position Calibration, verifying the benefit of adapting to both spectral heterogeneity and physical time distortion caused by dynamic patching.
+(iii) Decoding Strategy: The Multi-Patch Decoding shows clear superiority over standard single-patch autoregressive decoding.
+The integration of both components yielded the best performance across 97 GIFT-Eval tasks, which validates the superiority of our designs.
+Furthermore, the performance gap widens as the forecast horizon increases, demonstrating Kairos’s superior long-term forecasting capability.
+Table 2: Ablation study comparing encoder design, positional embedding and decoder strategy.
+Models were evaluated using the normalized MASE (detailed in Section 4.1.1) across individual prediction horizons and as an aggregate across all tasks.
+Model Variants Short Medium Long AVG
+Impact of Encoder Design
+   w/ Fixed Patch Size ($P=32$) 0.724 0.802 0.820 0.761
+   w/ Mixture-of-Size (w/o Null Experts) 0.720 0.770 0.800 0.748
+Impact of Positional Embedding
+   w/ Standard RoPE (Fixed $\theta$) 0.729 0.807 0.835 0.767
+   w/ Only Instance-Level Frequency Modulation 0.719 0.767 0.797 0.746
+   w/ Only Granularity-Aware Position Calibration 0.727 0.796 0.807 0.758
+Impact of Decoder Strategy
+   w/ Single-Patch Autoregressive 0.705 0.794 0.848 0.753
+[…]
 4.4 Model Analysis
-4.4.1 Patch Distribution Analysis
+4.4.1 Mixture-of-Size Tokenization Analysis
+Figure 4: Adaptive tokenization visualization.
+Tokens are demarcated by dotted lines and varying colors, where deeper red curves indicate finer granularity.
+Kairos adapts granularity to information density, applying finer processing to volatile regions.
 Finding 3: Kairos demonstrates adaptive capability to varying information densities, enabling it to model complex temporal dynamics with appropriate granularity.
-For each coarsest patch, Kairos dynamically fuses information by selecting from a set of granularities (patch sizes).
-To quantify this dynamic routing, we calculate a weighted average patch size by multiplying the weight assigned to each granularity by its corresponding patch size.
-Consequently, a smaller value for this metric indicates the model’s preference for finer-grained processing of a given coarsest patch, while a larger value suggests a preference for coarser granularity.
-In Figure 4, this weighted average patch size is visualized by the background color, with darker shades corresponding to smaller values.
-As the figure illustrates, Kairos strategically employs smaller patch sizes for sequences with abrupt changes and larger patch sizes for stable sequences.
-Figure 4: Patch size preferences in GIFT-Eval test datasets.
-Darker shades indicate a smaller weighted average patch size, signifying the model’s preference for finer-grained processing in that region.
-4.4.2 IARoPE Analysis
-Figure 5: Causal analysis of adaptive modulation by IARoPE.
-This experiment validates the criticality of matching positional encodings to the unique characteristics of each time series instance by disrupting or removing this adaptation.
-We test this by manipulating the RoPE frequencies $\theta$ under several conditions: IARoPE (standard), Intra-Dataset Shuffle ($\theta$ modulations permuted between different instances within the same dataset), Inter-Dataset Shuffle ($\theta$ modulations from instances of other datasets), and Fixed RoPE (no modulation).
+Instead of enforcing a uniform resolution, Kairos dynamically routes segments to optimal granularity based on local information.
+As visualized in Figure 4, Kairos strategically allocates fine-grained tokens to regions with abrupt changes or high volatility, while employing coarse-grained tokens for stable trends.
+This adaptive allocation mechanism reduces computational redundancy on low-information segments, allowing the model to achieve superior performance with a compact parameter budget compared to rigid architectures.
+See Appendix C for a detailed inference speed analysis.
+4.4.2 DRoPE Analysis
+Table 3: Causal analysis of adaptive modulation by DRoPE.
+We evaluate the impact of disrupting or removing the instance-wise $\theta$ adjustment.
+Degradation indicates relative performance drop compared to DRoPE.
+Method DRoPE Intra-Dataset Standard RoPE Inter-Dataset
+(Ours) Shuffle (Fixed $\theta$) Shuffle
+MASE $\downarrow$ 0.738 0.751 0.767 0.947
+Degradation -1.76% -3.93% -28.32%
 Finding 4: Modulating RoPE frequencies $\theta$ instance-wise can indeed better model time series temporal relationships.
-To verify that the performance gains of IARoPE are indeed caused by its instance-specific RoPE frequencies $\theta$ modulation, we designed a series of control experiments that impair or ablate this mechanism by shuffling the modulation parameters (details in Figure 5; Appendix E.4).
-The results confirm IARoPE’s superior forecasting accuracy, demonstrating the advantage of tailoring PE to individual instance frequency characteristics.
-The Intra-Dataset Shuffle ($\theta$ modulations permuted between different instances within the same dataset) performed second best, suggesting that while instance-specific modulation is optimal, leveraging dataset-level frequency similarities still offers utility.
-Conversely, both Inter-Dataset Shuffle ($\theta$ modulations from instances of other datasets) and Fixed RoPE (no modulation) yielded poorer performance, highlighting that mismatched or static $\theta$ parameters are suboptimal as they fail to capture the specific temporal nature of each series.
+To verify that DRoPE’s performance gains stem from its instance-specific $\theta_{\text{inst}}$ modulation, we designed control experiments that disrupt this mechanism (Table 3; Appendix E.4): (i) Intra-Dataset Shuffle: randomly permute $\theta_{\text{inst}}$ across instances within the same dataset; (ii) Inter-Dataset Shuffle: assign $\theta_{\text{inst}}$ from instances of different datasets; (iii) Fixed RoPE: use standard RoPE without any modulation.
+The results reveal a clear performance gradient (DRoPE $\succ$ Intra-Shuffle $\succ$ Fixed RoPE $\succ$ Inter-Shuffle), indicating that DRoPE learns instance-specific representations to modulate $\theta$.
+Intra-Dataset Shuffle outperforms Fixed RoPE, as samples within the same dataset share similar spectral characteristics, allowing shuffled $\theta_{\text{inst}}$ to remain partially compatible.
+However, Intra-Dataset Shuffle still incurs a 1.76% degradation compared to DRoPE, suggesting that instance-level adaptation contributes to the performance gains beyond dataset-level statistics.
+Inter-Dataset Shuffle yields the worst results (28.32% degradation), as $\theta_{\text{inst}}$ from different datasets cannot capture the target series’ temporal dynamics.
+These results collectively validate the effectiveness of instance-level $\theta_{\text{inst}}$ adaptation in DRoPE.
 5 Conclusions
-We introduce Kairos, a time series foundation model that advances the design of modules specifically adapted to the unique characteristics of time series data.
-Kairos achieves this through key innovations, including Mixture-of-Size Dynamic Patching, which considers varying information densities for more precise and efficient modeling, and Instance-Adaptive RoPE, which provides more accurate positional information tailored to time series nuances.
-Our comprehensive experiments reveal superior performance achieved by Kairos, highlighting the importance of designing foundational models that are inherently aligned with the structural nuances of time series data.
-Appendix A Analysis of Multi-Patch Prediction
+We introduce Kairos, a parameter-efficient foundation model explicitly tailored for time series heterogeneity.
+Its architecture features three key innovations: a Mixture-of-Size Encoder for adaptive granularity, a Heterogeneity-Aware Transformer with Dynamic RoPE for instance-specific temporal calibration, and a Multi-Patch Decoder for efficient parallel forecasting.
+Extensive experiments demonstrate that Kairos achieves superior performance, validating the necessity of aligning architectural inductive biases with the structural nuances of time series data.
+One limitation of our current approach is that Kairos relies on channel-independent modeling (Nie et al., 2023) to support multivariate time series and does not explicitly capture inter-variable dependencies.
+Future work aims to model these interactions and extend Kairos to tasks beyond forecasting.
+Impact Statement
+Our work on Kairos is foundational research focused on advancing parameter-efficient and adaptive time series modeling.
+We anticipate several positive societal impacts: (i) promoting sustainable AI by significantly reducing the computational resources and energy required for large-scale deployment; (ii) enhancing decision-making in dynamic environments (e.g., energy, retail) through improved modeling of heterogeneous temporal patterns; and (iii) providing a high-quality data curation methodology via the PreSTS corpus to improve model performance.
+While we do not foresee immediate negative societal consequences, we recognize that advanced predictive technologies carry inherent risks of misuse.
+Consequently, rigorous ethical oversight is therefore essential to guarantee that such predictive capabilities serve the public interest.
+Appendix A Etymology of Kairos
+Kairos was chosen for its mythological significance as the Greek god of the “right or critical moment”, reflecting our model’s ability to select the ideal set of granularities and positional encoding for time series characterized by varying information densities and frequencies.
+Furthermore, this continues a thematic tradition seen in other foundational models we benchmark against, such as Chronos (Ansari et al., 2024) (the personification of time) and Moirai (Woo et al., 2024) (the personifications of destiny).
+Appendix B Analysis of Multi-Patch Prediction
+Figure 5: Performance analysis of multi-patch prediction on the GIFT-Eval benchmark across short, medium, and long horizons.
+We vary the number of forecast patches and find that a forecast patch number of 2 achieves the optimal trade-off, resulting in the best overall performance (lowest normalized MASE).
+Unlike task-specific, non-autoregressive models such as TFT (Lim et al., 2021) and Informer (Zhou et al., 2021), which predict all future time points simultaneously, their architecture confines them to forecasting a fixed length identical to the training sequence, precluding variable-length predictions.
+Conversely, TSFMs typically employ an iterative process: they predict a segment of the sequence, append it to the historical data, and then forecast the subsequent segment, thereby enabling autoregressive predictions of arbitrary length.
+To mitigate the cumulative errors inherent in this autoregressive methodology, Kairos introduces forecast tokens designed to predict multiple patches within each autoregressive step.
 In this section, we present an analysis of the multi-patch prediction strategy introduced in Section 3.4.
@@
 Ultimately, by evaluating the mean normalized MASE, we identified the optimal trade-off, selecting a forecast patch number of $J=2$ for Kairos.
-Figure 6: Performance analysis of multi-patch prediction on the GIFT-Eval benchmark across short, medium, and long horizons.
-We vary the number of forecast patches and find that a forecast patch number of 2 achieves the optimal trade-off, resulting in the best overall performance (lowest normalized MASE).
-Appendix B Analysis of Inference Speed
+Appendix C Analysis of Inference Speed
 To provide a comprehensive view of our model’s efficiency, we benchmarked the single-batch inference speed of Kairos against several state-of-the-art models.
@@
 Model Inference Time (s)
-TTM-Advanced* 0.012
-Timer-XL 0.021
-ChronosBolt-Base 0.062
-Kairos (Ours) 0.065
-Moirai-Large 0.077
-Time-MoE-Large 0.156
-TimesFM 0.210
+TTM-Advanced* 0.009
+Timer-XL 0.014
+ChronosBolt-Base 0.055
+Kairos-Base (Ours) 0.061
+Moirai-Large 0.070
+Sundial-Base 0.099
+Time-MoE-Large 0.148
+TimesFM-2.0 0.202
 Chronos-Large 4.901
-Appendix C Implementation Details
-C.1 Training Details
-C.1.1 Model Configurations
-We train Kairos in three sizes: mini (10M), small (23M) and base (50M) with the detailed model configurations are in Table 4.
+Toto-Base 13.717
+Appendix D Implementation Details
+D.1 Training Details
+D.1.1 Model Configurations
+We train Kairos in three sizes: mini (10M), small (23M) and base (53M) with the detailed model configurations are in Table 5.
 The base model are trained for 300,000 steps with batch sizes of 512.
 We employ the AdamW optimizer, a linear decay learning rate adjustment strategy for model optimization.
-The learning rate for parameters related to IARoPE is set to 1e-5, while the learning rate for others is set to 1e-3.
+The learning rate for parameters related to DRoPE is set to 1e-5, while the learning rate for others is set to 1e-3.
 Training is conducted on 4 × NVIDIA A100 GPUs using TF32 precision, which takes only 15 hours for base size.
 Table 5: Details of Kairos model configurations.
-Layers Heads $d_{\text{model}}$ $d_{\text{ff}}$ $d_{\text{expert}}$ $P$ $K$ $S$ $Z$ $\tau$ $\eta_{b}$ $w$ $\theta_{\text{init}}$ Params
+Layers Heads $d_{\text{model}}$ $d_{\text{ff}}$ $d_{\text{expert}}$ $P$ $K$ $G$ $Z$ $\tau$ $\eta_{b}$ $w$ $\theta$ Params
 $\textsc{Kairos}_{\text{mini}}$ 4 4 256 1024 1408 $\{32,64,128\}$ 3 3 2 $\{0.55,0.1,0.05,0.15,0.15\}$ 0.01 128 $10000^{-2j/64}$ 10M
 $\textsc{Kairos}_{\text{small}}$ 4 8 384 1536 1408 $\{32,64,128\}$ 3 3 2 $\{0.55,0.1,0.05,0.15,0.15\}$ 0.01 128 $10000^{-2j/64}$ 23M
-$\textsc{Kairos}_{\text{base}}$ 6 8 512 2048 1408 $\{32,64,128\}$ 3 3 2 $\{0.55,0.1,0.05,0.15,0.15\}$ 0.01 128 $10000^{-2j/64}$ 50M
-C.1.2 Loss Function
+$\textsc{Kairos}_{\text{base}}$ 6 8 512 2048 1408 $\{32,64,128,256\}$ 4 4 2 $\{0.55,0.1,0.05,0.15,0.15\}$ 0.01 128 $10000^{-2j/64}$ 53M
+D.1.2 Loss Function
 To better accommodate varied forecast horizons, and following the methodology of ElasTST (Zhang et al., 2024), we build upon the standard quantile loss by assigning distinct weights to each timestep, such that earlier predictions are given greater importance.
@@
 Empirically, we set $K=9$ and use quantile levels of $\alpha=0.1,0.2,\ldots,0.9.$
-C.2 Evaluation Details
-C.2.1 Metric
+D.2 Evaluation Details
+D.2.1 Metric
 GIFT-Eval.
@@
 where $x_{i}$ is the ground truth and $\hat{x}_{i}$ s the prediction for the $i$-th future time point.
-C.2.2 Stride
+D.2.2 Stride
 Due to the significant inference latency of our baseline, Chronos (Ansari et al., 2024), we set the evaluation stride to 96 to enhance evaluation efficiency without compromising fairness.
@@
 Therefore, we wish to emphasize a critical point: while the specific numerical results would likely vary with a different stride, our chosen protocol is applied uniformly to all models under evaluation, ensuring a fair and equitable comparison.
-C.3 Mixture-of-Size Dynamic Patching (MoS-DP)
-C.3.1 Auxiliary-Loss-Free Load Balancing
+D.3 Mixture-of-Size Encoder
+D.3.1 Auxiliary-Loss-Free Load Balancing
 In this section, we elaborate on the details of bias previously introduced in Section 3.2.
@@
 We define a target load distribution $\tau=(\tau_{1},\tau_{2},\ldots,\tau_{S})$, where $\tau_{i}$ is the desired proportion of the total load for expert $i$, satisfying $\sum_{j=1}^{S}\tau_{j}=1$.
-Next, we update the bias term $\bm{b}_{i}$ using $L_{i}$ and $\tau$:
-$\bm{b}_{i}\leftarrow\bm{b}_{i}+\eta_{b}\cdot\frac{\tau_{i}\cdot\sum_{j=1}^{S}L_{j}-L_{i}}{\sum_{j=1}^{S}L_{j}},$ (14)
+Next, we update the bias term $\boldsymbol{b}_{i}$ using $L_{i}$ and $\tau$:
+$\boldsymbol{b}_{i}\leftarrow\boldsymbol{b}_{i}+\eta_{b}\cdot\frac{\tau_{i}\cdot\sum_{j=1}^{S}L_{j}-L_{i}}{\sum_{j=1}^{S}L_{j}},$ (13)
 where $\eta_{b}$ is a hyper-parameter to govern the magnitude of this adjustment and is referred to as the bias update speed.
-In practical implementation, $\bm{b}_{i}$ is updated once per batch, i.e., $L_{i}$ denotes the sum of the normalized gating values for the $i$-th expert across all sequences within the entire batch.
-At the end of each step, the bias term $\bm{b}_{i}$ for each expert is adjusted.
-This dynamic adjustment of $\bm{b}_{i}$ aims to balance the workload across the experts according to the desired distribution by influencing future top-K selections, while also steering the patches-to-expert affinity scores $s^{\prime}_{n,i}$ in subsequent batches, ensuring that the actual load distribution $L_{i}$ progressively aligns with the target distribution $\tau_{i}$, thereby promoting balanced expert utilization over time.
-C.3.2 Definition of the ancestor function
-In this section, we provide the formal definition for the function $\mathcal{A}(\cdot,\cdot)$ within the Dynamic Patch Fusion (DPF) module, as introduced in Section 3.2.
-Consider the $n$-th coarsest patch, which is initially partitioned at the coarsest granularity, $p_{S}$.
-As this coarsest patch is processed by the Dynamic Patch Router, $K$ experts are activated.
-Let $p_{k}$ denote the minimum patch size among these activated experts.
-The coarsest patch is subsequently partitioned into $M_{n}=\frac{p_{S}}{p_{k}}$ non-overlapping finest patches.
-We denote the $m$-th such finest patch as $\bm{x}_{n,m}^{p_{k}}$.
-The function $\mathcal{A}(\bm{x}_{n,m}^{p_{k}},p_{i})$ is then defined as the operation that retrieves the ancestor patch of size $p_{i}$ that contains the finest patch $\bm{x}_{n,m}^{p_{k}}$.
-Formally, this is expressed as:
-$\mathcal{A}(\bm{x}_{n,m}^{p_{k}},p_{i})=[\bm{x}_{\lfloor\frac{(m-1)\cdot p_{k}}{p_{i}}\rfloor\frac{p_{i}}{p_{k}}+1}^{p_{k}},\dots,\bm{x}_{\lfloor\frac{(m-1)\cdot p_{k}}{p_{i}}+1\rfloor\frac{p_{i}}{p_{k}}}^{p_{k}}].$ (15)
+In practical implementation, $\boldsymbol{b}_{i}$ is updated once per batch, i.e., $L_{i}$ denotes the sum of the normalized gating values for the $i$-th expert across all sequences within the entire batch.
+At the end of each step, the bias term $\boldsymbol{b}_{i}$ for each expert is adjusted.
+This dynamic adjustment of $\boldsymbol{b}_{i}$ aims to balance the workload across the experts according to the desired distribution by influencing future top-K selections, while also steering the patches-to-expert affinity scores $s^{\prime}_{n,i}$ in subsequent batches, ensuring that the actual load distribution $L_{i}$ progressively aligns with the target distribution $\tau_{i}$, thereby promoting balanced expert utilization over time.
+D.3.2 Definition of the Expansion Function
+In this section, we provide the formal definition for the function $\mathrm{Expand}(\cdot)$ within the Mixture-of-Size Fusion stage, as introduced in Section 3.2.
+Consider a single time series segment $\mathbf{x}$ of length $P$.
+During the Top-$K$ Granularity Routing, a set of valid experts $\mathcal{I}_{\mathrm{valid}}$ is activated.
+Let $P_{\min}$ denote the finest granularity among these activated experts:
+$P_{\min}=\min\{P_{j}\mid j\in\mathcal{I}_{\mathrm{valid}}\}.$ (14)
+Consequently, the target sequence length for alignment is defined as $M=\lceil P/P_{\min}\rceil$.
+For an activated granularity level $i$ with patch size $P_{i}$, the encoding process yields a sequence of embeddings $\boldsymbol{E}_{i}=[\boldsymbol{e}_{i,1},\dots,\boldsymbol{e}_{i,M_{i}}]$, where $M_{i}=\lceil P/P_{i}\rceil$.
+Due to the nested design of patch sizes, the ratio $R_{i}=P_{i}/P_{\min}$ is strictly a positive integer.
+The function $\mathrm{Expand}(\boldsymbol{E}_{i})$ maps the coarser embedding sequence to the finest temporal resolution by repeating each embedding vector $R_{i}$ times.
+Formally, the $k$-th element of the expanded sequence $\tilde{\boldsymbol{E}}_{i}\in\mathbb{R}^{M\times D_{h}}$ is derived from the original sequence via index mapping:
+$\tilde{\boldsymbol{e}}_{i,m}=\boldsymbol{e}_{i,\lceil m/R_{i}\rceil},\quad\text{for }m=1,\dots,M.$ (15)
+In vector notation, this operation can be expressed as:
+$\mathrm{Expand}(\boldsymbol{E}_{i})=[\underbrace{\boldsymbol{e}_{i,1},\dots,\boldsymbol{e}_{i,1}}_{R_{i}\text{ times}},\dots,\underbrace{\boldsymbol{e}_{i,M_{i}},\dots,\boldsymbol{e}_{i,M_{i}}}_{R_{i}\text{ times}}].$ (16)
 We now illustrate the computational process with the following example. Let the setup be as follows:
-  • Set of available patch sizes: $\{p_{1},\dots,p_{S}\}=\{32,64,128\}$.
+  • Segment length: $P=128$.
+  • Set of available granularity patch sizes: $\{P_{1},P_{2},P_{3}\}=\{32,64,128\}$.
   • Number of null experts: $Z=2$.
-  • Number of activated experts: $K=3$.
-  • Sequence context length: $T=2048$.
-The sequence is partitioned using the maximum patch size $p_{S}=128$, resulting in $N=\lceil\frac{T}{p_{S}}\rceil=16$ coarsest patches.
-For the first coarsest patch ($n=1$), suppose the activated expert indices are {1, 3, 4}.
-Since experts 4 and 5 are designated as null experts, the fusion granularities are determined by the non-null activated experts, corresponding to patch sizes $p_{1}=32$ and $p_{3}=128$.
-The minimum patch size among these is $p_{k}=32$.
-Consequently, this patch is partitioned into $M_{n}=\frac{p_{S}}{p_{k}}=\frac{128}{32}=4$ non-overlapping finest patches.
-For the third finest patch ($m=3$), the ancestor patches are computed as follows:
-  • For $i=1$ and $p_{i}=p_{1}=32$:
-    $\mathcal{A}(\bm{x}_{1,3}^{32},32)=\left[\bm{x}_{\lfloor\frac{(3-1)\cdot 32}{32}\rfloor\frac{32}{32}+1}^{32},\dots,\bm{x}_{\lfloor\frac{(3-1)\cdot 32}{32}+1\rfloor\frac{32}{32}}^{32}\right]=[\bm{x}_{3}^{32}].$ (16)
-    The ancestor patch is the finest patch itself.
-  • For $i=2$ and $p_{i}=p_{2}=64$:
-    $\mathcal{A}(\bm{x}_{1,3}^{32},64)=\left[\bm{x}_{\lfloor\frac{(3-1)\cdot 32}{64}\rfloor\frac{64}{32}+1}^{32},\dots,\bm{x}_{\lfloor\frac{(3-1)\cdot 32}{64}+1\rfloor\frac{64}{32}}^{32}\right]=[\bm{x}_{3}^{32},\bm{x}_{4}^{32}].$ (17)
-    Expert 2 is not activated, so its normalized gating value, $\alpha_{n,i}$, in Equation 3 would be zero, meaning it does not contribute to the information fusion.
-  • For $i=3$ and $p_{i}=p_{3}=128$:
-    $\mathcal{A}(\bm{x}_{1,3}^{32},128)=\left[\bm{x}_{\lfloor\frac{(3-1)\cdot 32}{128}\rfloor\frac{128}{32}+1}^{32},\dots,\bm{x}_{\lfloor\frac{(3-1)\cdot 32}{128}+1\rfloor\frac{128}{32}}^{32}\right]=[\bm{x}_{1}^{32},\bm{x}_{2}^{32},\bm{x}_{3}^{32},\bm{x}_{4}^{32}].$ (18)
-This mechanism, enabled by the function $\mathcal{A}(\cdot,\cdot)$, allows for the dynamic selection of different granularities for information fusion, tailored to the information density of each coarsest patch.
-Consequently, it effectively addresses the challenge of time series heterogeneity inherent in TSFMs.
-C.4 Instance-Adaptive Rotary Position Embedding (IARoPE)
-In this section, we provide a more detailed description of the IARoPE implementation discussed in Section 3.3.
-We first introduce the principles of Rotary Position Embedding (RoPE) (Su et al., 2024) in detail, and then provide corresponding supplementary information on the IARoPE implementation.
-C.4.1 Details of RoPE
+  • Top-$K$ selection: $K=2$.
+Suppose for a specific segment $\mathbf{x}$, the router activates experts corresponding to granularity levels 1 and 3 (i.e., $\mathcal{I}_{\mathrm{valid}}=\{1,3\}$).
+The finest activated resolution is $P_{\min}=\min(32,128)=32$.
+The target aligned sequence length is $M=128/32=4$.
+The expansion process for each expert is computed as follows:
+  • For Expert 1 ($P_{1}=32$):
+    The number of original patches is $M_{1}=128/32=4$.
+The embedding sequence is $\boldsymbol{E}_{1}=[\boldsymbol{e}_{1,1},\boldsymbol{e}_{1,2},\boldsymbol{e}_{1,3},\boldsymbol{e}_{1,4}]$.
+The repetition factor is $R_{1}=32/32=1$.
+    $\tilde{\boldsymbol{E}}_{1}=\mathrm{Expand}(\boldsymbol{E}_{1})=[\boldsymbol{e}_{1,1},\boldsymbol{e}_{1,2},\boldsymbol{e}_{1,3},\boldsymbol{e}_{1,4}].$ (17)
+    Since this expert operates at the finest activated resolution, the expansion is an identity mapping.
+  • For Expert 3 ($P_{3}=128$):
+    The number of original patches is $M_{3}=128/128=1$. The embedding sequence consists of a single vector $\boldsymbol{E}_{3}=[\boldsymbol{e}_{3,1}]$. The repetition factor is $R_{3}=128/32=4$.
+    $\tilde{\boldsymbol{E}}_{3}=\mathrm{Expand}(\boldsymbol{E}_{3})=[\boldsymbol{e}_{3,1},\boldsymbol{e}_{3,1},\boldsymbol{e}_{3,1},\boldsymbol{e}_{3,1}].$ (18)
+    The single coarse embedding is broadcast across all 4 time steps to align with the target resolution.
+  • Hypothetical Case for Expert 2 ($P_{2}=64$):
+    Although Expert 2 was not activated in this example, if it were processed, $M_{2}=128/64=2$ and $R_{2}=64/32=2$. The expansion would be:
+    $\tilde{\boldsymbol{E}}_{2}=[\boldsymbol{e}_{2,1},\boldsymbol{e}_{2,1},\boldsymbol{e}_{2,2},\boldsymbol{e}_{2,2}].$ (19)
+By explicitly aligning all representations to the finest resolution $P_{\min}$ via $\mathrm{Expand}(\cdot)$, Kairos ensures that the subsequent weighted summation $\bar{\boldsymbol{E}}=\sum_{i\in\mathcal{I}_{\mathrm{valid}}}\alpha_{i}\,\tilde{\boldsymbol{E}}_{i}$ is performed on element-wise compatible tensors, effectively fusing global context with local details.
+D.4 Dynamic Rotary Position Embedding (DRoPE)
+In this section, we provide a more detailed description of the DRoPE implementation discussed in Section 3.3.
+We first introduce the principles of Rotary Position Embedding (RoPE) (Su et al., 2024) in detail, and then provide corresponding supplementary information on the DRoPE implementation.
+D.4.1 Details of RoPE
 The core idea of RoPE is to rotate segments of the query and key vectors by an angle proportional according to their absolute position in the sequence.
 This allows the model to discern relative positions through the geometry of these rotations, without needing explicit calculation of relative distances.
+Throughout this subsection, we use $m$ and $n$ to denote generic token positions.
 Specifically, RoPE operates on vectors of an even dimension, denoted as $D_{h}$.
-For a vector $\bm{e}$ (representing a query $\mathbf{q}$ or a key $\mathbf{k}$ in self-attention) at position $m$, it is transformed by a rotation matrix $\mathbf{R}_{m}$.
+For a vector $\boldsymbol{e}$ (representing a query $\mathbf{q}$ or a key $\mathbf{k}$ in self-attention) at position $m$, it is transformed by a rotation matrix $\mathbf{R}_{m}$.
 This matrix is block-diagonal, composed of $D_{h}/2$ individual $2\times 2$ rotation blocks.
-Each block $\mathbf{R}_{m,j}$ acts on a pair of dimensions $(e_{2j},e_{2j+1})$ of the vector:
-As introduced in Section 3.3, RoPE applies a rotation to input vector. This rotation is applied to pairs of dimensions $(e_{2j},e_{2j+1})$, and can be concisely expressed using complex form:
-$f_{\text{RoPE}}(\bm{e},m)=(e_{2j}+ie_{2j+1})e^{im\theta_{j}},$ (19)
-where $\theta_{j}$ is the angular frequency.
-This multiplication by $e^{im\theta_{j}}$ corresponds to a rotation in the complex plane.
-For the real-valued components $e_{2j}$ and $e_{2j+1}$, this operation is equivalent to applying the following $2\times 2$ rotation matrix $\mathbf{R}_{m,j}$:
-$\mathbf{R}_{m,j}=\begin{bmatrix}\cos(m\theta_{j})&-\sin(m\theta_{j})\\ \sin(m\theta_{j})&\cos(m\theta_{j})\end{bmatrix}$, (20)
+Each block $\mathbf{R}_{m,d}$ acts on a pair of dimensions $(e_{2d},e_{2d+1})$ of the vector:
+As introduced in Section 3.3, RoPE applies a rotation to input vector. This rotation is applied to pairs of dimensions $(e_{2d},e_{2d+1})$, and can be concisely expressed using complex form:
+$f_{\text{RoPE}}(\boldsymbol{z},m)=(z_{2d}+iz_{2d+1})e^{im\theta_{d}},$ (20)
+where $\theta_{d}$ is the angular frequency.
+Here, $i$ denotes the imaginary unit.
+This multiplication by $e^{im\theta_{d}}$ corresponds to a rotation in the complex plane.
+For the real-valued components $z_{2d}$ and $z_{2d+1}$, this operation is equivalent to applying the following $2\times 2$ rotation matrix $\mathbf{R}_{m,d}$:
+$\mathbf{R}_{m,d}=\begin{bmatrix}\cos(m\theta_{d})&-\sin(m\theta_{d})\\ \sin(m\theta_{d})&\cos(m\theta_{d})\end{bmatrix}$, (21)
 Here, $m$ is the absolute position of the token.
-The term $\theta_{j}$ represents a predefined angular frequency for the $j$-th pair of dimensions ($j\in[0,D_{h}/2-1]$), typically defined as $\theta_{j}=b^{-2j/D_{h}}$ where $b=10000$ in RoPE (Su et al., 2024) original setting.
+The term $\theta_{d}$ represents a predefined angular frequency for the $d$-th pair of dimensions ($d\in[0,D_{h}/2-1]$), typically defined as $\theta_{d}=b^{-2d/D_{h}}$
 The full rotation matrix $\mathbf{R}_{m}$ for position $m$ is thus:
@@
 This allows RoPE to integrate relative positional information into the self-attention without any additional learnable parameters or explicit relative position computations.
-C.4.2 Details of IARoPE
-The initial RoPE frequencies $\theta_{\text{init},j}$ typically exhibit a wide numerical range.
-For instance, in our setting $\theta_{\text{init},0}=1.0\ (j=0)$, while $\theta_{\text{init},31}\approx 1.33\times 10^{-4}$ ($j=31$, corresponding to $D_{h}/2-1$ and $D_{h}=64$).
+Equation 25 confirms that the interaction relies on the relative position nmn-m through the full rotation matrix $\mathbf{R}_{n-m}$.
+Since this matrix is block-diagonal, the total interaction is actually composed of independent rotations on disjoint 2D subspaces.
+The following theorem demonstrates the periodicity of the attention score within each subspace.
+Theorem D.1 (Periodic dependence of RoPE attention).
+Consider the $d$-th 2D subspace of the hidden state under RoPE, with angular frequency $\theta_{d}$, and let $\mathrm{atten}_{d}(m,n)$ denote this subspace’s contribution to the dot-product attention between positions $m$ and $n$.
+Then there exist an amplitude $A_{d}(m,n)\geq 0$ and a phase $\varphi_{d}(m,n)$, depending only on the content vectors at $m$ and $n$, such that
+$\begin{split}\mathrm{atten}_{d}(m,n)&=A_{d}(m,n)\,\cos\big(\theta_{d}(m-n)+\varphi_{d}(m,n)\big)\\ &\;\propto\;\cos\big(\theta_{d}(m-n)+\varphi_{d}(m,n)\big).\end{split}$ (26)
+In particular, for fixed content, $\mathrm{atten}_{d}(m,n)$ is a periodic function of the relative distance (mn)(m-n) with period $2\pi/\theta_{d}$.
+Proof.
+Let $q_{m,d},k_{n,d}\in\mathbb{C}$ be the complex-valued representations of the $d$-th 2D subspace for the query and key at positions $m$ and $n$, respectively, as defined above.
+Applying RoPE multiplies these components by phase factors $e^{im\theta_{d}}$ and $e^{in\theta_{d}}$, yielding
+$q^{\prime}_{m,d}=q_{m,d}e^{im\theta_{d}},\qquad k^{\prime}_{n,d}=k_{n,d}e^{in\theta_{d}}.$ (27)
+The contribution of this subspace to the dot-product attention is the real part of $q^{\prime}_{m,d}(k^{\prime}_{n,d})^{*}$:
+$\mathrm{atten}_{d}(m,n)=\mathrm{Re}\!\big[q^{\prime}_{m,d}(k^{\prime}_{n,d})^{*}\big]=\mathrm{Re}\!\big[q_{m,d}k_{n,d}^{*}e^{i\theta_{d}(m-n)}\big].$ (28)
+Writing $q_{m,d}k_{n,d}^{*}$ in polar form as $q_{m,d}k_{n,d}^{*}=A_{d}(m,n)e^{i\varphi_{d}(m,n)}$ with $A_{d}(m,n)\geq 0$ and $\varphi_{d}(m,n)\in\mathbb{R}$, we obtain
+$\begin{split}\mathrm{atten}_{d}(m,n)&=\mathrm{Re}\!\big[A_{d}(m,n)e^{i(\theta_{d}(m-n)+\varphi_{d}(m,n))}\big]\\ &=A_{d}(m,n)\,\cos\big(\theta_{d}(m-n)+\varphi_{d}(m,n)\big),\end{split}$ (29)
+which shows that $\mathrm{atten}_{d}(m,n)$ is proportional to a cosine function of the relative distance (mn)(m-n), with angular frequency $\theta_{d}$.
+The periodicity with period $2\pi/\theta_{d}$ follows immediately from the periodicity of the cosine function.
+∎
+D.4.2 Details of DRoPE
+The initial RoPE frequencies $\theta_{d}$ typically exhibit a wide numerical range.
+For instance, in our setting $\theta_{0}=1.0\ (d=0)$, while $\theta_{31}\approx 1.33\times 10^{-4}$ ($d=31$, corresponding to $D_{h}/2-1$ and $D_{h}=64$).
 Directly applying affine modulation to these values could lead to numerical instability or disproportionate adjustments due to the vast scale differences.
-To address this and ensure stable and effective modulation across the entire range of base frequencies, our IARoPE performs the adaptation in log-space.
-The layer-specific modulation parameters, $\gamma^{(l)}_{j}$ and $\beta^{(l)}_{j}$, predicted by the Multilayer Perceptron (MLP) for layer $l$ following Algorithm 1, are applied to the log-transformed base frequencies $\log\theta_{\text{init},j}$ via an element-wise affine transformation as
-$\log\theta^{\prime(l)}_{j}=\gamma^{(l)}_{j}\odot\log\theta_{\text{init},j}+\beta^{(l)}_{j}$, (25)
-where $\log\theta^{\prime(l)}_{j}$ represents the modulated log-frequencies for layer $l$ and dimension pair $j$.
-Then, these modulated log-frequencies are transformed back to their original scale by exponentiation to obtain the adaptive rotation frequencies $\theta^{\prime(l)}_{j}$ used in the IARoPE calculation for layer $l$ as
-$\theta^{\prime(l)}_{j}=\exp(\log\theta^{\prime(l)}_{j})$. (26)
+To address this and ensure stable and effective modulation across the entire range of base frequencies, our DRoPE performs the adaptation in log-space.
+The layer-specific modulation parameters, $\gamma^{(l)}_{d}$ and $\beta^{(l)}_{d}$, predicted by the Multilayer Perceptron (MLP) for layer $l$ following Algorithm 1, are applied to the log-transformed base frequencies $\log\theta_{d}$ via an element-wise affine transformation as
+$\log\theta^{\prime(l)}_{\text{inst},d}=\gamma^{(l)}_{d}\odot\log\theta_{d}+\beta^{(l)}_{d}$, (30)
+where $\log\theta^{\prime(l)}_{\text{inst},d}$ represents the modulated log-frequencies for layer $l$ and dimension pair $d$.
+Then, these modulated log-frequencies are transformed back to their original scale by exponentiation to obtain the adaptive rotation frequencies $\theta^{\prime(l)}_{\text{inst},d}$ used in the DRoPE calculation for layer $l$ as
+$\theta^{\prime(l)}_{\text{inst},d}=\exp(\log\theta^{\prime(l)}_{\text{inst},d})$. (31)
 Algorithm 1 Generating Instance Adaptive Parameters ($\gamma^{(l)},\beta^{(l)}$)
-1:Time series instance $X\in\mathbb{R}^{B\times L}$, mask $M\in\{0,1\}^{B\times L}$, batch size $B$, sequence length $L$, FFT feature dimension $w$, MLP network $f_{\text{MLP}}$
-2:Adaptive parameters $\gamma^{(l)}\in\mathbb{R}^{B\times D_{h}/2}$, $\beta^{(l)}\in\mathbb{R}^{B\times D_{h}/2}$
+0: Time series instance $X\in\mathbb{R}^{B\times L}$, mask $M\in\{0,1\}^{B\times L}$, batch size $B$, sequence length $L$, FFT feature dimension $w$, MLP network $f_{\text{MLP}}$
+0: Adaptive parameters $\gamma^{(l)}\in\mathbb{R}^{B\times D_{h}/2}$, $\beta^{(l)}\in\mathbb{R}^{B\times D_{h}/2}$
 […]
@@
 […]
-9:return $\gamma^{(l)},\beta^{(l)}$
-Appendix D Calculation of information density
-In this section, we present the formula for the information density illustrated in Figure 1(b).
-To effectively characterize and compare different time series datasets, we introduce a method to quantify their information density and the variation of this density over time.
-For this, similar to TimeMixer (Wang et al., 2024a), we utilize spectral entropy, a metric that measures the complexity and compressibility of a signal.
-In this context, a signal with low spectral entropy, such as one with a few dominant periodic components, is considered to have low information density due to its redundant and predictable nature.
-Conversely, a signal with high spectral entropy, resembling white noise, has its power spread broadly across the frequency spectrum, indicating a high degree of randomness and thus a higher information density.
-Given the observations $\mathbf{x}_{1:T}=(x_{1},\ldots,x_{T})\in\mathbb{R}^{T}$ of a specific dataset, we analyze the time series using a sliding window approach.
-The series is segmented into $N$ windows, $\mathbf{w}_{i}$, each of size $M$ with a step size of $S$.
-In our analysis, we use a non-overlapping configuration where both the window size and the step size are set to 128 (i.e., $M=128$, $S=128$).
-The total number of windows is $N=\lfloor(T-M)/S\rfloor+1$.
-To mitigate spectral leakage from windowing, we apply a Hamming window to each segment $\mathbf{w}_{i}$:
-$h[n]=0.5-0.5\cos\left(\frac{2\pi n}{M-1}\right),\quad 0\leq n\leq M-1.$ (27)
-For each resulting windowed segment $\mathbf{w}^{\prime}_{i}=\mathbf{w}_{i}\cdot h$, we compute its normalized power spectral density, $p_{i}[k]$, which describes how the signal’s power is distributed over different frequencies.
-The spectral entropy for the $i$-th window, $H_{SE}(\mathbf{w}_{i})$, is then calculated using the Shannon entropy formula:
-$H_{SE}(\mathbf{w}_{i})=-\sum_{k=0}^{M-1}p_{i}[k]\log_{2}(p_{i}[k]).$ (28)
-This process yields a sequence of entropy values, $(H_{SE}(\mathbf{w}_{1}),\ldots,H_{SE}(\mathbf{w}_{N}))$, where each value represents the localized information density of its corresponding time segment.
-To obtain a holistic view of the dataset’s characteristics, we compute two key statistics from this entropy sequence.
-First, the mean spectral entropy ($\mu_{SE}$) serves as a measure of the average information density of the entire dataset.
-A higher $\mu_{SE}$ suggests that the dataset, on average, contains more complex and less predictable patterns.
-This allows for a direct comparison of the overall information content between different datasets.
-$\mu_{SE}=\frac{1}{N}\sum_{i=1}^{N}H_{SE}(\mathbf{w}_{i}).$ (29)
-Second, the standard deviation of spectral entropy ($\sigma_{SE}$) quantifies the variability of information density within the dataset.
-A small $\sigma_{SE}$ indicates that the dataset is stationary in its complexity, with a consistent level of information density throughout.
-A large $\sigma_{SE}$, however, reveals a non-stationary character, signifying substantial fluctuations in the signal’s complexity and information content over time.
-$\sigma_{SE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(H_{SE}(\mathbf{w}_{i})-\mu_{SE})^{2}}.$ (30)
-Together, $\mu_{SE}$ and $\sigma_{SE}$ provide a concise yet powerful summary of a dataset’s informational characteristics, enabling a quantitative assessment of its average complexity and internal dynamics.
+7:return $\gamma^{(l)},\beta^{(l)}$
+D.4.3 Interpretation of DRoPE
+From the theorem D.1, the base RoPE parameters $\{\theta_{d}\}_{d=0}^{D_{h}/2-1}$ define a fixed bank of sinusoidal kernels over relative positions.
+Each $\theta_{d}$ controls the frequency of one cosine kernel.
+This fixed frequencies works reasonably well when sequences share a homogeneous notion of temporal scale, but it can become suboptimal for a TSFM that must handle highly heterogeneous sampling intervals and spectral patterns across domains.
+DRoPE addresses this by making these frequencies instance-adaptive.
+The FFT-based module in Algorithm 1 extracts simple spectral statistics $X^{\prime}_{\text{FFT}}$ from each input sequence and maps them, via an MLP, to instance- and layer-specific parameters $\gamma^{(l)}_{d}$ and $\beta^{(l)}_{d}$.
+These parameters modulate the base frequencies in log-space:
+$\log\theta^{\prime(l)}_{\text{inst},d}=\gamma^{(l)}_{d}\odot\log\theta_{d}+\beta^{(l)}_{d},\quad\theta^{\prime(l)}_{\text{inst},d}=\exp\big(\log\theta^{\prime(l)}_{\text{inst},d}\big)$, (32)
+yielding an adapted frequency grid $\{\theta^{\prime(l)}_{d}\}$ for each layer and sequence.
+Intuitively, this log-space affine transformation stretches or compresses the original RoPE frequency grid in a sequence-dependent way, while preserving the relative-position nature of RoPE.
+The resulting $\theta^{\prime(l)}_{\text{inst},d}$ still define sinusoidal kernels over relative lags, but their effective frequencies are now gently steered by the spectrum of the current input.
+Crucially, DRoPE does not aim to recover a single true period for each series; real-world time series are often multi-periodic and non-stationary.
+Instead, we interpret $\{\theta^{\prime(l)}_{\text{inst},d}\}$ as a sequence-dependent frequency profile that shapes how attention depends on relative lags, providing a more flexible and data-driven positional bias than fixed RoPE.
 Appendix E Additional Details of Experiment Setting
@@
 The remaining datasets were classified into Tiers 4 and 5, based on a composite assessment of their size and pattern regularity.
-The specific details of these datasets, categorized by their respective sampling frequencies, are presented in Tables 5-8.
+The specific details of these datasets, categorized by their respective sampling frequencies, are presented in Tables 6-9.
 Synthetic data.
@@
 These feature a constant baseline with repeating events like trapezoidal spikes or inverted-U shaped dips, where the period, amplitude, and width of the events remain fixed across the entire series with no random jitter, as outlined in Algorithm 3.
+Synthetic dataset cases are provided in Appendix H.1.
+Distribution Analysis.
+To investigate whether the synthetic dataset supplements distributions absent in the real-world data, we adopted the statistical methodology of Toto (Cohen et al., 2025).
+We employed the ARCH-LM Statistic and Spectral Entropy to quantify time-varying volatility and information density (complexity), respectively.
+These metrics align closely with the design motivation for the Mixture-of-Size Encoding and DRoPE components of Kairos.
+As illustrated in Figures 6 (a) and (b), we observed that a significant portion of the real-world dataset exhibits ARCH-LM Statistic values approaching zero and Spectral Entropy values nearing one, indicating high time-varying volatility and complexity.
+In contrast, the synthetic dataset exhibits the opposite characteristics, exposing the model during training to time series with more constant volatility and greater regularity.
+This exposure enhances the model’s generalization capabilities.
+Furthermore, we statistically compared the periodicity of the real-world and synthetic datasets by plotting their respective Cumulative Distribution Functions (CDFs).
+As shown in Figure 6 (c), the dominant periods in the real-world dataset are heavily concentrated.
+This can cause the model to overfit to these specific periodicities and fail to learn generalizable periodic patterns.
+Conversely, the synthetic dataset presents a significantly smoother distribution of periods, enabling the model to generalize effectively across arbitrary periodicities.
+Consequently, the synthetic dataset effectively mitigates these biases present in the real-world data, thereby exposing the model to a more comprehensive and diverse set of distributions.
+Figure 6: Comparison of real and synthetic datasets across (a) ARCH-LM statistic, (b) spectral entropy, and (c) dominant period distributions, illustrating complementary volatility, complexity, and periodicity characteristics.
 Table 6: Detailed descriptions of second-level, minute-level, and hourly datasets.
@@
 Residential PV Power Energy T 699 376,016,850
-Loop Seattle Transport 5T 323 33,953,760
 Los-Loop Transport 5T 207 7,094,304
@@
 SHMetro Transport 15T 576 5,073,984
-SZ-Taxi Transport 15T 156 464,256
-M DENSE Transport 15T 30 525,600
-Electricity (15 Min.) Energy 15T 370 41,936,458
 Beijing Subway Transport 30T 552 867,744
@@
 Taxi Transport 30T 2428 3,589,798
-Solar Energy H 5,166 543,049,920
 BDG-2 Bear Energy H 91 1,482,312
@@
 IDEAL Energy H 217 1,255,253
-KDD Cup 2018 Energy H 270 2,897,004
 Low Carbon London Energy H 713 9,543,553
@@
 ERCOT Load Energy H 8 1,238,976
-M4 Hourly Econ/Fin H 414 373,372
 Mexico City Bikes Transport H 494 38,687,004
-Electricity (Hourly) Energy H 321 8,443,584
 Beijing Air Quality Nature H 132 4,628,448
-KDD Cup 2018 Nature H 270 2,942,364
 Pedestrian Counts Transport H 66 3,132,346
@@
 Traffic Transport H 862 15,122,928
-Solar (Hourly) Energy H 5,166 45,254,160
 Taxi (Hourly) Transport H 2,428 1,794,292
@@
 Favorita Transactions Sales D 54 84,408
-Hierarchical Sales Sales D 118 212,164
-Restaurant Sales D 809 289,402
 Subseasonal Climate D 3,448 56,788,560
@@
 Dominick Retail D 100,014 29,652,492
-M4 Daily Econ/Fin D 4,227 10,023,836
 M5 Sales D 30,490 47,649,940
-COVID Deaths Healthcare D 266 56,392
-Temperature Rain Nature D 32,072 22,290,040
 Monash Weather Climate D 3,010 43,032,000
@@
 Traffic Weekly Transport W 862 82,752
-M4 Weekly Econ/Fin W 359 371,579
-Electricity (Weekly) Energy W 321 50,076
 NN5 Weekly Econ/Fin W 111 12,543
@@
 GoDaddy Econ/Fin M 6,270 257,070
-M4 Monthly Econ/Fin M 48,000 11,246,411
-Car Parts Sales M 2,674 136,374
 CIF 2016 Econ/Fin M 72 7,108
 FRED MD Econ/Fin M 107 77,896
-Hospital Healthcare M 767 64,428
 M1 Monthly Econ/Fin M 617 55,998
@@
 M3 Other Econ/Fin Q 174 11,933
-M4 Quarterly Econ/Fin Q 24,000 2,406,108
 M1 Quarterly Econ/Fin Q 203 9,944
@@
 Tourism Quarterly Econ/Fin Q 427 42,544
-M4 Yearly Econ/Fin Y 23,000 858,458
 M1 Yearly Econ/Fin Y 181 4,515
@@
 Algorithm 2 Composite Time Series Generation
-1:Time series length $L=4096$, Primary period set $\mathcal{P}_{1}=\{24,48,288,360\}$, Harmonic multiplier $n=7$, Trend types $\mathcal{T}_{\text{types}}=\{\text{linear, exp, ARMA}\}$, Seasonal patterns $\mathcal{S}_{\text{patterns}}=\{\text{spike, interpolated segment}\}$.
-2:A synthetic time series $\bm{x}_{1:L}$.
+0: Time series length $L=4096$, Primary period set $\mathcal{P}_{1}=\{24,48,288,360\}$, Harmonic multiplier $n=7$, Trend types $\mathcal{T}_{\text{types}}=\{\text{linear, exp, ARMA}\}$, Seasonal patterns $\mathcal{S}_{\text{patterns}}=\{\text{spike, interpolated segment}\}$.
+0: A synthetic time series $\boldsymbol{x}_{1:L}$.
 […]
-4:Sample flags $f_{s},f_{t},f_{n}$ $\triangleright$ Determine component inclusion, ensuring $f_{s}\lor f_{t}$ is true
-5:if $f_{s}$ is true then
-6:  $k\sim\text{Bernoulli}(0.2)$ $\triangleright$ $k=1$ for double period (20% prob), $k=0$ for single
-7:  $p_{1}\sim\mathcal{U}(\mathcal{P}_{1})$ $\triangleright$ Sample primary period
-8:  $\mathcal{P}_{\text{active}}\leftarrow\{p_{1}\}$
-9:  if $k=1$ then
-10:   $\mathcal{P}_{\text{active}}\leftarrow\mathcal{P}_{\text{active}}\cup\{n\cdot p_{1}\}$
-11:  end if
-12:  for $p$ in $\mathcal{P}_{\text{active}}$ do
-13:   $a\sim\mathcal{U}(1.0,3.0)$ $\triangleright$ Sample amplitude for this component
-14:   $pattern\sim\mathcal{U}(\mathcal{S}_{\text{patterns}})$
-15:   $\bm{c}\leftarrow\text{GeneratePattern}(pattern,p,a)$ $\triangleright$ Create a single cycle of the pattern
-16:   $\bm{s}\leftarrow\bm{s}+\text{Tile}(\bm{c},L)$ $\triangleright$ Tile the cycle to length $L$ and add to seasonal component
-17:  end for
-18:end if
-19:if $f_{t}$ is true then
-20:  $type\sim\mathcal{U}(\mathcal{T}_{\text{types}})$
-21:  $\bm{t}\leftarrow\text{GenerateTrend}(type,L)$
-22:  if $f_{s}$ is true then
-23:   $\lambda\sim\mathcal{U}(0.1,0.3)$ $\triangleright$ Reduce trend strength when seasonality is present
-24:   $\bm{t}\leftarrow\lambda\cdot\bm{t}$
-25:  end if
-26:end if
-27:if $f_{n}$ is true then
-28:  $\sigma_{n}\sim\mathcal{U}(0.01,0.1)$
-29:  $\bm{n}\sim\mathcal{N}(\bm{0},\sigma_{n}^{2}\mathbf{I})$ $\triangleright$ Generate white Gaussian noise
-30:end if
+2: Sample flags $f_{s},f_{t},f_{n}$ {Determine component inclusion, ensuring $f_{s}\lor f_{t}$ is true}
+3:if $f_{s}$ is true then
+4:  $k\sim\text{Bernoulli}(0.2)$ {$k=1$ for double period (20% prob), $k=0$ for single}
+5:  $p_{1}\sim\mathcal{U}(\mathcal{P}_{1})$ {Sample primary period}
+6:  $\mathcal{P}_{\text{active}}\leftarrow\{p_{1}\}$
+7:  if $k=1$ then
+8:   $\mathcal{P}_{\text{active}}\leftarrow\mathcal{P}_{\text{active}}\cup\{n\cdot p_{1}\}$
+9:  end if
+10:  for all $p$ in $\mathcal{P}_{\text{active}}$ do
+11:   $a\sim\mathcal{U}(1.0,3.0)$ {Sample amplitude for this component}
+12:   $pattern\sim\mathcal{U}(\mathcal{S}_{\text{patterns}})$
+13:   $\boldsymbol{c}\leftarrow\text{GeneratePattern}(pattern,p,a)$ {Create a single cycle of the pattern}
+14:   $\boldsymbol{s}\leftarrow\boldsymbol{s}+\text{Tile}(\boldsymbol{c},L)$ {Tile the cycle to length $L$ and add to seasonal component}
+15:  end for
+16:end if
+17:if $f_{t}$ is true then
+18:  $type\sim\mathcal{U}(\mathcal{T}_{\text{types}})$
+19:  $\boldsymbol{t}\leftarrow\text{GenerateTrend}(type,L)$
+20:  if $f_{s}$ is true then
+21:   $\lambda\sim\mathcal{U}(0.1,0.3)$ {Reduce trend strength when seasonality is present}
+22:   $\boldsymbol{t}\leftarrow\lambda\cdot\boldsymbol{t}$
+23:  end if
+24:end if
+25:if $f_{n}$ is true then
+26:  $\sigma_{n}\sim\mathcal{U}(0.01,0.1)$
+27:  $\boldsymbol{n}\sim\mathcal{N}(\boldsymbol{0},\sigma_{n}^{2}\mathbf{I})$ {Generate white Gaussian noise}
+28:end if
 […]
-32:return $\bm{x}_{1:L}$
+30:return $\boldsymbol{x}_{1:L}$
 Algorithm 3 Idealized Industrial Signal Generation
-1:Time series length $L=4096$, Pattern types $\mathcal{P}_{\text{types}}=\{\text{inverted\_u, spikes}\}$.
-2:A synthetic time series $\bm{x}_{1:L}$.
+0: Time series length $L=4096$, Pattern types $\mathcal{P}_{\text{types}}=\{\text{inverted\_u, spikes}\}$.
+0: A synthetic time series $\boldsymbol{x}_{1:L}$.
 […]
@@
 […]
-7:if $type=\text{inverted\_u}$ then
-8:  $sign\leftarrow-1$
-9:else
-10:  $sign\leftarrow+1$
-11:end if
-12:for $i\leftarrow 0,p,2p,\dots$ up to L1L-1 do
-13:  $start\leftarrow i$, $end\leftarrow\min(i+w,L)$
-14:  $\bm{x}_{start:end}\leftarrow\bm{x}_{start:end}+sign\cdot\bm{e}_{1:end-start}$ $\triangleright$ Add or subtract event shape at periodic intervals
-15:end for
-16:if $\sigma_{n}>0$ then
-17:  $\bm{x}\leftarrow\bm{x}+\mathcal{N}(\bm{0},\sigma_{n}^{2}\mathbf{I})$ $\triangleright$ Add global Gaussian noise
-18:end if
-19:return $\bm{x}_{1:L}$
+5:if $type=\text{inverted\_u}$ then
+6:  $sign\leftarrow-1$
+7:else
+8:  $sign\leftarrow+1$
+9:end if
+10:for all $i\leftarrow 0,p,2p,\dots$ up to L1L-1 do
+11:  $start\leftarrow i$, $end\leftarrow\min(i+w,L)$
+12:  $\boldsymbol{x}_{start:end}\leftarrow\boldsymbol{x}_{start:end}+sign\cdot\boldsymbol{e}_{1:end-start}$ {Add or subtract event shape at periodic intervals}
+13:end for
+14:if $\sigma_{n}>0$ then
+15:  $\boldsymbol{x}\leftarrow\boldsymbol{x}+\mathcal{N}(\boldsymbol{0},\sigma_{n}^{2}\mathbf{I})$ {Add global Gaussian noise}
+16:end if
+17:return $\boldsymbol{x}_{1:L}$
 E.2 Evaluation Datasets
@@
 Context length {96, 2048} {96, 2048} {96, 2048} {336, 512, 2048} {96, 2048} 512 2048 2048 2048 1536 2048 2048
-E.4 Details of IARoPE Analysis
-In this section, we explain in more detail the setting of the IARoPE analysis experiment discussed in Section 4.4.2.
-To more definitively ascertain whether these instance-specific modulations truly capture and leverage beneficial positional information derived from an instance’s FFT features, we designed a series of “shuffle” experiments.
+E.4 Details of DRoPE Analysis
+In this section, we explain in more detail the setting of the DRoPE analysis experiment discussed in Section 4.4.2.
+To more definitively ascertain whether these instance-specific modulations truly capture and leverage beneficial positional information derived from an instance’s FFT features, we designed a series of ?shuffle?
+experiments.
 These experiments function as a form of causal intervention analysis.
-The underlying hypothesis is: if the instance-derived $\theta$ modulations are crucial for IARoPE’s performance, then disrupting this linkage by applying $\theta$ modulations from one instance to another (shuffling) should lead to a noticeable degradation in forecasting accuracy.
+The underlying hypothesis is: if the instance-derived $\theta_{\text{inst}}$ modulations are crucial for DRoPE’s performance, then disrupting this linkage by applying $\theta_{\text{inst}}$ modulations from one instance to another (shuffling) should lead to a noticeable degradation in forecasting accuracy.
 The details of each group in the experiment are as follows:
-  • IARoPE: Our proposed method, where each time series instance utilizes its own FFT-derived features to independently modulate its RoPE $\theta$ parameters at each layer.
-  • Intra-Dataset Shuffle: During inference, the learned layer-specific $\theta$ modulation parameters ($\gamma^{(l)},\beta^{(l)}$) are randomly permuted among instances within the same batch and originating from the same dataset.
-The modulated $\theta$ are always shuffled in the same layer.
-  • Inter-Dataset Shuffle: For instances of a target dataset, layer-specific $\theta$ modulation parameters ($\gamma^{(l)},\beta^{(l)}$) are randomly sampled from a pre-computed collection derived from a different source dataset and applied to instances of the target dataset.
+  • DRoPE: Our proposed method, where each time series instance utilizes its own FFT-derived features to independently modulate its RoPE $\theta_{\text{inst}}$ parameters at each layer.
+  • Intra-Dataset Shuffle: During inference, the learned layer-specific $\theta_{\text{inst}}$ modulation parameters ($\gamma^{(l)},\beta^{(l)}$) are randomly permuted among instances within the same batch and originating from the same dataset.
+The modulated $\theta_{\text{inst}}$ are always shuffled in the same layer.
+  • Inter-Dataset Shuffle: For instances of a target dataset, layer-specific $\theta_{\text{inst}}$ modulation parameters ($\gamma^{(l)},\beta^{(l)}$) are randomly sampled from a pre-computed collection derived from a different source dataset and applied to instances of the target dataset.
 This process also ensures layer-wise correspondence of the applied modulations.
-  • Fixed RoPE: We use the fully trained IARoPE model but disable its instance-adaptive modulation during inference.
-Specifically, the prediction of $\gamma^{(l)}$ and $\beta^{(l)}$ is bypassed, and all instances revert to using the initial RoPE $\theta_{\text{init}}$ across all layers.
-This approach keeps the base model architecture and other learned weights identical to IARoPE, isolating the effect of the modulation.
-By comparing IARoPE’s performance against these shuffled and fixed $\theta$ configurations, we can better attribute performance gains or losses to the effectiveness of the instance-adaptive $\theta$ modulation process.
+  • Fixed RoPE: We use the fully trained DRoPE model but disable its instance-adaptive modulation during inference.
+Specifically, the prediction of $\gamma^{(l)}$ and $\beta^{(l)}$ is bypassed, and all instances revert to using the initial RoPE $\theta$ across all layers.
+This approach keeps the base model architecture and other learned weights identical to DRoPE, isolating the effect of the modulation.
+By comparing DRoPE’s performance against these shuffled and fixed $\theta$ configurations, we can better attribute performance gains or losses to the effectiveness of the instance-adaptive $\theta_{\text{inst}}$ modulation process.
 Appendix F Full Evaluation Results
@@
 96 192 336 720 AVG 96 192 336 720 AVG 96 192 336 720 AVG 96 192 336 720 AVG 96 192 336 720 AVG 96 192 336 720 AVG 96 192 336 AVG
-DLinear MSE 0.377 0.416 0.446 0.482 0.430 0.28 0.398 0.505 0.796 0.495 0.312 0.339 0.364 0.401 0.354 0.174 0.228 0.287 0.396 0.271 0.174 0.214 0.256 0.309 0.238 0.829 0.832 0.826 0.858 0.836 0.820 0.849 0.781 0.817 0.480 -
-MAE 0.391 0.416 0.44 0.495 0.436 0.358 0.434 0.495 0.634 0.480 0.359 0.375 0.392 0.415 0.385 0.272 0.312 0.355 0.428 0.342 0.238 0.272 0.304 0.346 0.290 0.502 0.503 0.506 0.513 0.506 0.569 0.571 0.551 0.564 0.424
-iTransformer MSE 0.383 0.441 0.478 0.47 0.443 0.282 0.367 0.402 0.406 0.364 0.327 0.363 0.390 0.424 0.376 0.336 0.369 0.398 0.428 0.383 0.171 0.218 0.278 0.357 0.256 0.778 0.795 0.800 0.839 0.803 0.988 1.041 1.007 1.012 0.501 -
-MAE 0.396 0.429 0.449 0.469 0.436 0.344 0.392 0.422 0.433 0.398 0.355 0.381 0.411 0.442 0.397 0.355 0.381 0.411 0.442 0.397 0.21 0.256 0.299 0.351 0.279 0.457 0.463 0.482 0.490 0.473 0.626 0.64 0.642 0.636 0.423
+DLinear MSE 0.377 0.416 0.446 0.482 0.430 0.280 0.398 0.505 0.796 0.495 0.312 0.339 0.364 0.401 0.354 0.174 0.228 0.287 0.396 0.271 0.174 0.214 0.256 0.309 0.238 0.829 0.832 0.826 0.858 0.836 0.820 0.849 0.781 0.817 0.480 -
+MAE 0.391 0.416 0.440 0.495 0.436 0.358 0.434 0.495 0.634 0.480 0.359 0.375 0.392 0.415 0.385 0.272 0.312 0.355 0.428 0.342 0.238 0.272 0.304 0.346 0.290 0.502 0.503 0.506 0.513 0.506 0.569 0.571 0.551 0.564 0.424
+iTransformer MSE 0.383 0.441 0.478 0.470 0.443 0.282 0.367 0.402 0.406 0.364 0.327 0.363 0.390 0.424 0.376 0.336 0.369 0.398 0.428 0.383 0.171 0.218 0.278 0.357 0.256 0.778 0.795 0.800 0.839 0.803 0.988 1.041 1.007 1.012 0.501 -
+MAE 0.396 0.429 0.449 0.469 0.436 0.344 0.392 0.422 0.433 0.398 0.355 0.381 0.411 0.442 0.397 0.355 0.381 0.411 0.442 0.397 0.210 0.256 0.299 0.351 0.279 0.457 0.463 0.482 0.490 0.473 0.626 0.640 0.642 0.636 0.423
 TimesNet MSE 0.408 0.456 0.508 0.567 0.485 0.315 0.435 0.481 0.455 0.422 0.358 0.428 0.494 0.550 0.458 0.175 0.251 0.307 0.412 0.286 0.174 0.229 0.282 0.360 0.261 1.177 1.268 1.129 1.136 1.178 0.817 0.777 0.751 0.782 0.544 -
@@
 MAE 0.404 0.427 0.436 0.445 0.428 0.324 0.369 0.386 0.405 0.371 0.360 0.379 0.399 0.422 0.390 0.263 0.308 0.340 0.397 0.327 0.194 0.238 0.275 0.324 0.258 0.479 0.493 0.491 0.493 0.489 0.560 0.562 0.535 0.552 0.397
-$\text{Moirai}_{b}$ MSE 0.429 0.494 0.531 0.544 0.500 0.275 0.347 0.379 0.388 0.347 0.456 0.459 0.482 0.518 0.479 0.194 0.251 0.292 0.354 0.273 0.170 0.218 0.270 0.344 0.251 1.129 1.175 1.142 1.177 1.156 1.208 1.243 1.171 1.207 0.579 1
+$\text{Moirai}_{b}$ MSE 0.429 0.494 0.531 0.544 0.500 0.275 0.347 0.379 0.388 0.347 0.456 0.459 0.482 0.518 0.479 0.194 0.251 0.292 0.354 0.273 0.170 0.218 0.270 0.344 0.251 1.129 1.175 1.142 1.177 1.156 1.208 1.243 1.171 1.207 0.579 2
 MAE 0.412 0.447 0.467 0.501 0.457 0.319 0.365 0.388 0.408 0.370 0.385 0.400 0.415 0.435 0.409 0.266 0.306 0.333 0.375 0.320 0.203 0.249 0.288 0.339 0.270 0.472 0.495 0.485 0.485 0.484 0.557 0.550 0.538 0.548 0.403
@@
 MAE 0.397 0.415 0.424 0.441 0.419 0.334 0.378 0.400 0.421 0.383 0.350 0.374 0.390 0.410 0.381 0.254 0.296 0.329 0.382 0.315 0.197 0.242 0.284 0.336 0.265 0.511 0.518 0.523 0.541 0.523 0.628 0.628 0.608 0.621 0.408
-$\text{ChronosBolt}_{s}$ MSE 0.405 0.475 0.514 0.516 0.478 0.255 0.339 0.388 0.400 0.346 0.309 0.375 0.420 0.512 0.404 0.166 0.231 0.289 0.400 0.272 0.152 0.205 0.264 0.357 0.245 0.863 0.956 1.001 1.103 0.981 0.762 0.802 0.749 0.771 0.489 3
+$\text{ChronosBolt}_{s}$ MSE 0.405 0.475 0.514 0.516 0.478 0.255 0.339 0.388 0.400 0.346 0.309 0.375 0.420 0.512 0.404 0.166 0.231 0.289 0.400 0.272 0.152 0.205 0.264 0.357 0.245 0.863 0.956 1.001 1.103 0.981 0.762 0.802 0.749 0.771 0.489 2
 MAE 0.388 0.426 0.453 0.485 0.438 0.307 0.362 0.398 0.418 0.371 0.317 0.357 0.389 0.441 0.376 0.237 0.284 0.324 0.393 0.310 0.186 0.237 0.279 0.338 0.260 0.410 0.458 0.492 0.551 0.478 0.433 0.446 0.473 0.451 0.381
@@
 MAE 0.384 0.411 0.422 0.425 0.411 0.314 0.364 0.391 0.407 0.369 0.329 0.362 0.384 0.407 0.371 0.246 0.290 0.323 0.378 0.309 0.181 0.229 0.273 0.331 0.254 0.424 0.439 0.443 0.450 0.439 0.437 0.444 0.426 0.436 0.367
-$\textsc{Kairos}_{s}$ MSE 0.385 0.429 0.446 0.422 0.421 0.269 0.340 0.380 0.394 0.346 0.305 0.356 0.381 0.404 0.362 0.163 0.226 0.278 0.382 0.262 0.151 0.197 0.253 0.340 0.235 0.920 0.943 0.942 1.002 0.952 0.808 0.851 0.815 0.825 0.473 4
+$\textsc{Kairos}_{s}$ MSE 0.385 0.429 0.446 0.422 0.421 0.269 0.340 0.380 0.394 0.346 0.305 0.356 0.381 0.404 0.362 0.163 0.226 0.278 0.382 0.262 0.151 0.197 0.253 0.340 0.235 0.920 0.943 0.942 1.002 0.952 0.808 0.851 0.815 0.825 0.473 6
 MAE 0.383 0.408 0.418 0.424 0.408 0.320 0.368 0.397 0.416 0.375 0.326 0.360 0.382 0.404 0.368 0.238 0.284 0.320 0.384 0.307 0.186 0.232 0.274 0.329 0.255 0.410 0.426 0.429 0.442 0.427 0.447 0.446 0.429 0.441 0.366
-$\textsc{Kairos}_{b}$ MSE 0.396 0.434 0.452 0.426 0.427 0.276 0.353 0.385 0.386 0.350 0.295 0.336 0.367 0.393 0.348 0.160 0.220 0.271 0.357 0.252 0.146 0.192 0.248 0.338 0.231 0.927 0.955 0.945 0.989 0.954 0.817 0.843 0.804 0.821 0.471 28
-MAE 0.385 0.408 0.418 0.427 0.410 0.319 0.369 0.394 0.412 0.374 0.322 0.355 0.380 0.402 0.365 0.238 0.283 0.317 0.373 0.303 0.182 0.228 0.270 0.330 0.253 0.414 0.425 0.427 0.432 0.425 0.446 0.449 0.434 0.443 0.364
-Appendix G Showcases
+$\textsc{Kairos}_{b}$ MSE 0.396 0.446 0.472 0.434 0.437 0.270 0.348 0.367 0.376 0.340 0.291 0.337 0.368 0.408 0.351 0.163 0.224 0.274 0.361 0.255 0.147 0.191 0.248 0.333 0.230 0.936 0.972 0.963 1.014 0.971 0.786 0.817 0.766 0.790 0.471 26
+MAE 0.385 0.414 0.429 0.432 0.415 0.325 0.378 0.395 0.410 0.377 0.321 0.355 0.378 0.405 0.365 0.239 0.283 0.317 0.376 0.304 0.181 0.227 0.270 0.326 0.251 0.406 0.424 0.428 0.437 0.424 0.445 0.446 0.429 0.440 0.365
+Appendix G Calculation of information density
+In this section, we present the formula for the information density illustrated in Figure 1(b).
+To effectively characterize and compare different time series datasets, we introduce a method to quantify their information density and the variation of this density over time.
+For this, similar to TimeMixer (Wang et al., 2024a), we utilize spectral entropy, a metric that measures the complexity and compressibility of a signal.
+In this context, a signal with low spectral entropy, such as one with a few dominant periodic components, is considered to have low information density due to its redundant and predictable nature.
+Conversely, a signal with high spectral entropy, resembling white noise, has its power spread broadly across the frequency spectrum, indicating a high degree of randomness and thus a higher information density.
+Given the observations $\mathbf{x}_{1:T}=(x_{1},\ldots,x_{T})\in\mathbb{R}^{T}$ of a specific dataset, we analyze the time series using a sliding window approach.
+The series is segmented into $N$ windows, $\mathbf{w}_{i}$, each of size $M$ with a step size of $S$.
+In our analysis, we use a non-overlapping configuration where both the window size and the step size are set to 128 (i.e., $M=128$, $S=128$).
+The total number of windows is $N=\lfloor(T-M)/S\rfloor+1$.
+To mitigate spectral leakage from windowing, we apply a Hamming window to each segment $\mathbf{w}_{i}$:
+$h[n]=0.5-0.5\cos\left(\frac{2\pi n}{M-1}\right),\quad 0\leq n\leq M-1.$ (33)
+For each resulting windowed segment $\mathbf{w}^{\prime}_{i}=\mathbf{w}_{i}\cdot h$, we compute its normalized power spectral density, $p_{i}[k]$, which describes how the signal’s power is distributed over different frequencies.
+The spectral entropy for the $i$-th window, $H_{SE}(\mathbf{w}_{i})$, is then calculated using the Shannon entropy formula:
+$H_{SE}(\mathbf{w}_{i})=-\sum_{k=0}^{M-1}p_{i}[k]\log_{2}(p_{i}[k]).$ (34)
+This process yields a sequence of entropy values, $(H_{SE}(\mathbf{w}_{1}),\ldots,H_{SE}(\mathbf{w}_{N}))$, where each value represents the localized information density of its corresponding time segment.
+To obtain a holistic view of the dataset’s characteristics, we compute two key statistics from this entropy sequence.
+First, the mean spectral entropy ($\mu_{SE}$) serves as a measure of the average information density of the entire dataset.
+A higher $\mu_{SE}$ suggests that the dataset, on average, contains more complex and less predictable patterns.
+This allows for a direct comparison of the overall information content between different datasets.
+$\mu_{SE}=\frac{1}{N}\sum_{i=1}^{N}H_{SE}(\mathbf{w}_{i}).$ (35)
+Second, the standard deviation of spectral entropy ($\sigma_{SE}$) quantifies the variability of information density within the dataset.
+A small $\sigma_{SE}$ indicates that the dataset is stationary in its complexity, with a consistent level of information density throughout.
+A large $\sigma_{SE}$, however, reveals a non-stationary character, signifying substantial fluctuations in the signal’s complexity and information content over time.
+$\sigma_{SE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(H_{SE}(\mathbf{w}_{i})-\mu_{SE})^{2}}.$ (36)
+Together, $\mu_{SE}$ and $\sigma_{SE}$ provide a concise yet powerful summary of a dataset’s informational characteristics, enabling a quantitative assessment of its average complexity and internal dynamics.
+Appendix H Showcases
+H.1 Showcases of Synthetic Data
+Figure 7 showcases representative examples generated by the algorithm using the synthetic dataset.
+As detailed in Appendix E.1, the generated synthetic dataset is classified into two categories.
+The first is a composite type, formed from a combination of seasonal, trend, and noise components, which is designated Custom in the figures.
+The second consists of idealized industrial signals, designated Perfect periodic.
+Figure 7: Several distinct case types are generated by the synthetic dataset algorithm.
+Specifically, the Custom designation refers to a composite signal type, which is systematically constructed by combining seasonal, trend, and noise components.
+Concurrently, the Perfect periodic designation denotes synthesized, idealized industrial signals.
+H.2 Showcases of Kairos
 We present several prediction examples generated by Kairos during testing, as illustrated in Figure 8.
@@
 Figure 8: Example of forecasts from $\textsc{Kairos}_{b}$ on the test datasets used in experiments.
-Appendix H Broader Impacts
-Our work on Kairos is foundational research focused on advancing time series modeling.
-While Kairos itself is not designed for direct societal applications with immediate negative impacts, any powerful predictive technology could be misused.
-So beyond general risks of advanced AI, our model has no specific negative societal impacts need to be discussed here.

### v3

@@
-
+Kairos: Toward Adaptive and Parameter-Efficient Time Series Foundation Models
 Abstract
@@
 Existing TSFMs predominantly rely on massive parameterization to absorb such heterogeneity, as their static tokenization and positional encoding schemes entangle diverse temporal patterns into a fixed representation space, encouraging memorization rather than adaptation.
-To address this limitation, we propose Kairos, a flexible and parameter-efficient TSFM that decouples temporal heterogeneity from model capacity through a novel tokenization perspective.
+To address this limitation, we propose Kairos, a flexible and parameter-efficient TSFM dedicated to forecasting tasks, which decouples temporal heterogeneity from model capacity through a novel tokenization perspective.
 Kairos introduces a dynamic patching tokenizer and a mixture-of-size encoding that adapt observational granularity to local information density, enabling fine-grained temporal abstraction without increasing model width or depth.
@@
 The project page is at https://foundation-model-research.github.io/Kairos.
-Figure 1: (a) Comparison on the GIFT-Eval benchmark (Aksu et al., 2024) demonstrates that our Kairos achieves superior zero-shot forecasting performance (lower normalized MASE) while requiring significantly fewer parameters than existing TSFMs.
+Figure 1: (a) Comparison on the GIFT-Eval benchmark [3] demonstrates that our Kairos achieves superior zero-shot forecasting performance (lower normalized MASE) while requiring significantly fewer parameters than existing TSFMs.
 (b) (c) Significant variation exists in information density (i.e., signal complexity) across and within different time series datasets.
@@
 1 Introduction
-Time series forecasting is a core component of many real-world applications (Box et al., 2015; Wang et al., 2024b).
+Time series forecasting is a core component of many real-world applications [6, 39].
 Although time series data are ubiquitous, many individual datasets suffer from scarcity, making it difficult to train models independently, particularly for deep neural networks that require large amounts of data.
-This challenge has long motivated the development of few-shot and zero-shot forecasting methods (Oreshkin et al., 2020).
-More recently, inspired by the success of large-scale foundation models (Achiam et al., 2023; Kirillov et al., 2023), researchers have turned to building Time Series Foundation Models (TSFMs) for forecasting.
-These models leverage vast and diverse time series corpora to acquire strong generalization capabilities, showing promising results across various downstream tasks (Ansari et al., 2024; Woo et al., 2024; Shi et al., 2024).
-Despite this progress, a key limitation persists: current TSFMs rely on non-adaptive processing pipelines that fail to reflect the dynamic and heterogeneous nature of time series, severely compromising their parameter efficiency.
-As shown in Figure 1(b)(c), the statistical information density (details in Appendix G) varies not only across datasets but also across segments within the same series.
+This challenge has long motivated the development of few-shot and zero-shot forecasting methods [30].
+More recently, inspired by the success of large-scale foundation models [1, 21], researchers have turned to building Time Series Foundation Models (TSFMs) for forecasting.
+These models leverage vast and diverse time series corpora to acquire strong generalization capabilities, showing promising results across various downstream tasks [4, 40, 34].
+Despite this progress, a key limitation persists: current TSFMs for forecasting rely on non-adaptive processing pipelines that fail to reflect the dynamic and heterogeneous nature of time series, severely compromising their parameter efficiency.
+As shown in Figure 1(b)(c), the statistical information density (details in Appendix K) varies not only across datasets but also across segments within the same series.
 By applying rigid architectures (exemplified in Figure 1(d)) to this inherently heterogeneous data, models are forced to use massive parameters to memorize specific patterns rather than acquiring adaptive, generalizable rules.
-Unfortunately, this strategy diminishes the overall utility of the model’s capacity, rendering it less effective for real-world deployment.
+Unfortunately, this strategy diminishes the overall utility of the model’s capacity, rendering it less effective for real-world forecasting applications.
 This deficiency manifests primarily in two fundamental aspects of model design: (i) Representation learning: TSFMs have yet to adopt dynamic tokenization strategies analogous to those in Large Language Models (LLMs).
-While LLMs utilize subword tokenization (Gage, 1994) to efficiently compress semantic units of varying lengths, TSFMs predominantly persist with static patching, whether through point-wise tokenization (Ansari et al., 2024; Shi et al., 2024) or fixed-size patches (Das et al., 2024).
-By coupling token generation to fixed time intervals rather than information content, these approaches forcibly compress diverse temporal variations into a rigid representation space.
+While LLMs utilize subword tokenization [14] to efficiently compress semantic units of varying lengths, TSFMs predominantly persist with static patching, whether through point-wise tokenization [4, 34] or fixed-size patches [11].
+By coupling token generation to fixed time intervals rather than information content, these approaches compress diverse temporal variations into a rigid representation space.
 Consequently, the model wastes computational capacity on low-information segments while struggling to capture high-frequency details within the same fixed budget, leading to significant parameter inefficiency.
 (ii) Temporal dependency modeling: Conventional positional encodings in TSFMs impose a unified temporal scale across sequences, neglecting differences in periodicity, trend and seasonality.
-As a result, existing TSFMs like (Liu et al., 2025) struggle to adapt to heterogeneous domains, for example, power consumption data measured hourly versus retail sales data measured daily, each exhibiting distinct temporal dynamics, as we empirically validate in Section 4.3.
+As a result, existing TSFMs struggle to adapt to heterogeneous domains, for example, power consumption data measured hourly versus retail sales data measured daily, each exhibiting distinct temporal dynamics, as we validate in Section 4.3.
 Consequently, the model requires a deeper, more complex architecture to compensate for this misalignment, further exacerbating parameter inefficiency and hindering practical utility.
@@
 Kairos addresses time series heterogeneity through two primary mechanisms: (i) For intra-patch representation learning, the Mixture-of-Size Encoder adaptively models time series at multiple granularities based on local characteristics.
-Inspired by null experts in Mixture-of-Experts (Zeng et al., 2024; Jin et al., 2024), we introduce null patch sizes to dynamically adjust the number of active granularities.
+Inspired by null experts in Mixture-of-Experts [43, 20], we introduce computation-free null experts to dynamically adjust the number of active granularities.
 This approach moves beyond static tokenization strategies, enabling more expressive modeling of nuanced, time-varying dynamics.
 (ii) For inter-patch dependency modeling, we enhance the Transformer with Dynamic Rotary Position Embedding (DRoPE).
-Unlike standard RoPE (Su et al., 2024), which represents positional information with a fixed temporal scale, DRoPE modulates temporal scales using spectral features and granularity to generate instance-specific positional encodings.
+Unlike standard RoPE [35], which represents positional information with a fixed temporal scale, DRoPE modulates temporal scales using spectral features and granularity to generate instance-specific positional encodings.
 Flexible intra-patch and inter-patch modeling enable Kairos to effectively characterize heterogeneous time series with a substantially reduced parameter count (see Figure 1(a)).
 Additionally, the Multi-Patch Decoder uses learnable tokens to predict future patches in parallel, which mitigates cumulative errors in autoregressive generation and offers greater flexibility for variable-length prediction horizons.
-To further enhance training, we construct the Predictability-Stratified Time Series (PreSTS) corpus, a large-scale and diverse pretraining dataset curated through a targeted sampling strategy.
+To complement our architecture with high-quality supervision, we construct the Predictability-Stratified Time Series (PreSTS) corpus, a large-scale and diverse pretraining dataset curated through a targeted sampling strategy.
 By prioritizing more predictable sequences while maintaining broad coverage, PreSTS provides high-quality supervision that supports efficient model scaling.
-As demonstrated by the results on the GIFT-Eval benchmark (Aksu et al., 2024) in Figure 1(a), these combined innovations allow our Kairos to achieve superior performance while using fewer parameters.
-This strong performance is consistently observed on Time-Series-Library (TSLib) (Wang et al., 2024b) benchmarks as well.
-In summary, our main contributions are as follows: (i) We present Kairos, a parameter-efficient time series foundation model designed to handle the dynamic and heterogeneous nature of time series data.
-(ii) We introduce three key architectural innovations: the Mixture-of-Size Encoder for adaptive multi-granularity tokenization, DRoPE for instance-specific temporal calibration, and the Multi-Patch Decoder for flexible and efficient parallel forecasting.
+As demonstrated by the results on the GIFT-Eval benchmark [3] in Figure 1(a), these combined innovations allow our Kairos to achieve superior performance while using fewer parameters.
+This strong performance is consistently observed on Time-Series-Library (TSLib) [39] benchmarks as well.
+In summary, our main contributions are as follows: (i) We present Kairos, a parameter-efficient time series foundation model that introduces a novel architectural framework to handle the dynamic and heterogeneous nature of time series data.
+(ii) We curate the PreSTS corpus, a large-scale pre-training dataset of over 300 billion time points, featuring a predictability-based sampling strategy to provide high-quality and diverse supervision.
 (iii) Extensive empirical evaluations on the GIFT-Eval and Time-Series-Library benchmarks confirm the efficacy of Kairos, highlighting its superior zero-shot forecasting performance.
 2 Related Work
-Time Series Foundation Models (TSFMs).
+Time Series Foundation Models (TSFMs) for Forecasting.
 Inspired by the strong generalization of Large Language Models (LLMs), researchers have devoted increasing attention to TSFMs.
-Existing approaches typically involve adapting LLMs to time series (Gruver et al., 2023), training scalable Transformers on large corpora (Ansari et al., 2024; Das et al., 2024; Woo et al., 2024; Cohen et al., 2025; Liu et al., 2025), or exploring alternative architectures, including non-Transformer models (Auer et al., 2025; Graf et al., 2025) and lightweight Multi-Layer Perceptrons (Ekambaram et al., 2024).
+Existing approaches typically involve adapting LLMs to time series [18], training scalable Transformers on large corpora [4, 11, 40, 10, 28], or exploring alternative architectures, including non-Transformer models [5, 17] and lightweight Multi-Layer Perceptrons [13].
 However, these models often apply fixed tokenization strategies and position embedding to time series from diverse domains with distinct characteristics, limiting their cross-domain generalization.
@@
 Time Series Tokenization.
-Transformer-based models require converting raw data into discrete or continuous tokens (Sennrich et al., 2016; Agarwal et al., 2025).
+Transformer-based models require converting raw data into discrete or continuous tokens [32, 2].
 Current strategies generally fall into two static paradigms: (i) Fixed-size: The dominant paradigm relies on a uniform patch size across the sequence.
-This ranges from fine-grained point-wise tokenization (Ansari et al., 2024; Rasul et al., 2023; Shi et al., 2024) to coarser uni-size patching (Nie et al., 2023), which has been adopted by recent foundation models (Woo et al., 2024; Liu et al., 2025).
+This ranges from fine-grained point-wise tokenization [4, 31, 34] to coarser uni-size patching [29], which has been adopted by recent foundation models [40, 28].
 While the latter improves upon the computational inefficiency of point-wise methods, both remain rigid.
-(ii) Multi-size: To capture broader temporal dynamics, models like Pathformer (Chen et al., 2024) and TTM (Ekambaram et al., 2024) incorporate multiple predefined patch sizes.
-However, these scales are applied statically, rather than adapting to the time series instance.
-All the above approaches partition sequences based on fixed time intervals rather than local information.
-This rigidity forces the model to allocate equal parameters to segments of varying complexity.
-This imbalance inevitably leads to suboptimal parameter efficiency and limited generalization.
+(ii) Multi-size: To capture broader temporal dynamics, models like Pathformer [8] and TTM [13] incorporate multiple predefined patch sizes.
+However, these approaches apply selected granularities uniformly across the entire sequence based on global features rather than local information.
+This rigidity forces the model to allocate equal computational cost to segments of varying complexity.
+This imbalance leads to suboptimal parameter efficiency and limited generalization.
 In contrast, Kairos employs a mixture-of-size strategy.
-By allocating fewer tokens to simple patterns and more to complex regions, Kairos significantly enhances parameter efficiency while maintaining robust generalization.
+By allocating fewer tokens to simple patterns and more to complex regions, Kairos enhances parameter efficiency while maintaining robust generalization.
 Position Embedding.
-Due to the position-unaware characteristic of self-attention mechanism, Transformers Vaswani et al.
-(2017) rely on position embeddings (PEs) to model temporal information.
-However, most TSFMs adopt existing designs for PE in Natural Language Processing (NLP) domains (Das et al., 2024; Shi et al., 2024), which typically emphasize long-term decay through methods like sinusoidal calculations or by suppressing long-range positional information (Vaswani et al., 2017; Su et al., 2024).
+Due to the position-unaware characteristic of self-attention mechanism, Transformers [36] rely on position embeddings (PEs) to model temporal information.
+However, most TSFMs adopt existing designs for PE in Natural Language Processing (NLP) domains [11, 34], which typically emphasize long-term decay through methods like sinusoidal calculations or by suppressing long-range positional information [36, 35].
 By imposing a uniform temporal scale, these PEs struggle to model the heterogeneous temporal relationships of time series.
-Existing adaptive PE methods in NLP (Zheng et al., 2024; Lin et al., 2024) are also insufficient, as they target context extrapolation rather than the complex temporal structures of time series.
-Therefore, ElasTST (Zhang et al., 2024) proposed a tunable RoPE to better adapt to time series data.
+Existing adaptive PE methods in NLP [45, 23] are also insufficient, as they target context extrapolation rather than the complex temporal structures of time series.
+Therefore, ElasTST [44] proposed a tunable RoPE to better adapt to time series data.
 However, its adaptation uses per-dataset training rather than dynamic adjustment.
-To address this limitation, we propose to dynamically modulate PE tailored to the intrinsic characteristics of each input time series, enhancing the parameter efficiency and overall performance of TSFMs.
+To address this limitation, we propose to dynamically modulate PE tailored to the intrinsic characteristics of each input time series, effectively decoupling temporal heterogeneity from model capacity, therefore enhancing the parameter efficiency and overall performance of TSFMs.
 Figure 2: The architecture of Kairos, which includes (i) The Mixture-of-Size Encoder (right) adaptively tokenizes the input series by routing segments to optimal granularity experts and fusing their embeddings into a unified sequence $\boldsymbol{H}$.
@@
 The learning objective is to minimize the discrepancy between the ground truth $\mathbf{Y}$ and the prediction $\hat{\mathbf{Y}}=f_{\phi}(\mathbf{X})$.
-We detail it in Appendix D.1.2.
+We detail it in Appendix H.1.2.
 The architecture of Kairos, shown in Figure 2, consists of three components.
@@
 Granularity.
-For a given time series segment $\mathbf{x}$, we adopt a multi-granularity modeling perspective.
+For a time series segment $\mathbf{x}$, we adopt a multi-granularity perspective.
 Granularity refers to the temporal resolution at which the segment is tokenized and encoded, as defined below.
@@
 To address this challenge, we introduce a dynamic routing mechanism equipped with a lightweight router that adaptively selects an optimal subset of granularity levels for subsequent encoding of a given segment.
-Inspired by the Mixture-of-Experts (MoE) paradigm (Jacobs et al., 1991; Shazeer et al., 2017), we instantiate $G$ granularity-specific experts, each corresponding to a distinct granularity level, together with $Z\in\{1,\dots,K-1\}$ null experts that perform no computation.
+Inspired by the Mixture-of-Experts (MoE) paradigm [19, 33], we instantiate $G$ granularity-specific experts, each corresponding to a distinct granularity level, together with $Z\in\{1,\dots,K-1\}$ computation-free null experts.
 The inclusion of null experts allows the model to implicitly skip encoding at certain granularities, enabling adaptive control over the effective number of active granularities for each input segment.
-The routing weights are computed via a learnable scoring function.
-Specifically, for each expert $i\in\{1,\dots,G+Z\}$, an affinity score is first computed as $s_{i}=\mathbf{w}_{i}^{\top}\mathbf{x},$ which is then normalized with a bias-corrected softmax $\tilde{s}_{i}=\left(\exp(s_{i}+\boldsymbol{b}_{i})\right)/\left(\sum_{j=1}^{G+Z}\exp(s_{j}+\boldsymbol{b}_{j})\right),$ where $\mathbf{w}_{i}$ denotes learnable parameters and $\boldsymbol{b}_{i}$ is a bias term introduced for auxiliary-loss-free load balancing (Liu et al., 2024a) (details in Appendix D.3.1).
-The router thus produces a set of active granularity indices $\mathcal{I}_{\mathrm{valid}}=\operatorname*{argtop\text{-}K}_{i\in[G+Z]}\penalty 10000\ \tilde{s}_{i}\cap\{1,\dots,G\}.$ That is, only the Top-$K$ granularity experts are activated, while null experts contribute no computation even when selected.
-By selectively activating granularity experts and allowing null experts to absorb routing probability mass, Kairos dynamically adapts the effective granularity resolution for each time series segment according to its local information, while maintaining a sparse and parameter-efficient architecture.
+For each expert $i\in\{1,\dots,G+Z\}$, the routing weights are computed via a learnable affinity score $s_{i}=\mathbf{w}_{i}^{\top}\mathbf{x}$, normalized with a bias-corrected softmax $\tilde{s}_{i}=\left(\exp(s_{i}+\boldsymbol{b}_{i})\right)/\left(\sum_{j=1}^{G+Z}\exp(s_{j}+\boldsymbol{b}_{j})\right),$ where $\mathbf{w}_{i}$ denotes learnable parameters and $\boldsymbol{b}_{i}$ is a bias for auxiliary-loss-free load balancing [24] (details in Appendix H.3.1).
+The router then activates the indices $\mathcal{I}_{\mathrm{valid}}=\operatorname*{argtop\text{-}K}_{i\in[G+Z]}\penalty 10000\ \tilde{s}_{i}\cap\{1,\dots,G\}$.
+By allowing null experts to dynamically prune the number of active experts, Kairos dynamically adapts the effective resolution to local information while maintaining a sparse, parameter-efficient architecture.
 Step 2: Granularity-wise Encoding.
@@
 $\quad\boldsymbol{e}_{i,m}=\pi_{i}(\boldsymbol{p}_{i,m})\in\mathbb{R}^{D_{h}},m\in[1,M_{i}]\penalty 10000\ .$ (2)
-where $\pi_{i}$ denotes a granularity-specific projection function, implemented as a two-layer multilayer perceptron (MLP).
+where $\pi_{i}$ denotes a granularity-specific projection function, implemented as a two-layer MLP.
 Step 3: Mixture-of-Size Fusion.
@@
 To enable efficient alignment, we design the patch sizes to be nested, i.e., $P_{i+1}$ is divisible by $P_{i}$.
-Under this constraint, each embedding sequence $\boldsymbol{E}_{i}\in\mathbb{R}^{M_{i}\times D_{h}}$ can be deterministically expanded to the target resolution via repetition (details and an example in Appendix D.3.2): $\tilde{\boldsymbol{E}}_{i}=\mathrm{Expand}(\boldsymbol{E}_{i})\in\mathbb{R}^{M\times D_{h}}.$ The fused representation is obtained by a gated weighted summation over the activated granularities: $\bar{\boldsymbol{E}}=\sum_{i\in\mathcal{I}_{\mathrm{valid}}}\alpha_{i}\,\tilde{\boldsymbol{E}}_{i},\alpha_{i}=\tilde{s}_{i}/{\sum_{j\in\mathcal{I}_{\mathrm{valid}}}\tilde{s}_{j}}.$
+Under this constraint, each embedding sequence $\boldsymbol{E}_{i}\in\mathbb{R}^{M_{i}\times D_{h}}$ can be deterministically expanded to the target resolution via repetition (details and an example in Appendix H.3.2): $\tilde{\boldsymbol{E}}_{i}=\mathrm{Expand}(\boldsymbol{E}_{i})\in\mathbb{R}^{M\times D_{h}}.$ The fused representation is obtained by a gated weighted summation over the activated granularities: $\bar{\boldsymbol{E}}=\sum_{i\in\mathcal{I}_{\mathrm{valid}}}c_{i}\,\tilde{\boldsymbol{E}}_{i},c_{i}=\tilde{s}_{i}/{\sum_{j\in\mathcal{I}_{\mathrm{valid}}}\tilde{s}_{j}}.$
 Sample-level Result. Finally, the representations from all coarse segments are concatenated to form the encoder output
@@
 Unlike prior approaches such as Chen et al.
-(2024); Zhang et al.
-(2024), which impose a static multi-scale partitioning over the entire time series, our method (i) dynamically adapts the tokenization granularity at the segment level.
+[8], Zhang et al.
+[44], which impose a static multi-scale partitioning over the entire time series, our method (i) dynamically adapts the tokenization granularity at the segment level.
 (ii) decouples temporal heterogeneity from a fixed representational scale by explicitly and dynamically assigning patch size via routing and multi-granularity encoding, enabling more effective intra-patch representation learning.
-This adaptive behavior cannot be achieved via globally predefined scales applied statically and is empirically validated in Section 4.4.1.
+This adaptive behavior cannot be achieved via globally predefined scales applied statically and is empirically validated in Section 4.4.1 and Appendix E.
 As a result, varying temporal patterns are modeled via the lightweight Mixture-of-Size Encoder, reducing the burden of subsequent model parameterization.
 3.3 Heterogeneity-Aware Transformer
-Following Mixture-of-Size Encoder, the fused representation $\boldsymbol{H}\in\mathbb{R}^{\tilde{T}\times D_{h}}$ is processed by a Heterogeneity-Aware Transformer, a stack of standard Transformer (Vaswani et al., 2017) encoder blocks, yielding representations $\boldsymbol{H}^{\text{en}}\in\mathbb{R}^{\tilde{T}\times D_{h}}$ that are subsequently fed to the prediction head.
+Following Mixture-of-Size Encoder, the fused representation $\boldsymbol{H}\in\mathbb{R}^{\tilde{T}\times D_{h}}$ is processed by a Heterogeneity-Aware Transformer, a stack of standard Transformer [36] encoder blocks, yielding representations $\boldsymbol{H}^{\text{en}}\in\mathbb{R}^{\tilde{T}\times D_{h}}$ that are subsequently fed to the prediction head.
 To inject temporal information into the self-attention mechanism, we propose Dynamic Rotary Position Embedding (DRoPE).
 Unlike static encodings such as Vaswani et al.
-(2017); Su et al.
-(2024), DRoPE adaptively calibrates temporal structure to account for instance-specific periodicities and the varying physical durations of dynamic patches.
+[36], Su et al.
+[35], DRoPE adaptively calibrates temporal structure to account for instance-specific periodicities and the varying physical durations of dynamic patches.
 Base Structure of DRoPE.
-Our dynamic position encoding is based on the standard Rotary Position Embedding (RoPE) Su et al.
-(2024), where a hidden vector $\boldsymbol{z}\in\mathbb{R}^{D_{h}}$ (a query or key) at token index $t\in[1,\tilde{T}$] is treated as a set of $D_{h}/2$ complex numbers.
+Our dynamic position encoding is based on the standard Rotary Position Embedding (RoPE) [35], where a hidden vector $\boldsymbol{z}\in\mathbb{R}^{D_{h}}$ (a query or key) at token index $t\in[1,\tilde{T}$] is treated as a set of $D_{h}/2$ complex numbers.
 Each pair is rotated by an angle proportional to the position $t$ and a fixed frequency $\theta_{d}$ as below, where $d\in[0,D_{h}/2-1]$ indexes the dimension:
@@
 DRoPE generalizes this formulation by dynamically transforming both the frequency $\theta$ and the token position $t$.
-A detailed derivation of the RoPE calculation is provided in Appendix D.4.
+A detailed derivation of the RoPE calculation is provided in Appendix H.4.
 Instance-Level Frequency Modulation.
-Standard RoPE uses a uniform frequency $\theta$ for all sequences, failing to capture the unique periodic structures of heterogeneous time series (analysis in Theorem D.1).
+Standard RoPE uses a uniform frequency $\theta$ for all sequences, failing to capture the unique periodic structures of heterogeneous time series (analysis in Theorem H.1).
 We propose to modulate the base frequencies using instance-specific spectral features.
 Specifically, given a time series input $\mathbf{X}$, we extract the magnitudes of the first $\omega$ low-frequency components via Fast Fourier Transform, denoted as $\mathbf{X}_{\mathrm{fft}}$.
-A lightweight MLP maps these features to a scaling factor $\gamma$ and a bias $\beta$, to produce adaptive frequencies $\theta_{\mathrm{inst}}$:
-$\theta_{\text{inst},d}=\gamma_{d}\cdot\theta_{d}+\beta_{d}\penalty 10000\ ,\text{where\penalty 10000\ }\gamma,\beta=\mathrm{MLP}(\mathbf{X}_{\mathrm{fft}})$ (5)
-This affine transformation allows Kairos to re-scale the temporal rotation based on the dominant information like trends and seasonalities in the specific time series instance.
+A lightweight MLP maps these features to a scaling factor $\gamma$ and a bias $\beta$ to produce adaptive frequencies $\theta_{\mathrm{inst}}$.
+To accommodate the exponentially decaying nature of RoPE’s base frequencies, we perform this affine transformation in log-space:
+$\log\theta_{\text{inst},d}=\gamma_{d}\cdot\log\theta_{d}+\beta_{d}\penalty 10000\ ,\text{where\penalty 10000\ }\gamma,\beta=\mathrm{MLP}(\mathbf{X}_{\mathrm{fft}})$ (5)
+This log-space adaptation allows Kairos to robustly re-scale the temporal rotation based on the dominant information like trends and seasonalities in the specific time series instance (see Appendix H.4.2 for a detailed discussion).
 Granularity-Aware Position Calibration.
 In our multi-granularity framework, the discrete token index $t$ is no longer an accurate proxy for time because different tokens represent different physical durations, a direct consequence of the granularity routing and tokenization in our Mixture-of-Size Encoder, as described in Section 3.2.
-To maintain temporal consistency across mixed patch sizes, we replace the integer index $t$ with a calibrated position $\tilde{t}=\sum_{k=1}^{t-1}{\bar{P}_{k}}/{\bar{P}_{\mathrm{g\_min}}}$, where $\bar{P}_{k}$ is the patch size of the $k$-th token and $\bar{P}_{\mathrm{g\_min}}=\min\{P_{i}\}_{i=1}^{G}$ is the globally minimum patch size of the finest granularity.
-The resulting DRoPE operation unifies these two components, rotating the hidden vector $\boldsymbol{z}$ at position $t$ as follows:
+To maintain temporal consistency across mixed patch sizes, we replace the index $t$ with a calibrated position $\tilde{t}=\sum_{k=1}^{t-1}{\bar{P}_{k}}/{\bar{P}_{\mathrm{g\_min}}}$, where $\bar{P}_{k}$ is the patch size of the $k$-th token and $\bar{P}_{\mathrm{g\_min}}=\min\{P_{i}\}_{i=1}^{G}$ is the globally minimum patch size of the finest granularity.
+Unifying these two components, rotating the hidden vector $\boldsymbol{z}$ at position $t$ as follows:
 $f_{\text{DRoPE}}(\boldsymbol{z},t)=(\boldsymbol{z}_{2d}+i\boldsymbol{z}_{2d+1})e^{i\cdot\tilde{t}\cdot\theta_{\text{inst},d}}$ (6)
@@
 Multi-Patch Decoder processes these token embeddings with the latent representations $\boldsymbol{H}^{\text{en}}\in\mathbb{R}^{\tilde{T}\times D_{h}}$ from the above Heterogeneity-Aware Transformer through cross-attention, to produce a sequence of latent forecasting embeddings $\boldsymbol{h}^{\text{de}}\in\mathbb{R}^{J\times D_{h}}$.
-Each embedding $\boldsymbol{h}^{\text{de}}_{j}$ is then projected into the observation space via a shared prediction head, which is implemented as a residual FFN block Das et al.
-(2024), to generate a future patch of length $H_{s}=\lceil H/J\rceil$ as
+Each embedding $\boldsymbol{h}^{\text{de}}_{j}$ is then projected into the observation space via a shared prediction head, which is implemented as a residual FFN block [11], to generate a future patch of length $H_{s}=\lceil H/J\rceil$ as
 $\hat{\mathbf{Y}}_{(j-1)H_{s}+1:jH_{s}}=\text{ResidualFFN}(\boldsymbol{h}^{\text{de}}_{j}),\quad j\in[1,J]$ (7)
-This multi-token design is strategically for two objectives regarding predominant forecasting paradigms.
-(i) Error mitigation and efficiency: Compared to conventional single-step autoregressive models (Ansari et al., 2024; Liu et al., 2024c), predicting $J$ patches in parallel improves computational efficiency and mitigates cumulative error by reducing iterative decoding steps.
-(ii) Horizon flexibility: Unlike models that generate a single larger patch (Das et al., 2024), our approach offers superior granularity for variable prediction horizons.
-By assembling the forecast from the specific subset of tokens required, Kairos ensures high-fidelity alignment with the target length and avoids the imprecise truncation associated with oversized outputs.
-We provide a detailed analysis of this design in Appendix B.
-By combining the Mixture-of-Size Encoder for adaptive input processing with this Multi-Patch Decoder for flexible output generation, Kairos achieves a unified and parameter-efficient framework for zero-shot time series forecasting.
+This multi-token design serves two primary objectives: (i) error mitigation and computational efficiency via parallel patch prediction, and (ii) horizon flexibility for exact alignment with various target lengths.
+We provide a detailed comparative analysis of this design in Appendix D.
+By pairing this flexible Multi-Patch Decoder with the adaptive Mixture-of-Size Encoder, Kairos achieves a unified and parameter-efficient framework for zero-shot time series forecasting.
 3.5 Training Corpus
@@
 PreSTS integrates large-scale real-world time series from diverse domains and a complementary synthetic dataset to ensure broad coverage.
-All test sets from the GIFT-Eval (Aksu et al., 2024) and TSLib (Wang et al., 2024b) benchmarks are explicitly excluded to prevent data leakage.
-While dataset diversity is widely regarded as essential for TSFM training, recent studies indicate that low-predictability or anomalous sequences can substantially impair forecasting (Cheng et al., 2024).
+All test sets from the GIFT-Eval [3] and TSLib [39] benchmarks are explicitly excluded to prevent data leakage.
+While dataset diversity is widely regarded as essential for TSFM training, recent studies indicate that low-predictability or anomalous sequences can substantially impair forecasting [9].
 To address this, we stratify real-world datasets into five tiers based on predictability and assign higher sampling probabilities to more predictable data during training.
 This strategy promotes stable and high-quality pre-training without compromising data diversity.
-Further details are provided in Appendix E.1.
+Further details on dataset construction, along with an experimental analysis decoupling the contribution of the corpus from our model architecture, are provided in Appendix I.1 and Appendix G, respectively.
 4 Experiment
@@
 Table 1: Performance evaluation on the GIFT-Eval.
-We evaluate using normalized MASE and CRPS (Section 4.1.1).
+We evaluate using normalized MASE and CRPS (Section 4.1).
 Baselines encompass statistical, Deep Learning (DL), and TSFM methods, with TSFMs categorized into TestData Leakage and Zero-Shot following official protocols.
@@
 The subscripts $l$, $b$, $s$, $m$, and $a$ represent model sizes of large, base, small, mini and advanced, respectively.
-The complete experimental results are presented in Appendix F.
+The complete experimental results are presented in Appendix J.
 4.1 Experiment Settings
-4.1.1 Evaluation Datasets and Metric
-To comprehensively evaluate our proposed method, we conducted zero-shot assessments on two well-known public benchmarks: GIFT-Eval (Aksu et al., 2024) and Time-Series-Library (TSLib) (Wang et al., 2024b).
+To comprehensively evaluate our proposed method, we conducted zero-shot assessments on two well-known public benchmarks: GIFT-Eval [3] and Time-Series-Library (TSLib) [39].
 We train Kairos in three sizes: mini (10M), small (23M) and base (53M).
-The training details are provided in Appendix D.1.
+The training details are provided in Appendix H.1.
 Benchmarks.
 The GIFT-Eval benchmark comprises 97 tasks spanning short- (55), medium- (21), and long-term (21) horizons, facilitating a comprehensive assessment.
-For TSLib, we selected the ETT and Weather (Zhou et al., 2021) datasets, consistent with Time-MoE (Shi et al., 2024).
-To test adaptability to diverse frequencies, we also included the Saugeen datasets (Godahewa et al., 2021), denoted as Saugeen (D) and Saugeen (W) for daily and weekly intervals, respectively.
-The detailed description of the evaluation datasets, splitting configurations, and evaluation length selection are provided in Appendix E.2 and E.3.
+For TSLib, we selected the ETT and Weather [46] datasets, consistent with Time-MoE [34].
+To test adaptability to diverse frequencies, we included the Saugeen datasets [15], denoted as Saugeen (D) and Saugeen (W) for daily and weekly intervals, respectively.
+Details of the compared methods, evaluation datasets, splitting configurations, and evaluation length selection are provided in Appendices I.2, I.3, and I.4.
 Metric.
-In our evaluation of GIFT-Eval, consistent with prior research (Auer et al., 2025; Liu et al., 2025), we employ the official Mean Absolute Scaled Error (MASE) and Continuous Ranked Probability Score (CRPS) metrics to assess point and probabilistic forecasting performance, respectively.
+In our evaluation of GIFT-Eval, consistent with prior research [5, 28], we employ the official Mean Absolute Scaled Error (MASE) and Continuous Ranked Probability Score (CRPS) metrics to assess point and probabilistic forecasting performance, respectively.
 These metrics are first normalized by the Seasonal Naïve baseline for each individual task.
 The final score is then computed as the geometric mean of these normalized values.
-For the TSLib benchmark, we followed previous works (Liu et al., 2023; 2024c; Nie et al., 2023) by adopting Mean Squared Error (MSE) and Mean Absolute Error (MAE) to evaluate time series forecasting performance.
-The metric calculations are detailed in Appendix D.2.1.
-4.1.2 Compared Methods
-We compare Kairos with several state-of-the-art models, including TSFMs such as Sundial (Liu et al., 2025), Toto (Cohen et al., 2025), YingLong (Wang et al., 2025), VisionTS (Chen et al., 2025), Time-MoE (Shi et al., 2024), ChronosBolt (Ansari et al., 2024), TTM (Ekambaram et al., 2024), Moirai (Woo et al., 2024), Timer-XL (Liu et al., 2024c), TimesFM-2.0 (Das et al., 2024), Chronos (Ansari et al., 2024), and advanced full-shot deep learning models, including PatchTST (Nie et al., 2023), DLinear (Zeng et al., 2022), iTransformer (Liu et al., 2023), Pathformer (Chen et al., 2024), and TimesNet (Wu et al., 2022).
+For the TSLib benchmark, we followed previous works [26, 27, 29] by adopting Mean Squared Error (MSE) and Mean Absolute Error (MAE) to evaluate time series forecasting performance.
+The metric calculations are detailed in Appendix H.2.1.
 4.2 Zero-shot Evaluation
-Finding 1: Kairos achieves superior zero-shot forecasting performance with significantly higher parameter efficiency.
+Finding 1: Kairos achieves superior zero-shot forecasting performance with significantly higher parameter efficiency and learns highly transferable representations..
 On the GIFT-Eval benchmark, as presented in Table 1, $\textsc{Kairos}_{\text{base}}$ (53M) achieves the best MASE and the second-best CRPS among state-of-the-art task-specific deep learning methods and other TSFMs.
@@
 These results confirm that Kairos delivers a significant improvement in predictive capability without relying on massive scale.
+Furthermore, we demonstrate in Appendix C that these learned representations are highly transferable to downstream classification tasks.
 4.3 Ablation Study
-Finding 2: The synergistic integration of Mixture-of-Size Encoding, DRoPE, and the Multi-Patch Decoding is critical for adapting to heterogeneous time series and achieving superior forecasting performance.
-Table 2 details the component-wise contributions.
-(i) Encoder Design: Kairos outperforms variants using fixed patch sizes or excluding null experts, confirming that dynamic, sparse multi-granularity modeling is critical for capturing diverse temporal patterns.
-(ii) Positional Embedding: DRoPE proves superior to standard RoPE (Su et al., 2024).
-Our ablation further isolates the gains from Instance-Level Frequency Modulation and Granularity-Aware Position Calibration, verifying the benefit of adapting to both spectral heterogeneity and physical time distortion caused by dynamic patching.
-(iii) Decoding Strategy: The Multi-Patch Decoding shows clear superiority over standard single-patch autoregressive decoding.
-The integration of both components yielded the best performance across 97 GIFT-Eval tasks, which validates the superiority of our designs.
-Furthermore, the performance gap widens as the forecast horizon increases, demonstrating Kairos’s superior long-term forecasting capability.
 Table 2: Ablation study comparing encoder design, positional embedding and decoder strategy.
-Models were evaluated using the normalized MASE (detailed in Section 4.1.1) across individual prediction horizons and as an aggregate across all tasks.
+Models were evaluated using the normalized MASE (detailed in Section 4.1) across prediction horizons and as an aggregate across all tasks.
 Model Variants Short Medium Long AVG
@@
 […]
+Finding 2: The synergistic integration of Mixture-of-Size Encoding, DRoPE, and the Multi-Patch Decoding is critical for adapting to heterogeneous time series and achieving superior forecasting performance.
+Table 2 details the component-wise contributions.
+(i) Encoder Design: Kairos outperforms variants using fixed patch sizes or excluding null experts, confirming that dynamic, sparse multi-granularity modeling is critical for capturing diverse temporal patterns.
+(ii) Positional Embedding: DRoPE proves superior to standard RoPE [35].
+Our ablation further isolates the gains from Instance-Level Frequency Modulation and Granularity-Aware Position Calibration, verifying the benefit of adapting to both spectral heterogeneity and physical time distortion caused by dynamic patching.
+(iii) Decoding Strategy: The Multi-Patch Decoding shows clear superiority over standard single-patch autoregressive decoding.
+The integration yielded the best performance across 97 GIFT-Eval tasks, which validates the superiority of our designs.
+Furthermore, the performance gap widens as the forecast horizon increases, demonstrating Kairos’s superior long-term forecasting capability.
 4.4 Model Analysis
@@
 This adaptive allocation mechanism reduces computational redundancy on low-information segments, allowing the model to achieve superior performance with a compact parameter budget compared to rigid architectures.
-See Appendix C for a detailed inference speed analysis.
+See Appendix F for a detailed inference speed analysis.
+Furthermore, to demonstrate that Kairos learns genuine segment-level adaptation rather than merely relying on a generic local multi-scale inductive bias, we provide additional experimental evidence in Appendix E.
 4.4.2 DRoPE Analysis
@@
 Finding 4: Modulating RoPE frequencies $\theta$ instance-wise can indeed better model time series temporal relationships.
-To verify that DRoPE’s performance gains stem from its instance-specific $\theta_{\text{inst}}$ modulation, we designed control experiments that disrupt this mechanism (Table 3; Appendix E.4): (i) Intra-Dataset Shuffle: randomly permute $\theta_{\text{inst}}$ across instances within the same dataset; (ii) Inter-Dataset Shuffle: assign $\theta_{\text{inst}}$ from instances of different datasets; (iii) Fixed RoPE: use standard RoPE without any modulation.
+To verify that DRoPE’s performance gains stem from its instance-specific $\theta_{\text{inst}}$ modulation, we designed control experiments that disrupt this mechanism (Table 3; Appendix I.5): (i) Intra-Dataset Shuffle: randomly permute $\theta_{\text{inst}}$ across instances within the same dataset; (ii) Inter-Dataset Shuffle: assign $\theta_{\text{inst}}$ from instances of different datasets; (iii) Fixed RoPE: use standard RoPE without any modulation.
 The results reveal a clear performance gradient (DRoPE $\succ$ Intra-Shuffle $\succ$ Fixed RoPE $\succ$ Inter-Shuffle), indicating that DRoPE learns instance-specific representations to modulate $\theta$.
@@
 However, Intra-Dataset Shuffle still incurs a 1.76% degradation compared to DRoPE, suggesting that instance-level adaptation contributes to the performance gains beyond dataset-level statistics.
-Inter-Dataset Shuffle yields the worst results (28.32% degradation), as $\theta_{\text{inst}}$ from different datasets cannot capture the target series’ temporal dynamics.
+Inter-Dataset Shuffle yields the worst results (28.32% degradation), as $\theta_{\text{inst}}$ from different datasets cannot capture the temporal dynamics of the target series.
 These results collectively validate the effectiveness of instance-level $\theta_{\text{inst}}$ adaptation in DRoPE.
@@
 Extensive experiments demonstrate that Kairos achieves superior performance, validating the necessity of aligning architectural inductive biases with the structural nuances of time series data.
-One limitation of our current approach is that Kairos relies on channel-independent modeling (Nie et al., 2023) to support multivariate time series and does not explicitly capture inter-variable dependencies.
+One limitation of our current approach is that Kairos relies on channel-independent modeling [29] to support multivariate time series and does not explicitly capture inter-variable dependencies.
 Future work aims to model these interactions and extend Kairos to tasks beyond forecasting.
-Impact Statement
-Our work on Kairos is foundational research focused on advancing parameter-efficient and adaptive time series modeling.
-We anticipate several positive societal impacts: (i) promoting sustainable AI by significantly reducing the computational resources and energy required for large-scale deployment; (ii) enhancing decision-making in dynamic environments (e.g., energy, retail) through improved modeling of heterogeneous temporal patterns; and (iii) providing a high-quality data curation methodology via the PreSTS corpus to improve model performance.
-While we do not foresee immediate negative societal consequences, we recognize that advanced predictive technologies carry inherent risks of misuse.
-Consequently, rigorous ethical oversight is therefore essential to guarantee that such predictive capabilities serve the public interest.
-Appendix A Etymology of Kairos
+Appendix A Limitations and Future Work
+While Kairos demonstrates superior parameter efficiency and zero-shot performance, we acknowledge several limitations that provide avenues for future research.
+Channel-Independent Modeling.
+Currently, Kairos adopts a channel-independent modeling approach [29] to handle multivariate time series, treating each variable as an individual sequence without explicitly modeling the relationships between variables.
+However, our architectural innovations are orthogonal to the channel-independence assumption.
+It is easy to incorporate channel dependency modeling for further enhancement.
+For instance, a channel-mixer block can be easily integrated to explicitly capture cross-channel correlations.
+We plan to explore these extensions in future work.
+Evaluation Scope.
+Our empirical evaluation primarily focuses on zero-shot forecasting, which is the primary evaluation setting for TSFMs [28, 10].
+As a generative task, forecasting directly reflects whether a model can capture temporal structures such as trend, seasonality, and local dynamics.
+Furthermore, we provide evaluation results on classification tasks in Appendix C, which demonstrate that Kairos successfully learns transferable representations.
+We plan to further extend Kairos to tasks such as anomaly detection and imputation in future versions.
+Appendix B Etymology of Kairos
 Kairos was chosen for its mythological significance as the Greek god of the “right or critical moment”, reflecting our model’s ability to select the ideal set of granularities and positional encoding for time series characterized by varying information densities and frequencies.
-Furthermore, this continues a thematic tradition seen in other foundational models we benchmark against, such as Chronos (Ansari et al., 2024) (the personification of time) and Moirai (Woo et al., 2024) (the personifications of destiny).
-Appendix B Analysis of Multi-Patch Prediction
+Furthermore, this continues a thematic tradition seen in other foundational models we benchmark against, such as Chronos [4] (the personification of time) and Moirai [40] (the personifications of destiny).
+Appendix C Transferability on Classification Tasks
+To verify whether Kairos learns generalizable features that transfer beyond forecasting tasks, we evaluate its performance on downstream time series classification.
+Experimental Setup.
+Following MOMENT [16], we evaluate on 91 datasets from the UCR archive [12].
+To isolate representation quality, Kairos-Base and MOMENT use a frozen encoder with a lightweight classifier, comparing them against fully fine-tuned baselines (GPT4TS [47], TimesNet [41]).
+Results.
+The classification results are summarized in Table 4.
+Despite relying solely on frozen representations, Kairos consistently outperforms both MOMENT and the fully fine-tuned baselines.
+These results provide evidence that our architectural innovations do not merely overfit to forecasting dynamics.
+Instead, by addressing temporal heterogeneity, Kairos extracts high-level semantic features that are highly transferable to adaptation-heavy settings like classification.
+This solidifies Kairos’s capacity as a general-purpose foundation model capable of supporting diverse downstream time series applications.
+Table 4: Representation transfer results on 91 UCR classification datasets. Baseline results follow MOMENT [16].
+Model Accuracy $\uparrow$ Rank Mean $\downarrow$ Rank Std $\downarrow$
+MOMENT (Frozen) 0.794 1.90 0.80
+GPT4TS (Fine-tuned) 0.567 3.27 0.81
+TimesNet (Fine-tuned) 0.573 3.32 0.74
+[…]
+Appendix D Analysis of Multi-Patch Prediction
 Figure 5: Performance analysis of multi-patch prediction on the GIFT-Eval benchmark across short, medium, and long horizons.
 We vary the number of forecast patches and find that a forecast patch number of 2 achieves the optimal trade-off, resulting in the best overall performance (lowest normalized MASE).
-Unlike task-specific, non-autoregressive models such as TFT (Lim et al., 2021) and Informer (Zhou et al., 2021), which predict all future time points simultaneously, their architecture confines them to forecasting a fixed length identical to the training sequence, precluding variable-length predictions.
+Unlike task-specific, non-autoregressive models such as TFT [22] and Informer [46], which predict all future time points simultaneously, their architecture confines them to forecasting a fixed length identical to the predefined prediction horizon used during training, precluding variable-length predictions during inference.
 Conversely, TSFMs typically employ an iterative process: they predict a segment of the sequence, append it to the historical data, and then forecast the subsequent segment, thereby enabling autoregressive predictions of arbitrary length.
 To mitigate the cumulative errors inherent in this autoregressive methodology, Kairos introduces forecast tokens designed to predict multiple patches within each autoregressive step.
-In this section, we present an analysis of the multi-patch prediction strategy introduced in Section 3.4.
+As briefly mentioned in Section 3.4, the multi-token design of our Multi-Patch Decoder is strategically motivated by two limitations in predominant forecasting paradigms: (i) Error mitigation and efficiency: Compared to conventional single-step autoregressive models [4, 27], predicting $J$ patches in parallel improves computational efficiency and mitigates cumulative error by reducing iterative decoding steps.
+(ii) Horizon flexibility: Unlike models that generate a single larger patch [11], our approach offers superior granularity for variable prediction horizons.
+By assembling the forecast from the specific subset of tokens required, Kairos ensures high-fidelity alignment with the target length and avoids the imprecise truncation associated with oversized outputs.
+To empirically validate this design and determine the optimal setting, we present a detailed analysis of the multi-patch prediction strategy.
 During the training phase, we experimented with a range of forecast patch numbers, specifically $J=\{1,2,4,6,12\}$.
 The corresponding evaluation results on the GIFT-Eval benchmark are illustrated in Figure 5.
-Our observations reveal that when $J=1$, which corresponds to the conventional approach of predicting a single patch (Ansari et al., 2024; Liu et al., 2024c; Das et al., 2024), the model achieves optimal performance in short-term forecasting.
+Our observations reveal that when $J=1$, which corresponds to the conventional approach of predicting a single patch [4, 27, 11], the model achieves optimal performance in short-term forecasting.
 However, this method necessitates multiple iterations of autoregressive prediction, leading to a significant degradation in performance for medium- and long-term forecasting.
@@
 Ultimately, by evaluating the mean normalized MASE, we identified the optimal trade-off, selecting a forecast patch number of $J=2$ for Kairos.
-Appendix C Analysis of Inference Speed
+Appendix E Validation of Learned Segment-Level Adaptation
+To explicitly demonstrate that the performance gains of the Mixture-of-Size Encoder stem from genuine segment-level adaptation rather than merely a generic local multi-scale inductive bias, we conduct two dedicated analyses: a structural comparison with a sequence-level multi-scale baseline, and a direct causal intervention on the learned routing decisions.
+Advantage of Intra-Sequence Adaptation.
+We separate the benefits of local intra-sequence adaptation from global multi-scale availability.
+To achieve this, we replaced our Mixture-of-Size Encoder with the sequence-level multi-scale routing mechanism proposed in Pathformer [8].
+While Pathformer adaptively selects patch sizes for each sequence, it applies these selected granularities uniformly across the entire input, ignoring intra-sequence heterogeneity.
+As shown in Table 5, this replacement degrades GIFT-Eval MASE from 0.738 to 0.759, underscoring the advantage of local intra-sequence adaptation over global sequence-level multi-scale modeling.
+Causal Intervention on Routing Decisions.
+To more directly test whether the router makes meaningful local decisions, we performed causal interventions on the routing mechanism at inference time.
+We kept all model weights and architectural components unchanged, but actively manipulated the router’s outputs:
+  • Uniform Weights: Assigning uniform weights across all granularities within each segment, which removes adaptive selection while preserving the multi-scale structure.
+  • Shuffled Routing: Randomly shuffling the router-produced granularity weights within each segment, which disrupts the learned routing decisions.
+Assigning uniform weights across all granularities degrades performance from 0.738 to 0.831, indicating that the gain cannot be explained by multi-scale modeling alone.
+When we further disrupt the learned routing policy by randomly shuffling the granularity weights, performance deteriorates much more sharply to 1.205.
+Taken together, these results provide direct intervention evidence that Kairos benefits not simply from a generic multi-scale gain, but from assigning granularities appropriately to each segment.
+Table 5: Empirical validation of segment-level adaptation on GIFT-Eval. Results report the normalized MASE across prediction horizons and as an aggregate across all tasks.
+Setting Short Medium Long AVG
+[…]
+Structural Replacement
+   w/ Sequence-Level Routing (Pathformer) 0.730 0.782 0.819 0.759
+Causal Interventions at Inference
+   w/ Uniform Granularity Weights 0.856 0.780 0.819 0.831
+   w/ Shuffled Routing Decisions 1.154 1.282 1.271 1.205
+Appendix F Analysis of Inference Speed
 To provide a comprehensive view of our model’s efficiency, we benchmarked the single-batch inference speed of Kairos against several state-of-the-art models.
 The experimental setup involved an input sequence length of 2048 and a prediction horizon of 96 on a single NVIDIA TITAN RTX GPU.
-For TTM-Advanced, its maximum supported input length of 1536 was used.
-As shown in Table 4, the inference time of Kairos is on the same order of magnitude as other highly efficient models like ChronosBolt and Moirai.
-This demonstrates that Kairos maintains competitive computational efficiency while delivering the superior forecasting accuracy detailed in the main paper.
-Notably, it is significantly faster than models like Time-MoE, TimesFM, and especially Chronos.
-Table 4: Comparison of single-batch inference speeds.
-All models were tested with an input length of 2048 and an output length of 96, with the exception of TTM-Advanced (*), which used its maximum input length of 1536.
-Times are averaged per batch and reported in seconds.
-Model Inference Time (s)
-TTM-Advanced* 0.009
-Timer-XL 0.014
-ChronosBolt-Base 0.055
-Kairos-Base (Ours) 0.061
-Moirai-Large 0.070
-Sundial-Base 0.099
-Time-MoE-Large 0.148
-TimesFM-2.0 0.202
-Chronos-Large 4.901
-Toto-Base 13.717
-Appendix D Implementation Details
-D.1 Training Details
-D.1.1 Model Configurations
+For TTM-Advanced [13], its maximum supported input length of 1536 was used.
+As shown in Table 6, Kairos-Base occupies a superior position on the efficiency-performance frontier.
+Its inference time (0.061s) is on the same order of magnitude as other highly efficient models like ChronosBolt and Moirai, yet it delivers significantly better zero-shot accuracy (0.738 MASE).
+Compared to high-capacity models like Toto and Chronos, Kairos is several orders of magnitude faster while maintaining higher forecasting precision.
+Table 6: Comparison of single-batch inference speeds and zero-shot performance on GIFT-Eval.
+All models were tested with an input length of 2048 and an output length of 96, except for TTM-Advanced (*), which used its maximum input length of 1536.
+Inference times are averaged per batch on a single GPU.
+Model Inference Time (s) $\downarrow$ MASE $\downarrow$
+TTM-Advanced* 0.009 0.812
+Timer-XL 0.014 0.795
+ChronosBolt-Base 0.055 0.781
+[…]
+Moirai-Large 0.070 0.765
+Sundial-Base 0.099 0.750
+Time-MoE-Large 0.148 0.755
+TimesFM-2.0 0.202 0.760
+Chronos-Large 4.901 0.876
+Toto-Base 13.717 0.750
+Appendix G Decoupling Architecture and Data Contributions
+In this section, we provide a dedicated experimental analysis to explicitly decouple the performance contributions of our proposed PreSTS corpus from the architectural innovations of Kairos.
+Ablation on Data Curation.
+We isolate the impact of our data curation components, specifically the predictability-based tier-stratified sampling and the inclusion of synthetic data.
+As shown in Table 7, removing the data curation components degrades the performance on the GIFT-Eval benchmark, confirming their utility in providing high-quality supervision.
+However, this degradation is notably smaller than the performance drop observed when removing the main architectural components.
+This indicates that while the PreSTS data curation strategy provides a complementary gain, the architectural design serves as the primary contributor to the overall performance of Kairos.
+Table 7: Ablation study decoupling data curation and architectural components. Models were evaluated using the normalized MASE on GIFT-Eval.
+Variant MASE $\downarrow$
+Data Curation Ablations
+   w/o tier-stratified sampling 0.748
+   w/o tier-stratified sampling & synthetic data 0.755
+Architectural Ablations
+   w/o Multi-Patch Decoder 0.753
+   w/o Mixture-of-Size Encoder 0.761
+   w/o DRoPE 0.767
+[…]
+Matched-Data Comparison.
+To further disentangle the architecture from the pre-training data, we perform a matched-data comparison.
+Specifically, we train Kairos on the original Chronos corpus [4] and train the advanced baseline ChronosBolt on our proposed PreSTS corpus.
+The results are summarized in Table 8.
+Table 8: Matched-data comparison on the GIFT-Eval benchmark separating architecture and training data contributions.
+Model Params Training Data MASE $\downarrow$
+[…]
+Kairos 53M Chronos corpus 0.761
+ChronosBolt 205M PreSTS 0.781
+ChronosBolt 205M Chronos corpus 0.808
+Chronos 709M Chronos corpus 0.870
+These matched-data results demonstrate that the architectural contribution significantly outweighs the data-curation contribution in our setting:
+[…]
+  • Under the same Chronos corpus, the gap between Kairos and ChronosBolt is 0.047 (0.761 vs.
+0.808), whereas for ChronosBolt the gain from replacing the Chronos corpus with PreSTS is 0.027 (0.808 to 0.781).
+This suggests that, in our experiments, architectural innovation is the primary source of improvement, while PreSTS provides an additional but smaller gain.
+Appendix H Implementation Details
+H.1 Training Details
+H.1.1 Model Configurations
 We train Kairos in three sizes: mini (10M), small (23M) and base (53M) with the detailed model configurations are in Table 9.
@@
 $\textsc{Kairos}_{\text{base}}$ 6 8 512 2048 1408 $\{32,64,128,256\}$ 4 4 2 $\{0.55,0.1,0.05,0.15,0.15\}$ 0.01 128 $10000^{-2j/64}$ 53M
-D.1.2 Loss Function
-To better accommodate varied forecast horizons, and following the methodology of ElasTST (Zhang et al., 2024), we build upon the standard quantile loss by assigning distinct weights to each timestep, such that earlier predictions are given greater importance.
+H.1.2 Loss Function
+To better accommodate varied forecast horizons, and following the methodology of ElasTST [44], we build upon the standard quantile loss by assigning distinct weights to each timestep, such that earlier predictions are given greater importance.
 The model’s parameters are optimized by minimizing the quantile loss with weight decay, formulated as:
-$\mathcal{L}=\frac{1}{B}\sum_{i=1}^{B}\sum_{t=1}^{H}\frac{1}{K}\sum_{k=1}^{K}\omega(t)L_{\alpha_{k}}(y_{i,t},q_{i,t}(\alpha_{k})),$ (8)
-$\omega(t)=\frac{1}{H}(\ln(H)-\ln(t)),\quad L_{\alpha}(y,q)=(\alpha-\mathbf{1}_{\{y<q\}})(y-q),$ (9)
-where $B,H,K$ are the batch size, forecast horizon, and number of quantiles, respectively; $y_{i,t}$ is the ground truth value, $q_{i,t}(\alpha_{k})$ is the $k$-th quantile forecast, $\omega(t)$ is the weight at each time step $t$, and $L_{\alpha_{k}}$ denotes the pinball loss function.
-Empirically, we set $K=9$ and use quantile levels of $\alpha=0.1,0.2,\ldots,0.9.$
-D.2 Evaluation Details
-D.2.1 Metric
+$\mathcal{L}=\frac{1}{B}\sum_{b=1}^{B}\sum_{t=1}^{H}\frac{1}{Q}\sum_{k=1}^{Q}\omega(t)L_{\alpha_{k}}(y_{b,t},q_{b,t}(\alpha_{k})),$ (8)
+$\omega(t)=\frac{1}{H}(\ln(H)-\ln(t^{\prime})),\quad L_{\alpha}(y,q)=(\alpha-\mathbf{1}_{\{y<q\}})(y-q),$ (9)
+where $B,H,Q$ are the batch size, forecast horizon, and total number of quantiles, respectively; $y_{b,t}$ is the ground truth value, $q_{b,t}(\alpha_{k})$ is the $k$-th quantile forecast, $\omega(t)$ is the weight at each time step $t$, and $L_{\alpha_{k}}$ denotes the pinball loss function.
+To prevent the loss weight from evaluating to zero at the final prediction step ($t=H$), we evaluate the logarithmic decay using $t^{\prime}$, which maps the discrete timesteps $t\in\{1,\dots,H\}$ to an evenly spaced continuous sequence from $1+\epsilon_{1}$ to $H-\epsilon_{2}$ (specifically, $\epsilon_{1}=10^{-5}$ and $\epsilon_{2}=10^{-3}$ in our implementation).
+Empirically, we set $Q=9$ and use quantile levels of $\alpha\in\{0.1,0.2,\dots,0.9\}$.
+H.2 Evaluation Details
+H.2.1 Metric
 GIFT-Eval.
@@
 TSLib. We adopt mean square error (MSE) and mean absolute error (MAE) as evaluation metrics. These metrics are calculated as follows:
-$\text{MSE}=\frac{1}{H}\sum_{i=1}^{H}(x_{i}-\hat{x})^{2},\quad\text{MAE}=\frac{1}{H}\sum_{i=1}^{H}\lvert x_{i}-\hat{x}\rvert$ (10)
-where $x_{i}$ is the ground truth and $\hat{x}_{i}$ s the prediction for the $i$-th future time point.
-D.2.2 Stride
-Due to the significant inference latency of our baseline, Chronos (Ansari et al., 2024), we set the evaluation stride to 96 to enhance evaluation efficiency without compromising fairness.
-This setting is consistent with the protocols of Moirai (Woo et al., 2024), Moirai-MoE (Liu et al., 2024b), and the GIFT-Eval benchmark (Aksu et al., 2024), all of which set the stride equal to the prediction length.
+$\text{MSE}=\frac{1}{H}\sum_{i=1}^{H}(x_{i}-\hat{x_{i}})^{2},\quad\text{MAE}=\frac{1}{H}\sum_{i=1}^{H}\lvert x_{i}-\hat{x_{i}}\rvert$ (10)
+where $x_{i}$ is the ground truth and $\hat{x}_{i}$ is the prediction for the $i$-th future time point.
+H.2.2 Stride
+Due to the significant inference latency of our baseline, Chronos [4], we set the evaluation stride to 96 to enhance evaluation efficiency without compromising fairness.
+This setting is consistent with the protocols of Moirai [40], Moirai-MoE [25], and the GIFT-Eval benchmark [3], all of which set the stride equal to the prediction length.
 Additionally, the original Chronos paper evaluates the model on only a single window per dataset.
 Therefore, we wish to emphasize a critical point: while the specific numerical results would likely vary with a different stride, our chosen protocol is applied uniformly to all models under evaluation, ensuring a fair and equitable comparison.
-D.3 Mixture-of-Size Encoder
-D.3.1 Auxiliary-Loss-Free Load Balancing
+H.3 Mixture-of-Size Encoder
+H.3.1 Auxiliary-Loss-Free Load Balancing
 In this section, we elaborate on the details of bias previously introduced in Section 3.2.
-In order to ensure that different experts are adequately trained and to control the distribution ratios of various patch sizes, we employ an Auxiliary-Loss-Free Load Balancing method similar to that used in DeepSeek-V3 (Liu et al., 2024a).
-Specifically, we sum the normalized gating value of the $i$-th expert activated throughout the entire sequence to obtain $L_{i}$:
-$L_{i}=\sum_{n=1}^{N}\alpha_{n,i},$ (11)
-where
-$\alpha_{n,i}=\frac{g_{n,i}}{\sum_{j=1}^{S}g_{n,j}}.$ (12)
-We define a target load distribution $\tau=(\tau_{1},\tau_{2},\ldots,\tau_{S})$, where $\tau_{i}$ is the desired proportion of the total load for expert $i$, satisfying $\sum_{j=1}^{S}\tau_{j}=1$.
-Next, we update the bias term $\boldsymbol{b}_{i}$ using $L_{i}$ and $\tau$:
-$\boldsymbol{b}_{i}\leftarrow\boldsymbol{b}_{i}+\eta_{b}\cdot\frac{\tau_{i}\cdot\sum_{j=1}^{S}L_{j}-L_{i}}{\sum_{j=1}^{S}L_{j}},$ (13)
-where $\eta_{b}$ is a hyper-parameter to govern the magnitude of this adjustment and is referred to as the bias update speed.
-In practical implementation, $\boldsymbol{b}_{i}$ is updated once per batch, i.e., $L_{i}$ denotes the sum of the normalized gating values for the $i$-th expert across all sequences within the entire batch.
-At the end of each step, the bias term $\boldsymbol{b}_{i}$ for each expert is adjusted.
+In order to ensure that different experts are adequately trained and to control the distribution ratios of various patch sizes, we employ an Auxiliary-Loss-Free Load Balancing method similar to that used in DeepSeek-V3 [24].
+Specifically, we compute the empirical load $L_{i}$ for the $i$-th expert by summing its normalized routing weights across all $N$ segments of all $B$ sequences within a training batch:
+$L_{i}=\sum_{b=1}^{B}\sum_{n=1}^{N}\tilde{s}_{b,n,i},$ (11)
+where $\tilde{s}_{b,n,i}$ denotes the routing weights for the $i$-th expert on the $n$-th segment of the $b$-th sequence, which directly corresponds to the definition of $\tilde{s}_{i}$ in Section 3.2.
+We define a target load distribution $\tau=(\tau_{1},\dots,\tau_{G+Z})$, where $\tau_{i}$ is the desired proportion of the total load for expert $i$, satisfying $\sum_{j=1}^{G+Z}\tau_{j}=1$.
+At the end of each training step, we update the bias term $\boldsymbol{b}_{i}$ using the empirical load $L_{i}$ and the target $\tau$:
+$\boldsymbol{b}_{i}\leftarrow\boldsymbol{b}_{i}+\eta_{b}\cdot\frac{\tau_{i}\cdot\sum_{j=1}^{G+Z}L_{j}-L_{i}}{\sum_{j=1}^{G+Z}L_{j}},$ (12)
+where $\eta_{b}$ is a hyper-parameter governing the magnitude of this adjustment, referred to as the bias update speed.
 This dynamic adjustment of $\boldsymbol{b}_{i}$ aims to balance the workload across the experts according to the desired distribution by influencing future top-K selections, while also steering the patches-to-expert affinity scores $s^{\prime}_{n,i}$ in subsequent batches, ensuring that the actual load distribution $L_{i}$ progressively aligns with the target distribution $\tau_{i}$, thereby promoting balanced expert utilization over time.
-D.3.2 Definition of the Expansion Function
+H.3.2 Definition of the Expansion Function
 In this section, we provide the formal definition for the function $\mathrm{Expand}(\cdot)$ within the Mixture-of-Size Fusion stage, as introduced in Section 3.2.
@@
   • Number of null experts: $Z=2$.
-  • Top-$K$ selection: $K=2$.
-Suppose for a specific segment $\mathbf{x}$, the router activates experts corresponding to granularity levels 1 and 3 (i.e., $\mathcal{I}_{\mathrm{valid}}=\{1,3\}$).
+  • Top-$K$ selection: $K=3$.
+Suppose for a specific segment $\mathbf{x}$, the router activates experts corresponding to granularity levels 1 and 3, alongside one null expert (i.e., the selected set is $\{1,3,\text{null}\}$ while $\mathcal{I}_{\mathrm{valid}}=\{1,3\}$).
 The finest activated resolution is $P_{\min}=\min(32,128)=32$.
@@
 By explicitly aligning all representations to the finest resolution $P_{\min}$ via $\mathrm{Expand}(\cdot)$, Kairos ensures that the subsequent weighted summation $\bar{\boldsymbol{E}}=\sum_{i\in\mathcal{I}_{\mathrm{valid}}}\alpha_{i}\,\tilde{\boldsymbol{E}}_{i}$ is performed on element-wise compatible tensors, effectively fusing global context with local details.
-D.4 Dynamic Rotary Position Embedding (DRoPE)
+H.4 Dynamic Rotary Position Embedding (DRoPE)
 In this section, we provide a more detailed description of the DRoPE implementation discussed in Section 3.3.
-We first introduce the principles of Rotary Position Embedding (RoPE) (Su et al., 2024) in detail, and then provide corresponding supplementary information on the DRoPE implementation.
-D.4.1 Details of RoPE
+We first introduce the principles of Rotary Position Embedding (RoPE) [35] in detail, and then provide corresponding supplementary information on the DRoPE implementation.
+H.4.1 Details of RoPE
 The core idea of RoPE is to rotate segments of the query and key vectors by an angle proportional according to their absolute position in the sequence.
@@
 The following theorem demonstrates the periodicity of the attention score within each subspace.
-Theorem D.1 (Periodic dependence of RoPE attention).
+Theorem H.1 (Periodic dependence of RoPE attention).
 Consider the $d$-th 2D subspace of the hidden state under RoPE, with angular frequency $\theta_{d}$, and let $\mathrm{atten}_{d}(m,n)$ denote this subspace’s contribution to the dot-product attention between positions $m$ and $n$.
@@
 ∎
-D.4.2 Details of DRoPE
+H.4.2 Details of DRoPE
 The initial RoPE frequencies $\theta_{d}$ typically exhibit a wide numerical range.
@@
 The layer-specific modulation parameters, $\gamma^{(l)}_{d}$ and $\beta^{(l)}_{d}$, predicted by the Multilayer Perceptron (MLP) for layer $l$ following Algorithm 1, are applied to the log-transformed base frequencies $\log\theta_{d}$ via an element-wise affine transformation as
-$\log\theta^{\prime(l)}_{\text{inst},d}=\gamma^{(l)}_{d}\odot\log\theta_{d}+\beta^{(l)}_{d}$, (30)
+$\log\theta^{\prime(l)}_{\text{inst},d}=\gamma^{(l)}_{d}\cdot\log\theta_{d}+\beta^{(l)}_{d}$, (29)
 where $\log\theta^{\prime(l)}_{\text{inst},d}$ represents the modulated log-frequencies for layer $l$ and dimension pair $d$.
@@
 Algorithm 1 Generating Instance Adaptive Parameters ($\gamma^{(l)},\beta^{(l)}$)
-0: Time series instance $X\in\mathbb{R}^{B\times L}$, mask $M\in\{0,1\}^{B\times L}$, batch size $B$, sequence length $L$, FFT feature dimension $w$, MLP network $f_{\text{MLP}}$
+0: Time series instance $X\in\mathbb{R}^{B\times L}$, mask $M\in\{0,1\}^{B\times L}$, batch size $B$, sequence length $L$, FFT feature dimension $\omega$, MLP network $f_{\text{MLP}}$
 0: Adaptive parameters $\gamma^{(l)}\in\mathbb{R}^{B\times D_{h}/2}$, $\beta^{(l)}\in\mathbb{R}^{B\times D_{h}/2}$
@@
 7:return $\gamma^{(l)},\beta^{(l)}$
-D.4.3 Interpretation of DRoPE
-From the theorem D.1, the base RoPE parameters $\{\theta_{d}\}_{d=0}^{D_{h}/2-1}$ define a fixed bank of sinusoidal kernels over relative positions.
+H.4.3 Interpretation of DRoPE
+From the theorem H.1, the base RoPE parameters $\{\theta_{d}\}_{d=0}^{D_{h}/2-1}$ define a fixed bank of sinusoidal kernels over relative positions.
 Each $\theta_{d}$ controls the frequency of one cosine kernel.
@@
 These parameters modulate the base frequencies in log-space:
-$\log\theta^{\prime(l)}_{\text{inst},d}=\gamma^{(l)}_{d}\odot\log\theta_{d}+\beta^{(l)}_{d},\quad\theta^{\prime(l)}_{\text{inst},d}=\exp\big(\log\theta^{\prime(l)}_{\text{inst},d}\big)$, (32)
+$\log\theta^{\prime(l)}_{\text{inst},d}=\gamma^{(l)}_{d}\cdot\log\theta_{d}+\beta^{(l)}_{d},\quad\theta^{\prime(l)}_{\text{inst},d}=\exp\big(\log\theta^{\prime(l)}_{\text{inst},d}\big)$, (31)
 yielding an adapted frequency grid $\{\theta^{\prime(l)}_{d}\}$ for each layer and sequence.
@@
 Instead, we interpret $\{\theta^{\prime(l)}_{\text{inst},d}\}$ as a sequence-dependent frequency profile that shapes how attention depends on relative lags, providing a more flexible and data-driven positional bias than fixed RoPE.
-Appendix E Additional Details of Experiment Setting
-E.1 Pre-training Datasets
-We trained Kairos on the Predictability-Stratified Time Series (PreSTS) corpus, which consists of over 300B real-world time series observations from Chronos (Ansari et al., 2024) and Moirai (Woo et al., 2024) in conjunction with 15B synthetic time points.
-Following (Das et al., 2024), the training loader samples 80% real data and 20% synthetic data.
+Appendix I Additional Details of Experiment Setting
+I.1 Pre-training Datasets
+We trained Kairos on the Predictability-Stratified Time Series (PreSTS) corpus, which consists of over 300B real-world time series observations from Chronos [4] and Moirai [40] in conjunction with 15B synthetic time points.
+Following [11], the training loader samples 80% real data and 20% synthetic data.
 Real-world data.
@@
 The remaining datasets were classified into Tiers 4 and 5, based on a composite assessment of their size and pattern regularity.
-The specific details of these datasets, categorized by their respective sampling frequencies, are presented in Tables 6-9.
+The specific details of these datasets, categorized by their respective sampling frequencies, are presented in Tables 10-13.
 Synthetic data.
@@
 These feature a constant baseline with repeating events like trapezoidal spikes or inverted-U shaped dips, where the period, amplitude, and width of the events remain fixed across the entire series with no random jitter, as outlined in Algorithm 3.
-Synthetic dataset cases are provided in Appendix H.1.
+Synthetic dataset cases are provided in Appendix L.1.
 Distribution Analysis.
-To investigate whether the synthetic dataset supplements distributions absent in the real-world data, we adopted the statistical methodology of Toto (Cohen et al., 2025).
+To investigate whether the synthetic dataset supplements distributions absent in the real-world data, we adopted the statistical methodology of Toto [10].
 We employed the ARCH-LM Statistic and Spectral Entropy to quantify time-varying volatility and information density (complexity), respectively.
@@
 17:return $\boldsymbol{x}_{1:L}$
-E.2 Evaluation Datasets
+I.2 Compared Methods
+We compare Kairos with several state-of-the-art models, including TSFMs such as Sundial [28], Toto [10], YingLong [38], VisionTS [7], Time-MoE [34], ChronosBolt [4], TTM [13], Moirai [40], Timer-XL [27], TimesFM-2.0 [11], Chronos [4], and advanced full-shot deep learning models, including PatchTST [29], DLinear [42], iTransformer [26], Pathformer [8], and TimesNet [41].
+I.3 Evaluation Datasets
 We select datasets from diverse domains and with varying sampling frequencies as evaluation datasets. The details are summarized in Table 14.
@@
 For the evaluation on the TSLib benchmark, the datasets were split into training, validation, and test sets.
-The split for the ETT and Weather datasets follows the configuration adopted by iTransformer (Liu et al., 2023).
+The split for the ETT and Weather datasets follows the configuration adopted by iTransformer [26].
 All other datasets use a 70%/10%/20% ratio for the training, validation, and test sets, respectively.
-E.3 Evaluation Length Selection
-In this section, we provide further details regarding the selection of historical sequence lengths for the various models evaluated in the TSLib benchmark, supplementing the discussion in Section 4.1.1.
-To accommodate diverse application scenarios, an increasing number of TSFMs (Liu et al., 2025; Das et al., 2024; Liu et al., 2024c; Woo et al., 2024) have devoted attention to predicting over long contexts.
+I.4 Evaluation Length Selection
+In this section, we provide further details regarding the selection of historical sequence lengths for the various models evaluated in the TSLib benchmark, supplementing the discussion in Section 4.1.
+To accommodate diverse application scenarios, an increasing number of TSFMs [28, 11, 27, 40] have devoted attention to predicting over long contexts.
 Consequently, we evaluate Kairos and other TSFMs under a long‑context setting.
@@
 We compared the performance of each model using the context length from its original publication against a length of 2048, selecting the superior of the two.
-To be specific, lengths of 96 and 2048 were compared for DLinear (Zeng et al., 2022), iTransformer (Liu et al., 2023), TimesNet (Wu et al., 2022), and Pathformer, while lengths of 336, 512, and 2048 were compared for PatchTST (Nie et al., 2023).
+To be specific, lengths of 96 and 2048 were compared for DLinear [42], iTransformer [26], TimesNet [41], and Pathformer, while lengths of 336, 512, and 2048 were compared for PatchTST [29].
 The definitive context lengths adopted for each model are systematically tabulated in Table 15.
@@
 Context length {96, 2048} {96, 2048} {96, 2048} {336, 512, 2048} {96, 2048} 512 2048 2048 2048 1536 2048 2048
-E.4 Details of DRoPE Analysis
+I.5 Details of DRoPE Analysis
 In this section, we explain in more detail the setting of the DRoPE analysis experiment discussed in Section 4.4.2.
-To more definitively ascertain whether these instance-specific modulations truly capture and leverage beneficial positional information derived from an instance’s FFT features, we designed a series of ?shuffle?
-experiments.
+To more definitively ascertain whether these instance-specific modulations truly capture and leverage beneficial positional information derived from an instance’s FFT features, we designed a series of “shuffle” experiments.
 These experiments function as a form of causal intervention analysis.
@@
 This process also ensures layer-wise correspondence of the applied modulations.
-  • Fixed RoPE: We use the fully trained DRoPE model but disable its instance-adaptive modulation during inference.
-Specifically, the prediction of $\gamma^{(l)}$ and $\beta^{(l)}$ is bypassed, and all instances revert to using the initial RoPE $\theta$ across all layers.
-This approach keeps the base model architecture and other learned weights identical to DRoPE, isolating the effect of the modulation.
-By comparing DRoPE’s performance against these shuffled and fixed $\theta$ configurations, we can better attribute performance gains or losses to the effectiveness of the instance-adaptive $\theta_{\text{inst}}$ modulation process.
-Appendix F Full Evaluation Results
+  • Fixed RoPE: To evaluate the architectural contribution of our module, this setting serves as a structural baseline trained from scratch.
+We replace the DRoPE module with standard RoPE, entirely removing the parameter-predicting MLP.
+All instances use the static, predefined RoPE $\theta$ across all layers during both training and inference.
+By comparing DRoPE’s performance against the inference-time causal interventions (shuffled configurations) and the fixed $\theta$ baseline, we can definitively attribute the performance gains to the effectiveness of the instance-adaptive $\theta_{\text{inst}}$ modulation process.
+Appendix J Full Evaluation Results
 In this section, we present the detailed results in Section 4.2.
 Table 16 presents the full results of the zero-shot forecasting experiments conducted at each forecast horizon.
-For Time-MoE, we report the results as presented in the original paper (Shi et al., 2024).
+For Time-MoE, we report the results as presented in the original paper [34].
 Table 16: Full results of zero-shot forecasting experiments.
+Best results are highlighted in bold, and second best results are underlined.
+For fair comparison, the full-shot model is excluded from the ranking and is presented for reference only.
+The dash (-) indicates that the dataset was utilized by the model during its pretraining phase.
 Method Metric ETTh1 ETTh2 ETTm1 ETTm2 Weather Saugeen(D) Saugeen(W) AVG Total $1^{\textbf{st}}$
@@
 MAE 0.385 0.414 0.429 0.432 0.415 0.325 0.378 0.395 0.410 0.377 0.321 0.355 0.378 0.405 0.365 0.239 0.283 0.317 0.376 0.304 0.181 0.227 0.270 0.326 0.251 0.406 0.424 0.428 0.437 0.424 0.445 0.446 0.429 0.440 0.365
-Appendix G Calculation of information density
+Appendix K Calculation of information density
 In this section, we present the formula for the information density illustrated in Figure 1(b).
 To effectively characterize and compare different time series datasets, we introduce a method to quantify their information density and the variation of this density over time.
-For this, similar to TimeMixer (Wang et al., 2024a), we utilize spectral entropy, a metric that measures the complexity and compressibility of a signal.
+For this, similar to TimeMixer [37], we utilize spectral entropy, a metric that measures the complexity and compressibility of a signal.
 In this context, a signal with low spectral entropy, such as one with a few dominant periodic components, is considered to have low information density due to its redundant and predictable nature.
@@
 The total number of windows is $N=\lfloor(T-M)/S\rfloor+1$.
-To mitigate spectral leakage from windowing, we apply a Hamming window to each segment $\mathbf{w}_{i}$:
+To mitigate spectral leakage from windowing, we apply a Hann window to each segment $\mathbf{w}_{i}$:
 $h[n]=0.5-0.5\cos\left(\frac{2\pi n}{M-1}\right),\quad 0\leq n\leq M-1.$ (32)
@@
 Together, $\mu_{SE}$ and $\sigma_{SE}$ provide a concise yet powerful summary of a dataset’s informational characteristics, enabling a quantitative assessment of its average complexity and internal dynamics.
-Appendix H Showcases
-H.1 Showcases of Synthetic Data
+Appendix L Showcases
+L.1 Showcases of Synthetic Data
 Figure 7 showcases representative examples generated by the algorithm using the synthetic dataset.
-As detailed in Appendix E.1, the generated synthetic dataset is classified into two categories.
+As detailed in Appendix I.1, the generated synthetic dataset is classified into two categories.
 The first is a composite type, formed from a combination of seasonal, trend, and noise components, which is designated Custom in the figures.
@@
 Concurrently, the Perfect periodic designation denotes synthesized, idealized industrial signals.
-H.2 Showcases of Kairos
+L.2 Showcases of Kairos
 We present several prediction examples generated by Kairos during testing, as illustrated in Figure 8.
@@
 Figure 8: Example of forecasts from $\textsc{Kairos}_{b}$ on the test datasets used in experiments.
+Appendix M Broader impacts
+Our work on Kairos is foundational research focused on advancing parameter-efficient and adaptive time series modeling.
+We anticipate several positive societal impacts: (i) promoting sustainable AI by significantly reducing the computational resources and energy required for large-scale deployment; (ii) enhancing decision-making in dynamic environments (e.g., energy, retail) through improved modeling of heterogeneous temporal patterns; and (iii) providing a high-quality data curation methodology via the PreSTS corpus to improve model performance.
+While we do not foresee immediate negative societal consequences, we recognize that advanced predictive technologies carry inherent risks of misuse.
+Consequently, rigorous ethical oversight is therefore essential to guarantee that such predictive capabilities serve the public interest.